### v1

# Knowledge-guided machine learning model with soil moisture for corn yield prediction under drought conditions

Abstract
Remote sensing (RS) techniques, by enabling non-contact acquisition of extensive ground observations, have become a valuable tool for corn yield prediction. Traditional process-based (PB) models are limited by fixed input features and struggle to incorporate large volumes of RS data. In contrast, machine learning (ML) models are often criticized for being “black boxes” with limited interpretability. To address these limitations, we used Knowledge-Guided Machine Learning (KGML), which combined the strengths of both approaches and fully used RS data. However, previous KGML methods overlooked the crucial role of soil moisture in plant growth. To bridge this gap, we proposed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, using soil moisture as an intermediate variable to emphasize its key role in plant development. Additionally, based on the prior knowledge that the model may overestimate under drought conditions, we designed a drought-aware loss function that penalizes predicted yield in drought-affected areas. Our experiments showed that the KGML-SM model outperformed other ML models. Finally, we explored the relationships between drought, soil moisture, and corn yield prediction, assessing the importance of various features and analyzing how soil moisture impacts corn yield predictions across different regions and time periods.
keywords:
Corn yield prediction; Knowledge-guided Machine learning; Soil moisture; drought
journal: Journal
[…]
[inst1]organization=Biological Systems Engineering,addressline=University of Wisconsin-Madison, city=Madison, postcode=53706, state=WI, country=USA
[…]
[inst2]organization=Department of of Soil Science,addressline=University of Wisconsin-Madison, city=Madison, postcode=53706, state=WI, country=USA
[…]
[inst3]organization=Research and Development Division, National Agricultural Statistics Service, United States Department of Agriculture,addressline=Washington, postcode=20250, state=DC, country=USA
1 Introduction
Corn plays a vital place in the U.S. as a primary crop, supporting food security as well as the animal feed and biofuel industries (Graham et al., 2007; Thompson, 1969). Accurate yield prediction is crucial for effective resource management and economic stability (Kucharik and Ramankutty, 2005). However, achieving precise prediction across large areas and diverse environmental conditions is challenging, making it necessary to pursue advanced research in corn yield prediction techniques (Weiss et al., 2020). Traditional process-based (PB) models (Weiss et al., 2007) are built on a deep understanding of the physical, chemical, and biological processes within crop growth systems (Weiss et al., 2004). These models use equations derived from scientific principles to simulate the behavior of crops, weather, and soil interactions (Weiss et al., 2002), making them highly interpretable and reliable in scenarios where the underlying processes are well understood (Puntel et al., 2016; Shahhosseini et al., 2021). Several promising modeling frameworks are designed for simulating and predicting agricultural production systems, such as The Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), WOrld FOod STudies (WOFOST) (Van Diepen et al., 1989), and The Agricultural Policy / Environmental eXtender (APEX) (Williams and Izaurralde, 2010). Many recent studies (Zhen et al., 2023, 2022) have successfully used these models to predict crop yields. However, PB models have several limitations. Plant growth is highly complex, and these models rely on a limited set of fixed input features, which are insufficient to accurately capture the intricacies of this process. Moreover, PB models require extensive manual parameter tuning, limiting their applicability for large-scale applications across vast regions.
In recent years, remote sensing (RS) has provided several benefits for corn yield prediction (Lobell et al., 2015). It allows for the collection of large-scale and real-time data across vast agricultural areas, which, in conjunction with machine learning (ML) methods, provides an accurate and efficient way for large-scale yield prediction. Deep learning (Goodfellow et al., 2016), a widely used ML method, has the advantage of leveraging vast amounts of RS data by automatically learning features from large datasets. This removes the need for manual feature engineering, enabling the model to capture complex patterns and relationships between environmental factors and crop yield (He et al., 2016). Many studies have applied deep learning to crop yield prediction and shown its advantages. You et al. (2017) introduced a scalable approach to crop yield prediction using modern representation learning rather than traditional hand-crafted features, achieving impressive results. Wang et al. (2018) demonstrated soybean yield predictions in Argentina and obtained reliable results in Brazil using transfer learning, even with limited data. Ma et al. (2021a) developed a county-level corn yield prediction model using a Bayesian Neural Network across 12 states in the Corn Belt, showing the potential for large-scale corn yield prediction tasks. However, despite the popularity of deep learning models in corn yield prediction tasks, they have certain limitations. First, deep learning models require a large amount of data to learn the relationships between inputs and outputs. Second, deep learning models often function as a black box, lacking reliable interpretability of the prediction results. Therefore, finding ways to expand our data sources is a critical factor in achieving accurate predictions. Additionally, there are complex biological processes among soil, weather, and corn, which are challenging for deep learning models to fully capture. Given these challenges, combining PB models with deep learning models has become an important research topic.
Knowledge-guided machine learning (KGML) (Karpatne et al., 2022) aims to integrate scientific knowledge in ML frameworks to achieve better generalizability, scientific consistency, and explainability of results. KGML involves three main approaches: Knowledge-Guided Learning, Knowledge-Guided Architecture, and Knowledge-Guided Pretraining. Knowledge-Guided Learning incorporates scientific knowledge, such as physical laws, into the learning algorithm by modifying the loss function. It guides the optimization process toward scientifically consistent and generalizable solutions (Daw et al., 2022; Bao et al., 2021). Knowledge-Guided Architecture embeds scientific knowledge directly into the structure of the ML model. By designing knowledge-guided neural network architectures, researchers can enable models to automatically capture physical properties or conservation laws, which improves both interpretability and robustness (Dugdale et al., 2017; Luo et al., 2023). Knowledge-Guided Pretraining uses simulated data or self-supervised learning tasks to pretrain ML models. By initializing model parameters using data that encodes scientific principles, this method reduces the need for large amounts of observational data and improves performance in data-scarce scenarios (Licheng et al., 2022; Chen et al., 2023). Building on these approaches, KGML has been increasingly applied in various domains, including agriculture, hydrology, and climate science, enabling improved data-driven decision-making and predictive modeling.
Recently, many studies have used KGML to combine PB models and ML models for crop yield prediction (Kimball et al., 2023; Burroughs et al., 2023). Since PB models already provide insights into the crop growth process and generate high-quality simulated data, Knowledge-Guided Pretraining is the most popular approach (He et al., 2023; Yang et al., 2023). These studies aim to use ML to identify patterns between crop features and yield from large datasets, while also deriving interpretability from the data simulated by PB models. Specifically, KGML models leverage PB models to simulate intermediate variables representing key plant growth processes, enhancing their ability to capture domain-specific knowledge. Then, ML models are applied to simulate specific parts of the process. By incorporating these intermediate variables, the model can better represent the underlying biological and environmental mechanisms. For example, He et al. (2023) integrated physical knowledge from PB models and enabled the model to adjust for temporal data shifts over the years. Yang et al. (2023) used both historical and in-season RS data to better predict key agricultural variables such as grain yield and carbon cycling processes. Yang et al. (2024) presented a knowledge-guided computer vision framework designed to monitor and simulate the growth of individual strawberry fruits, which can predict growth at the individual fruit level by integrating image-based data and dynamic simulations.
Although these KGML models have made significant progress in crop yield prediction, all existing KGML-based studies on crop yield prediction have overlooked the critical role of soil, particularly how soil moisture contributes to plant growth and affects yield. Soil moisture is influenced by temperature, precipitation, and other environmental factors and directly affects crop growth. Unlike RS data, soil moisture has a cumulative impact on yield over time. As a key drought indicator, low soil moisture causes water stress, which limits corn growth and reduces yields (Unganai and Kogan, 1998). Nonetheless, both PB and ML models have certain limitations in modeling soil moisture. In PB models, weather data serve as input, while soil moisture is treated as output (McCown et al., 1996; Williams and Izaurralde, 2010), failing to capture its direct impact on crop yield. In ML models, a key challenge is the limited soil moisture data. For example, as the most commonly used soil moisture dataset, the SPL4SMGP.007 SMAP dataset (Entekhabi et al., 2010) only covers data from 2015 onward, limiting ML models’ ability to learn from extensive historical data and impacting the model performance. Therefore, we integrated PB and ML models through KGML to better model soil moisture for corn yield prediction. Additionally, we examined how soil moisture variations over time and across space affected corn yield in drought-prone areas, where drought often reduced yields and hindered prediction accuracy (Cooper et al., 2014). Researchers have long been interested in studying the relationship between soil moisture, drought, and corn yield prediction (Ines et al., 2013; Bushong et al., 2016; Mladenova et al., 2017). However, these studies either have a narrow scope (Vergopolan et al., 2021), resulting in limited data that cannot be effectively generalized to broader applications, or they focus solely on end-of-season yield prediction (Pignotti et al., 2023).
In this paper, we proposed the KGML-SM framework, which integrated PB and ML models for corn yield prediction while incorporating the role of soil moisture in plant growth. The framework consisted of two key components: a Weather-to-Soil (W2S) encoder that modeled the relationship between weather inputs and soil moisture outputs, and an attention module (Vaswani, 2017) that integrated features from multiple sources and dynamically weighted their influence on corn yield across different growth stages and drought conditions. Additionally, we designed a drought-aware yield prediction loss function, informed by the observed relationship between drought and yield reduction, to further enhance model performance under drought conditions. To train this model, we adopted a two-step process. First, we used the APSIM model to generate simulated soil moisture and corn yield data, which were then combined with the input weather data for pretraining. Then, we finetuned the model with soil moisture data from the SMAP product, corn yield data from official statistics, as well as weather data and vegetation indices (VIs) from Google Earth Engine (GEE) (Gorelick et al., 2017), which enabled it to adapt to real-world variability and enhance predictive accuracy. Our research aimed to improve the accuracy and interpretability of corn yield predictions, which ultimately contributed to more resilient agricultural systems in the face of climate variability.
2 Data acquisition
In this study, we developed two datasets for corn yield prediction. The first, a field-level dataset, was generated using APSIM (McCown et al., 1996) and used for pretraining. The second, a county-level dataset, was derived from GEE (Gorelick et al., 2017) and USDA NASS and employed for finetuning. The workflow was divided into three main steps: simulation, pretraining, and finetuning. A summary of all public datasets used in this study is provided in Table 1. In this section, we first introduce the study area (Sec. 2.1); then, we provide details of the APSIM field-level dataset (Sec. 2.2); finally, we describe the construction of the GEE county-level dataset (Sec. 2.3).
Table 1: Public datasets used in this study. a, b, and c indicate different purposes: a for APSIM simulation, b for model pretraining, and c for model finetuning.
Dataset Use Description References
Iowa Environmental Mesonet a, b In-situ weather data (Herzmann et al., 2004)
PRISM c Gridded weather data (Daly et al., 2015, 2008)
MODIS c Vegetation indices (Justice et al., 1998; Schaaf and Wang, 2015)
MODIS MCD18A1.061 c Radiation data (Wang, 2021)
USDA-CDL c Crop data layer (USDA-NASS, 2017)
SMAP c Soil moisture data (Entekhabi et al., 2010)
USDA NASS yield c County-level yield (USDA, 2020)
2.1 Study area
Our research focused on corn yield prediction across the U.S. Corn Belt, selecting twelve states as our study area: North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan. These states are crucial agricultural regions in the U.S., known for their significant contributions to corn production. We generated a five-year average yield map (Fig. 1) for these twelve states and considered them well suited for corn yield prediction research.
Figure 1: The 5-year average county-level corn yield map in selected states.
2.2 APSIM field-level dataset
In this section, we explain how APSIM was used to generate a field-level simulated dataset for model pretraining. For the simulation of the APSIM field-level dataset, we used data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides U.S. agricultural and environmental data. We downloaded all available station data from the IEM, which included data from thousands of stations across 12 states from 1980 to 2023 (Table 2).
Table 2: The number of stations selected for each state.
State IA IL IN KS MI MN MO ND NE OH SD WI
Station number 113 120 89 150 140 140 130 79 128 100 114 147
2.2.1 Input and output of APSIM
The APSIM model simulates crop yield using four weather data inputs: maximum temperature (Tmax), minimum temperature (Tmin), precipitation (PPT), and radiation (Radn). It also requires certain management parameters, which we describe in detail later (Sec. 3.3.1). Maximum and minimum temperatures influence the rate of plant development and growth stages, impacting processes like photosynthesis and respiration. Precipitation data are essential for modeling soil moisture levels, which directly affect water availability for crops. Radiation is a critical factor in photosynthesis, as it provides the energy needed for plant growth. These weather data are available from the IEM website (Herzmann et al., 2004), which collects and distributes environmental data from various observing networks. The APSIM model takes weather data as input to simulate root zone soil moisture (SM_rootzone), surface soil moisture (SM_surface), and corn yield. SM_rootzone is crucial for corn growth as it directly affects water availability for uptake, influencing plant development and yield. SM_surface plays a key role in seed germination and early growth stages.
2.2.2 Field-level dataset for pretraining
Besides APSIM’s input and output variables, we also extracted two useful features from IEM: prediction year and location (longitude and latitude). Then, we calculated the five-year historical average yield based on the average local simulated corn yield over the past five years. We believed these three features might contribute to improving the accuracy of our corn yield prediction model. After APSIM generated the simulation data, we combined its input weather data, output soil moisture, and corn yield, along with other features to construct the APSIM field-level dataset (Table 3). This dataset was used for pretraining our KGML model, aiming to help the model learn the internal processes of corn growth.
Table 3: Summary of variables comprising the APSIM field-level dataset, including inputs, outputs, and other features
Category Variables
Weather data Radn (MJ/m2) (input)
Tmax (${}^{\circ}\mathrm{C}$) (input)
Tmin (${}^{\circ}\mathrm{C}$) (input)
PPT (mm) (input)
Soil moisture SM_surface (simulated)
SM_rootzone (simulated)
Corn yield Corn yield (t/ha) (simulated)
Others Prediction year
Location (Latitude and longitude)
Historical average corn yield (t/ha) (simulated)
2.3 GEE county-level dataset
In this section, we introduce how to construct a county-level dataset for finetuning from GEE. All the variables in the APSIM field-level dataset were also included. To enhance the ML model with RS data, we included VIs from satellite imagery: Green Chlorophyll Index (GCI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI).
2.3.1 Vegetation indices
The Moderate Resolution Imaging Spectroradiometer (MODIS) dataset (Justice et al., 1998; Schaaf and Wang, 2015) is a vital source of satellite-derived data with 500 m resolution, providing consistent, high-quality observations of the Earth’s surface. MODIS captures data in multiple spectral bands and offers various products like VIs. These indices are essential for monitoring vegetation health, biomass, water content, and chlorophyll levels across different regions and time scales. They help improve yield prediction by indicating vegetation growth and crop health.
The Green Chlorophyll Index (GCI) is used to estimate the chlorophyll content in plants, which directly relates to plant health and productivity. It is calculated using the near-infrared (NIR) and green spectral bands. GCI is valuable for assessing vegetation health, especially for monitoring crop growth and detecting areas with nutrient deficiencies. High chlorophyll content typically indicates a healthy, productive plant, making GCI an essential tool in agricultural management. GCI is defined as:
$GCI=\frac{NIR}{Green}-1$ (1)
where $NIR$ represents the surface reflectance in the Near Infrared band, and $Green$ represents the surface reflectance in the Green band.
The Enhanced Vegetation Index (EVI) is designed to optimize the vegetation signal by enhancing sensitivity in high biomass areas and reducing atmospheric and canopy background noise. It uses the red, blue, and NIR spectral bands. EVI is particularly useful in regions with dense vegetation, where NDVI might saturate. It provides a more accurate representation of vegetation conditions, especially in areas with high leaf area index, making it ideal for monitoring forests and agricultural lands. EVI is defined as:
$EVI=2.5\times\frac{NIR-Red}{NIR+6\times Red-7.5\times Blue+1}$ (2)
where $Red$ and $Blue$ represent the surface reflectances of the Red band and Blue band, respectively.
The Normalized Difference Water Index (NDWI) measures the moisture content in vegetation and soil by using the NIR and shortwave infrared (SWIR) bands. NDWI is crucial for monitoring water stress in plants, assessing drought conditions, and detecting changes in water bodies. It aids in managing irrigation practices and understanding the impact of water availability on crop yields. NDWI is defined as:
$NDWI=\frac{NIR-SWIR}{NIR+SWIR}$ (3)
where $SWIR$ represents the shortwave infrared band.
The Normalized Difference Vegetation Index (NDVI) is one of the most widely used indices for measuring vegetation greenness and biomass. It is calculated using the NIR and red spectral bands. NDVI is a standard tool for monitoring plant growth, assessing vegetation cover, and detecting changes in biomass over time. It is particularly useful for tracking seasonal variations in vegetation and estimating crop yields. NDVI is defined as:
$NDVI=\frac{NIR-RED}{NIR+RED}$ (4)
2.3.2 Weather data
The Parameter-elevation Regressions on Independent Slopes Model (PRISM) dataset (Daly et al., 2015)(Daly et al., 2008) is a high-resolution weather dataset with 4 km resolution that provides detailed information on various climatic variables, including PPT, Tmax, and Tmin. This dataset is widely used in agricultural research, hydrology, and weather studies due to its fine spatial resolution and comprehensive coverage, making it an essential tool for understanding and predicting weather-related impacts on crop yield and other environmental processes.
The MCD18A1 Version 6.1 (Wang, 2021) is a MODIS Terra and Aqua combined Downward Shortwave Radiation gridded Level 3 product that provides us with reliable radiation data (Radn). It is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours. Downward Shortwave Radiation is incident solar radiation over land surfaces in the shortwave spectrum (300-4,000 nanometers) and is an important variable in land-surface models that address a variety of scientific and applied issues.
2.3.3 Soil moisture
The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Entekhabi et al., 2010) plays a critical role in our research on the relationship between drought and corn yield prediction. By providing detailed measurements of soil moisture at both the surface (0–5 cm) and root zone levels (0–100 cm), SMAP data allows us to assess the availability of water in the soil, a key factor influencing crop growth and resilience during drought conditions.
2.3.4 County-level dataset for finetuning
In addition to the primary weather and vegetation parameters, we also incorporated features such as the prediction year, location (latitude and longitude), and the 5-year historical average yield (USDA, 2020) into our GEE county-level dataset, as in the APSIM field-level dataset. By integrating these features, the model gained a more comprehensive understanding of the environmental and temporal factors affecting corn yield, leading to more reliable predictions. All the variables in GEE county-level dataset are listed in Table 4.
Table 4: Summary of variables in GEE county-level dataset
Category Variables Spatial resolution Source
Vegetation index Green Chlorophyll Index (GCI) 500m MODIS
Enhanced Vegetation Index (EVI)
Normalized Difference Water Index (NDWI)
Normalized Difference Vegetation Index (NDVI)
Weather data Radn (W/$m^{2}$)
Tmax (${}^{\circ}\mathrm{C}$) 4km PRISM
Tmin (${}^{\circ}\mathrm{C}$)
PPT (mm)
Soil moisture SM_surface 9km SMAP
SM_rootzone
Corn yield USDA NASS corn yield (t/ha) N/A USDA NASS
Others Prediction year
Location (Latitude and longitude)
Historical average yield (t/ha)
3 Methodology
The overall pipeline of our KGML-SM is shown in Fig. 2. Our deep learning model consisted of two main components: the W2S encoder, which learned the relationship between weather data and soil moisture, and the attention-based feature-weighting module, which captured how various features influence corn yield. We first pretrained the model using the APSIM field-level dataset, then finetuned it with the GEE county-level dataset.
Figure 2: The pipeline of the proposed KGML-SM model.
We begin with a formal Problem formulation of KGML-SM (Sec. 3.1), followed by an introduction to the model components (Sec. 3.2); next, we explain how field-level dataset is generated for pretraining (Sec. 3.3) and county-level dataset for finetuning and testing (Sec. 3.4); finally, we introduce how the KGML model is trained and used for prediction (Sec. 3.5).
3.1 Problem formulation
We formally define the corn yield prediction problem and present its mathematical formulation within a KGML framework. For a unique combination of county and year sample $i$, where $i\in\mathcal{N}$, $\mathcal{N}$ represents all counties across all years in the experiment, we define the input features as follows. The weather features are denoted as $\mathbf{w_{i}}=[w^{1}_{i},w^{2}_{i},\dots,w^{t}_{i}]$, and the other features (prediction year, location, and historical average yield) as $\mathbf{o_{i}}=[o^{1}_{i},o^{2}_{i},\dots,o^{t}_{i}]$, where $t\in\mathcal{T}$ and $\mathcal{T}$ represents the entire period from corn planting to harvest. We also have the simulated soil moisture as $\mathbf{s_{i}}=[s^{1}_{i},s^{2}_{i},\dots,s^{t}_{i}]$ and $y_{i}$ denotes the corn yield of each county in that year as supervision. The APSIM field-level dataset, $\mathcal{D}_{\text{sim}}=\{(\mathbf{w_{i}},\mathbf{o_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$, is used for pretraining. The vertical bar $\mid$ separates the input variables from the target labels.
Additionally, the VIs are represented as $\mathbf{v_{i}}=[v^{1}_{i},v^{2}_{i},\dots,v^{t}_{i}]$. We denote two subsets of the GEE county-level dataset as $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ respectively, where $\mathcal{D}_{\text{train}}$ is used for finetuning and $\mathcal{D}_{\text{test}}$ for testing, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}=\{(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$. Here, we omit the description of the validation set $\mathcal{D}_{\text{val}}$ but introduce its partitioning and usage details in the experimental setup (Sec. 4.1).
The objective is to first learn an encoder $f_{\text{W2S}}$ that maps weather inputs to soil moisture: $\mathbf{\hat{s}_{i}}=f_{\text{W2S}}(\mathbf{w_{i}})$. Then, the predicted soil moisture $\mathbf{\hat{s}_{i}}$ is used along with other input features to predict yield through an attention module $f_{\text{att}}$: $y_{i}=f_{\text{att}}(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}},\mathbf{\hat{s}_{i}})$. The model’s performance is evaluated by comparing the predicted yields $\hat{y}_{i}$ with the actual yields $y_{i}$.
3.2 KGML-SM model structure
3.2.1 Weather-to-Soil encoder
W2S encoder is a module designed to model the influence from weather conditions to soil moisture. By capturing the relationship between weather and soil moisture, the W2S encoder enhances the model’s understanding of soil dynamics and the impact of weather variability on soil conditions and corn yield.
The W2S encoder follows a U-Net-based encoder-decoder structure (Ronneberger et al., 2015), consisting of an encoding function $\mathcal{E}$, a decoding function $\mathcal{D}$, and a final fully connected layer $\mathcal{F}$ for feature transformation. Given an input time-series weather input $\mathbf{w_{i}}$, the encoding process can be formulated as:
$\mathbf{z}_{1}=\mathcal{E}_{1}(\mathbf{w_{i}}),\quad\mathbf{z}_{2}=\mathcal{E}_{2}(\mathcal{P}(\mathbf{z}_{1}))$ (5)
where $\mathcal{E}_{i}$ denotes the $i$-th convolutional block (LeCun et al., 1995), and $\mathcal{P}$ represents the max-pooling operation (Nagi et al., 2011) that reduces the temporal dimension. For simplicity we omit the subscript $i$ for $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$.
The decoding process reconstructs high-level features through upsampling and convolutional layers:
$\mathbf{\tilde{z}}_{2}=\mathcal{U}(\mathbf{z}_{2}),\quad\mathbf{\tilde{z}}_{1}=\mathcal{D}_{1}([\mathbf{\tilde{z}}_{2},\mathbf{z}_{1}])$ (6)
where $\mathcal{U}$ is a transposed convolution (upsampling) operation, $[\mathbf{\hat{z}}_{2},\mathbf{z}_{1}]$ represents skip connections concatenating encoder and decoder feature maps, and $\mathcal{D}_{1}$ is the decoder convolutional block.
The predicted soil moisture $\mathbf{\hat{s}_{i}}$ is obtained via:
$\mathbf{\hat{s}_{i}}=\mathcal{F}(\mathcal{D}_{1}([\mathbf{\tilde{z}}_{2},\mathbf{z}_{1}]))$ (7)
3.2.2 Attention module
The attention mechanism (Vaswani, 2017) is a powerful tool in ML that enables models to focus on the most relevant parts of the input data when making predictions. By assigning different levels of importance, or weights, to various input elements, the attention mechanism helps the model prioritize the most crucial information for the task. In KGML-SM, we aim to use the attention mechanism to weight different features, helping us understand each feature’s contribution to yield prediction across different dimensions.
Specifically, suppose that we have input $x_{i}=[\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}},\mathbf{\hat{s}_{i}}]$ for predicting corn yield. For each feature embedding $x_{i}$, we use the attention mechanism to learn the corresponding weight $\alpha_{i}$, which is then used to compute the final yield $y_{i}$.
First, we compute the query $Q_{i}$, key $K_{i}$, and value $V_{i}$ vectors from the feature $x_{i}$ using learned linear transformations:
$Q_{i}=W_{Q}x_{i},\quad K_{i}=W_{K}x_{i},\quad V_{i}=W_{V}x_{i}$ (8)
where $W_{Q}$, $W_{K}$, and $W_{V}$ are the learned weight matrices for the query, key, and value, respectively.
Next, we calculate the attention scores by taking the dot product of the query and key, scaled by the square root of the key’s dimension $d_{k}$:
$\text{Attention}(Q_{i},K_{i})=\frac{Q_{i}\cdot{K_{i}}^{\top}}{\sqrt{d_{k}}}$ (9)
These attention scores are then passed through a softmax function (Goodfellow et al., 2016) to obtain the weights $\alpha_{i}$:
$\alpha_{i}=\text{softmax}\left(\frac{Q_{i}\cdot{K_{i}}^{\top}}{\sqrt{d_{k}}}\right)$ (10)
And The softmax function (Goodfellow et al., 2016) is defined as:
$\text{softmax}(\mathbf{z})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{K}e^{z_{j}}}$ (11)
Where $\mathbf{z}=[z_{1},\dots,z_{K}]$ is the input vector for $i=1,\dots,K$. $e^{z_{i}}$ is the exponential of the $i$-th element of the input vector $\mathbf{z}$.
Finally, we compute the weighted sum of the values $V_{i}$ to obtain the final yield prediction:
$y_{i}=\sum_{t=1}\alpha_{i}V_{i}$ (12)
This process allows us to capture the importance of each feature $x_{i}$ in predicting the yield, dynamically adjusting the weights based on the learned attention scores.
3.3 Generating field-level dataset for pretraining
3.3.1 APSIM management parameters
In addition to weather input data, the APSIM model requires management parameters for calibration. Simulation methods vary depending on the scale of the study. For field-level crop areas, management information is obtained directly from farmers and used as input, while other unknown parameters are adjusted to optimize the model (Shahhosseini et al., 2021). For county-level crop areas, a common practice is to identify representative fields within those counties, average their management parameters, and apply these averages to represent the entire county (Puntel et al., 2016). When the study area is very large, as in our research on simulating all counties in the U.S. Corn Belt, calibrating each county individually became impractical. In this case, the typical approach is to define a range for each parameter based on empirical and statistical data, and then randomly combine these parameters within the specified ranges to simulate yield across all counties (Lobell et al., 2015). Given the vast area of cornfields in our study, it was not feasible to calibrate these parameters for every individual station. Therefore, we established a general range for these parameters to encompass all possible values across different fields. This approach ensured that the model remained applicable to a wide range of conditions while maintaining a reasonable level of accuracy in its predictions.
We selected an appropriate adjustment range for the most important management parameters (Table 5). Sowing density determines the number of plants per unit area, directly affecting competition for resources such as light, water, and nutrients. Based on statistics from the USDA NASS (USDA, 2020), the range for sowing density was set at 6-9 plants/m2. Sowing dates are crucial as they determine the crop’s growth cycle and its interaction with seasonal weather patterns. Typically, growers maximize corn yield by planting in late April or early May (Licht, 2021; Coulter, 2024). In the simulation, sowing started between April 20–25 and ended between May 15–20. Fertilizer application is essential for providing the necessary nutrients to support plant growth. The most commonly used nitrogen fertilizers for corn production in North America are anhydrous ammonia, urea, and urea-ammonium nitrate solutions (Herzmann et al., 2004). The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N). Initial soil water content is important for establishing the starting conditions for the model’s simulation of soil moisture dynamics throughout the growing season. The initial soil water content was set between 40% and 60%.
Table 5: Summary of management parameters in APSIM simulation
Factor Value range Source
Start of sowing window Apr-20 to Apr-25 (Licht, 2021; Coulter, 2024)
End of sowing window May-15 to May-20 (Licht, 2021; Coulter, 2024; Lobell et al., 2015)
Plant population 6-9 plants/m2 (USDA, 2020)
Fertilizer amount 200-300 kg/ha (Herzmann et al., 2004)
Intial soil water 40%-60% (Lobell et al., 2015)
3.3.2 Filtering simulated data based on soil moisture
PB models overlook certain environmental factors and real-world influences, leading to significant discrepancies between some simulated and real-world data. In the simulation process, the simulated intermediate variable is soil moisture, and the final simulation output is corn yield. It is reasonable to assume that with multiple intermediate variables, the final simulation results will have greater errors than the intermediate variables themselves. Additionally, soil moisture is influenced by external factors in a relatively direct manner, unlike yield, which depends on complex plant growth processes. Therefore, soil moisture was used as a benchmark to filter the simulated dataset. We trained a linear regression (LR) model on the GEE county-level dataset to predict soil moisture from weather data and used it to filter the most accurate simulated data in the APSIM field-level dataset. Simulated soil moisture with higher accuracy and its corresponding corn yield were retained to ensure that the model benefited from the most reliable simulations. Notably, we used only simulated data from years before the test year, eliminating the risk of future information leakage.
3.4 Generating county-level dataset for finetuning and testing
3.4.1 Feature Extraction Within Cropland
After downloading various feature data from GEE, we needed to identify the cropland areas. We used the cropland mask from the Cropland Data Layer (CDL) (USDA-NASS, 2017) to extract the feature data corresponding to cropland. The Cropland Data Layer is an annual, raster-based crop-specific land cover dataset produced by the USDA. It has a resolution of 30 meters and offers detailed crop type and location data across the United States. For our corn yield prediction study in twelve U.S. Corn Belt states, we used the CDL to generate a corn field mask in GEE. This mask identifies the specific areas where corn is grown, allowing us to focus our analysis on relevant agricultural fields.
3.4.2 Data preprocessing
Every 16 days, we aggregated all county-level features into a single value to represent agricultural conditions for that period. This method followed the experimental setup from our previous research (Ma et al., 2021b; Wang et al., 2025). We believed that a 16-day interval was sufficient to capture timely changes in weather and environmental conditions.
3.5 Developing the KGML framework
3.5.1 Pretraining with APSIM field-level dataset
The pretraining process began by using a W2S encoder $f_{W2S}$ to learn the relationship between input weather features $w_{i}$ and predicted soil moisture $\hat{s}_{i}$. This process was formulated as $\hat{s}_{i}=f_{W2S}(w_{i}),\ \text{where}\ w_{i}\in\mathcal{D}_{\text{sim}}$. Then, we used the simulated soil moisture $s_{i}\in\mathcal{D}_{\text{sim}}$ to guide the W2S encoder. The loss function $\mathcal{L}_{\text{SM}}$ was defined as:
$\mathcal{L}_{\text{SM}}=\frac{1}{N}\sum_{i=1}^{N}(s_{i}-\hat{s}_{i})^{2}$ (13)
where $N=\lvert\mathcal{N}\rvert$. Next, we concatenated the predicted soil moisture with weather data and other features to form the features embedding $x_{i}=[w_{i},o_{i},\hat{s}_{i}]$. To predict the final corn yield $\hat{y}_{i}$, we used the attention module $f_{att}$, which took the concatenated features $x_{i}$ as input:
$\hat{y}_{i}=f_{att}(x_{i})$ (14)
The model was trained using the simulated corn yield $y_{i}$ from the dataset $\mathcal{D}_{\text{sim}}$, ensuring that predictions align with PB simulations. The final yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation. It was formulated as:
$\mathcal{L}_{\text{yield}}=\frac{1}{N}\sum_{i=1}^{N}d_{i}\left[(y_{i}-\hat{y}_{i})^{2}+\lambda\max(0,\hat{y}_{i}-y_{i})^{2}\right]$ (15)
where $d_{i}$ is a drought-aware weighting factor, which was defined as:
$d_{i}=\frac{1}{\bar{s}_{i}+\varepsilon}$ (16)
where $\bar{s}_{i}$ is the average soil moisture over the growing season for sample $i$, and $\varepsilon$ is a small constant to prevent numerical instability. We set $\varepsilon=1$ in our experiment. Since soil moisture plays a critical role in crop growth and yield formation, the loss function assigned a higher penalty to drier conditions, encouraging the model to be more responsive to soil moisture variations.
Additionally, we incorporated an asymmetric penalty term (Ridnik et al., 2021) controlled by the factor $\lambda>0$, which amplifies the loss when the predicted yield $\hat{y}_{i}$ exceeds the true yield $y_{i}$:
$\lambda\max(0,\hat{y}_{i}-y_{i})^{2}$ (17)
This asymmetry discourages overestimation, particularly under drought conditions, where yield predictions tend to be more uncertain. By applying a stronger penalty to overestimated yields, the model is encouraged to be more conservative, reducing the risk of unrealistic predictions. We set $\lambda=2$ in our experiment.
Specifically, when $\hat{y}_{i}>y_{i}$, the predicted error is scaled by a penalty factor $\lambda$, amplifying the loss in these cases. This encourages the model to adopt a conservative approach, reducing the likelihood of overestimating yield, particularly in drought-prone regions where overestimation could lead to inaccurate agricultural planning.
This formulation ensured that the model not only learned accurate yield predictions but also captured the impact of soil moisture variability and drought stress, leading to more reliable and interpretable results.
3.5.2 Finetuning and testing with GEE county-level dataset
We used the GEE county-level dataset $\mathcal{D}_{\text{train}}$ for finetuning, which contained real-world input features and corresponding USDA NASS yield records. The loss functions $\mathcal{L}_{\text{yield}}$ and $\mathcal{L}_{\text{SM}}$ were optimized on this dataset. Finally, we evaluated the finetuned model on the testing set $\mathcal{D}_{\text{test}}$, where only the input features were available.
4 Experimental Results
4.1 Experimental setup
We conducted experiments on both traditional ML models and DL models. When predicting corn yield for a specific year, we trained the model using all data from preceding years, then split the dataset into 80% for training and 20% for validation, and tested it on the target year. Each experiment was conducted five times with different random seeds, and the final results represented the average across these runs to ensure robustness and reliability.
We implemented the deep learning models using the PyTorch framework (Paszke et al., 2019) and the traditional ML code with sklearn (Pedregosa et al., 2011). The models were run on A100-SXM4-40GB and A100-SXM4-80GB GPUs. For training, we used the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of 0.001 and the ReduceLROnPlateau scheduler.
Root mean square error ($RMSE$) and the coefficient of determination ($R^{2}$) were used to evaluate the performance of our model. $RMSE$ is a measure of the differences between predicted and actual values in a regression model. It provides an estimate of the standard deviation of the prediction errors. A lower $RMSE$ indicates a better fit of the model to the data. The formula for $RMSE$ is:
$RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (18)
$R^{2}$ is a statistical measure that indicates the proportion of the variance in the dependent variable that is predictable from the independent variables. An $R^{2}$ value of 1 indicates perfect prediction, while a value of 0 indicates that the model does not explain any of the variance in the response variable. The formula for $R^{2}$ is:
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (19)
where $n$ is the number of observations, $y_{i}$ is the actual value for the $i$-th observation, $\hat{y}_{i}$ is the predicted value for the $i$-th observation, and $\bar{y}$ is the mean of the actual values.
4.2 Statistical analysis
Before conducting the experiments, we analyzed the statistics of drought, soil moisture, and corn yield. We investigated drought conditions in the United States (Drought.gov, 2024) during August from 2019 to 2023 and present the corresponding maps (Fig. 3). While we focused on August, drought conditions throughout the entire corn growing season from May to October were generally similar. In 2019 (Fig. 3(a)), severe drought was primarily observed in Iowa and Michigan. In 2020 (Fig. 3(b)), severe drought affected Iowa, with North Dakota, South Dakota, Nebraska, and Illinois facing moderate drought. 2021 (Fig. 3(c)) saw worsening conditions, with severe drought covering large parts of North Dakota, South Dakota, and Minnesota, while Nebraska and Iowa experienced moderate drought. In 2022 (Fig. 3(d)), severe drought persisted in Nebraska and Kansas, with South Dakota and Iowa experiencing moderate drought. Finally, in 2023 (Fig. 3(e)), severe drought was widespread across Nebraska, Kansas, Minnesota, Iowa, and Wisconsin, while Missouri faced moderate drought conditions.
Figure 3: Drought map of the U.S. in August from (a) 2019 to (e) 2023.
Before studying the impact of soil moisture on corn yield and its prediction, we specified the relationship between soil moisture and drought. Fig. 4 showes the maps of rootzone and surface soil moisture in August from 2019 to 2023. The comparison of the drought maps above with soil moisture data from 2019 to 2023 revealed a strong correlation between drought-affected areas and lower soil moisture levels. We also noticed that rootzone moisture was more abundant than surface moisture and that the two spatial distributions were generally similar.
Figure 4: This figure shows the maps of (1) rootzone and (2) surface soil moisture in August from (a) 2019 to (e) 2023.
Next, we examined the specific impact of drought on corn yield. We plotted the difference in corn yield for each year from 2019 to 2023 relative to the previous year’s yield (Fig. 5), where negative values indicate a yield reduction. In 2019 (Fig. 5(a)), drought-affected areas included Iowa and Michigan, however, significant yield reductions were observed only in Michigan. In 2020 (Fig. 5(b)), Iowa faced drought along with a derecho storm (Hosseini et al., 2020), leading to noticeable yield losses. In 2021 (Fig. 5(c)), drought impacted North Dakota, South Dakota, Minnesota, Nebraska, and Iowa, yet significant yield reductions occurred only in North Dakota and South Dakota. In 2022 (Fig. 5(d)), drought-affected states included Nebraska, Kansas, South Dakota, and Iowa, but yield declines were primarily observed in North Dakota, Kansas, and South Dakota. In 2023 (Fig. 5(e)), although Nebraska and Kansas continued to experience drought, corn yield did not decrease significantly compared to the previous year, likely due to the already reduced yield in 2022. However, Iowa and Missouri experienced some yield losses due to ongoing drought conditions. Thus, we concluded that drought consistently led to a decline in corn yield. We summarized the drought-affected areas that experienced yield reductions in Table 6. In the experiment, we focused on these states that were affected by drought and experienced yield reductions.
Figure 5: The difference in corn yield between each year from (a) 2019 to (e) 2023 and the previous year, with negative values indicating a reduction in yield.
Table 6: Summary of obviously yield-reduced states within drought-affected areas.
Year Area type Drought-affected states Obviously yield-reduced states
2019 IA,MI MI
2020 IA,ND,SD,NE,IL IA
2021 ND,SD,MN,NE,IA ND,SD
2022 NE,KS,SD,IA NE,KS,SD
2023 NE,KS,MN,IA,WI,MO IA,MO
4.3 KGML-SM performance comparison with traditional ML models
To validate the superiority of our KGML-SM model, we compared it with four traditional machine learning models: LR, multilayer perceptron (MLP), ridge regression (RR), and random forest (RF). LR (Bishop and Nasrabadi, 2006) is a simple statistical method that models the relationship between a dependent variable and one or more independent variables by fitting a linear equation to the data. MLP (Goodfellow et al., 2016) is a neural network with multiple layers, including an input, hidden, and output layer. It captures non-linear relationships and is widely used for classification and regression. RR (Hoerl and Kennard, 1970) is an extension of linear regression that includes an L2 regularization term to prevent overfitting by penalizing large coefficients. RF (Breiman, 2001) is an ensemble learning method that constructs multiple decision trees during training and aggregates their predictions to improve accuracy and robustness.
The results showed that our KGML-SM model consistently outperformed the other ML models across all years (Table 7). Based on the results, RF performed the best among traditional ML models, indicating its strong ability to capture complex relationships in the data. RR performed slightly worse than RF, with slightly higher RMSE values, suggesting that regularization helped improve predictions but was not as effective as ensemble learning. LR ranked next, showing higher RMSE values, likely due to its inability to model non-linear relationships effectively. MLP performed the worst, with the highest RMSE values in most years, indicating that it struggled to generalize well, possibly due to overfitting or insufficient training data.
Table 7: Comparison with traditional ML models
Year Method KGML-SM LR MLP RR RF
RMSE R2 RMSE R2 RMSE R2 RMSE R2 RMSE R2
2019 0.964 0.741 1.328 0.621 1.169 0.607 1.214 0.607 1.040 0.712
2020 0.980 0.792 1.304 0.690 1.207 0.734 1.230 0.661 1.120 0.719
2021 1.104 0.836 1.167 0.790 1.247 0.761 1.129 0.808 1.236 0.794
2022 1.085 0.837 1.471 0.740 1.400 0.765 1.318 0.781 1.185 0.821
2023 1.071 0.807 1.226 0.737 1.225 0.738 1.140 0.776 1.196 0.791
4.4 Ablation study of different compoments in KGML-SM model
To further demonstrate the contribution of each module in our KGML-SM model, we conducted a series of ablation studies. We began with a baseline model trained using only the GEE county-level dataset and the attention module (Att) (Wang et al., 2025). Next, we incorporated the simulated dataset to evaluate its impact on model performance (Att+sim). We then introduced the W2S encoder to assess its effectiveness in integrating soil moisture information (Att+sim+W2S). Finally, the impact of the drought-aware loss function was evaluated by comparing the performance of the KGML-SM model with the (Att+sim+W2S) model. We specifically compared the differences among models and their components in Table 8.
Table 8: Comparison of different methods and their components.
Method Attention module Simulated data W2S encoder Drought-aware loss
Att
Att+sim
Att+sim+W2S
KGML-SM
Through an ablation study of all model components (Table 9), we found that the simulated dataset contributed the most to performance improvement. This indicated that our APSIM field-level dataset effectively captured real-world data patterns, playing a crucial role in enabling the model to learn the relationship between corn yield and environmental variables. Additionally, our drought-aware loss function significantly enhanced performance, demonstrating that incorporating prior knowledge of the real-world relationship between drought and corn yield can effectively guide the training and prediction of ML models.
Table 9: Ablation study of different components in KGML-SM model
Year Method KGML-SM Att+sim+W2S Att+sim Att
RMSE R2 RMSE R2 RMSE R2 RMSE R2
2019 0.964 0.741 0.974 0.752 1.087 0.685 1.268 0.57
2020 0.98 0.792 0.981 0.783 1.003 0.77 1.011 0.77
2021 1.104 0.836 1.127 0.811 1.143 0.814 1.195 0.808
2022 1.085 0.837 1.119 0.802 1.244 0.802 1.315 0.779
2023 1.071 0.807 1.101 0.805 1.114 0.781 1.201 0.766
4.5 Spatial Visualization of model performance
In Fig. 6, we plotted the error maps for the KGML-SM and LR model results from 2019 to 2023. In 2019 (Fig. 6 (a)), the lower part of Michigan shows noticeable overestimation in the attention model, whereas our KGML-SM model performed well. In 2020 (Fig. 6 (b)), although the derecho storm in Iowa also had some impact on our KGML-SM model, we found that the LR model exhibited more severe overestimation. In 2021 (Fig. 6 (c)), overestimation in the upper-left region, particularly in North Dakota and South Dakota, is significantly reduced. In 2022 (Fig. 6 (d)), overestimation decreases notably in South Dakota and Kansas, though the effect in Kansas is less pronounced. In 2023 (Fig. 6 (e)), Missouri showed a clear reduction in overestimation, while the effect in Iowa is less noticeable. Overall, our model shows a notable improvement in addressing the issue of overestimation in drought-affected areas.
Figure 6: The error map for (1) KGML-SM and (2) LR model from (a) 2019 to (e) 2023.
4.6 Understanding model bias and prediction-observation alignment
We generated scatter plots for the KGML model and the LR model to analyze their prediction performance (Fig. 7). These plots help visualize the relationship between observed and predicted values, revealing patterns of overestimation, underestimation, and potential prediction biases across different years. In 2019 (Fig. 7 (a)), the KGML-SM model produced a noticeably narrower distribution, indicating a lower spread in prediction errors. In 2020 (Fig. 7 (b)), The predictions of the KGML-SM model were noticeably more concentrated along the diagonal and exhibited symmetry on both sides, whereas the LR model produced more dispersed predictions in high-yield regions. In 2021 (Fig. 7 (c)), the LR model exhibited prediction collapse, where certain observed values correspond to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations. In 2022 (Fig. 7 (d)), the KGML-SM model maintained a narrower and more concentrated prediction distribution, indicating improved robustness. In 2023 (Fig. 7 (e)), the KGML-SM model continued to avoid prediction collapse, demonstrating more stable performance compared to the attention model. This comparison highlighted the advantage of the KGML-SM model in mitigating prediction collapse and improving overall robustness across different years.
Figure 7: The scatter plot for (1) KGML-SM and (2) LR model from (a) 2019 to (e) 2023.
5 Discussion
Specifically, we discussed the role of soil moisture in model prediction from three perspectives:
  • 1.
    How did soil moisture influence model prediction spatially?
  • 2.
    How did soil moisture affect model performance throughout the corn growth season?
  • 3.
    how did soil moisture contribute to model prediction in drought and non-drought regions?
We leveraged the attention mechanism to weight input features, capturing each feature’s contribution to the model, to explore these questions.
5.1 Spatial influence of soil moisture on model prediction
We visualized the attention scores of soil moisture across twelve states in the U.S. Corn Belt from June to August over the years 2019 to 2023 (Fig. 8). The attention scores indicated the relative importance assigned to soil moisture by the model in different regions, with higher scores suggesting a stronger influence on yield prediction. The attention map highlighted how the model’s reliance on soil moisture varied across different growth stages and drought conditions, allowing us to assess whether soil moisture has a greater impact on the model in drought-affected areas.
Figure 8: The attention visualization of soil moisture on corn yield prediction in different regions. The corn growth period was segmented into three stages, covering (1) June to (3) August, for the years (a) 2019 to (e) 2023. The attention visualization was computed by first calculating the attention values for all counties. These values were then normalized based on the maximum value in this year to ensure a consistent color scale, making the distribution more visually interpretable.
In June, during the early growth period, the attention scores for soil moisture were generally lower across all four years, suggesting that soil moisture is less critical at this stage. This might be due to cooler temperatures and lower evaporation rates, which reduced the immediate need for water. In July, attention scores increased, particularly in drought-affected areas, as soil moisture became more influential during active growth and vegetative stages. By August, attention scores peaked in drought-affected states, aligning with the critical reproductive phase of corn when adequate soil moisture was essential for kernel development.
From 2019 to 2023, the attention distribution of soil moisture in corn yield prediction exhibited noticeable variations, aligning with the drought severity in different years. Note that, for better visualization, we normalized the attention values within each year. Therefore, we only need to focus on the differences across regions within the same year, rather than making cross-year comparisons. In early 2019 (Fig. 8(a1)), the attention was concentrated in Iowa, reflecting the drought-affected states. Subsequently (Fig. 8(a3)), the regions with higher attention values gradually expanded, but Iowa remained the area with the highest attention. In 2020 (Fig. 8(b)), the high-attention regions expanded to include North Dakota, South Dakota, and Nebraska, corresponding to widespread drought in USDA drought map. In 2021 (Fig. 8(c)), the attention intensified over North Dakota, South Dakota, Minnesota, and Nebraska, mirroring severe drought conditions. By 2022 (Fig. 8(d)), high-attention areas were mainly in Nebraska, Kansas, South Dakota, and Iowa. Additionally, over time, there was a trend of drought expanding toward the central region. In 2023 (Fig. 8(e)), the attention peaked across Nebraska, Kansas, Minnesota, Iowa, Wisconsin, and Missouri. Finally, in August, it reached its peak, and soil moisture became more important in most states in the central region.
This above analysis underscored the dynamic role of soil moisture in model prediction, with its importance intensifying during key growth stages and under severe drought conditions. The elevated attention scores in drought-prone regions suggested that soil moisture serves as a crucial predictor for accurately estimating corn yield in water-stressed environments.
5.2 Temporal role of soil moisture during the corn growth season
To illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—every 16 days from June to September. We found that the VIs had the highest influence on the model around August, which aligned with findings from previous studies (Johnson, 2014). In this period, VIs shows the strongest correlation with corn yield because this period corresponds to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield. High VIs in this timeframe indicate optimal chlorophyll content, canopy development, and water availability, which are critical for photosynthesis and grain formation. Additionally, we found that weather data showed a significant increase in attention around July in 2020 and 2022, and the periods with noticeable attention spikes aligned with the trends of VIs. This phenomenon was more pronounced in soil moisture, as soil moisture attention showed a strong correlation with VIs in all years except 2022. In 2022, a slight increase in soil moisture attention could still be observed near the VIs peak. This may be because VIs during these periods are closely related to certain weather data and soil moisture data. For example, NDWI increases with higher PPT because more rainfall enhances soil moisture and plant water content, GCI decreases with excessive Tmax because extreme heat can degrade chlorophyll and reduce plant vitality, etc. This indicated that our attention mechanism effectively captured feature importance over time dimension.
Figure 9: 5 year attention values of different feature types in the time series. The attention visualization is computed by first calculating the attention values for all features across all time points. Then, for each time point, we average the attention values across all features within the same category (VIs, Weather, and SM).
5.3 Statistical impact of soil moisture in drought and non-drought regions
The box plot illustrates the distribution of soil moisture attention across all counties from 2019 to 2023 (Fig. 10). Each point in this box plot represents a county, displaying the comparison of attention between red-marked drought regions and blue-marked non-drought regions from 2019 to 2023. Through this visualization, we aimed to analyze the differences in soil moisture attention between drought-affected and non-drought counties across all years. In 2021, the most severe drought year, both the variance and magnitude of soil moisture attention were the highest. We also observed that attention values were notably high in 2019, which aligns with our experimental results, where our KGML-SM model achieved the most significant improvement over other models. Since our model better captured the impact of soil moisture, it achieved the most significant improvement in years when soil moisture has a greater influence. In contrast, the values for 2020, 2022, and 2023 were more concentrated, with fewer extreme values. Overall, in all years, soil moisture attention in drought-affected regions was consistently higher than in non-drought regions, indicating that soil moisture had a greater impact on the model in these areas.
Figure 10: The boxplot of soil moisture attention in drought-affected areas (D--A) and non-drought areas (N--D) across all counties.
6 Conclusion
In this study, we proposed the KGML-SM model, which first supplements the limited soil moisture data with simulated data. It then learned the PB relationship between weather and soil moisture. Additionally, based on prior knowledge that drought leads to yield overestimation, we introduced a drought-aware loss function to mitigate this issue in drought-affected areas. We analyzed 12 states in the U.S. Corn Belt to examine the impact of soil moisture on corn yield prediction. Our method consistently outperformed baseline models across multiple test years. We also explore the spatial and temporal impact of soil moisture on the model by visualizing attention, highlighting when and where the model focuses more on soil moisture. In the future, we aim to apply transfer learning to adapt models trained on well-studied regions with simulated data to areas with limited data availability.

### v2

@@
-Knowledge-guided machine learning model with soil moisture for corn yield prediction under drought conditions
+Knowledge-guided machine learning for county-level corn yield prediction under drought
 Abstract
-Remote sensing (RS) techniques, by enabling non-contact acquisition of extensive ground observations, have become a valuable tool for corn yield prediction.
-Traditional process-based (PB) models are limited by fixed input features and struggle to incorporate large volumes of RS data.
-In contrast, machine learning (ML) models are often criticized for being “black boxes” with limited interpretability.
-To address these limitations, we used Knowledge-Guided Machine Learning (KGML), which combined the strengths of both approaches and fully used RS data.
-However, previous KGML methods overlooked the crucial role of soil moisture in plant growth.
-To bridge this gap, we proposed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, using soil moisture as an intermediate variable to emphasize its key role in plant development.
-Additionally, based on the prior knowledge that the model may overestimate under drought conditions, we designed a drought-aware loss function that penalizes predicted yield in drought-affected areas.
-Our experiments showed that the KGML-SM model outperformed other ML models.
-Finally, we explored the relationships between drought, soil moisture, and corn yield prediction, assessing the importance of various features and analyzing how soil moisture impacts corn yield predictions across different regions and time periods.
+Remote sensing (RS) technique, enabling the non-contact acquisition of extensive ground observations, is a valuable tool for crop yield predictions.
+Traditional process-based models struggle to incorporate large volumes of RS data, and most users lack understanding of crop growth mechanisms.
+In contrast, machine learning (ML) models are often criticized as “black boxes” due to their limited interpretability.
+To address these limitations, we utilized Knowledge-Guided Machine Learning (KGML), a framework that leverages the strengths of both process-based and ML models.
+Existing works have either overlooked the role of soil moisture in corn growth or did not embed this effect into their models.
+To bridge this gap, we developed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, treating soil moisture as an intermediate variable in corn growth to emphasize its key role in plant development.
+Additionally, based on the prior knowledge that the model may overestimate under drought conditions, we designed a drought-aware loss function that penalized predicted yield in drought-affected areas.
+Our experiments showed that the KGML-SM model outperformed other traditional ML models.
+We explored the relationships between drought, soil moisture, and corn yield prediction by assessing the importance of different features within the model, and analyzing how soil moisture impacts predictions across different regions and time periods.
+Finally we provided interpretability for prediction errors to guide future model optimization.
 keywords:
-Corn yield prediction; Knowledge-guided Machine learning; Soil moisture; drought
+Corn yield prediction; Knowledge-guided Machine learning; Soil moisture; drought-aware learning
 journal: Journal
@@
 […]
-[inst2]organization=Department of of Soil Science,addressline=University of Wisconsin-Madison, city=Madison, postcode=53706, state=WI, country=USA
+[inst2]organization=Department of of Soil and Environmental Sciences,addressline=University of Wisconsin-Madison, city=Madison, postcode=53706, state=WI, country=USA
 […]
@@
 1 Introduction
-Corn plays a vital place in the U.S.
-as a primary crop, supporting food security as well as the animal feed and biofuel industries (Graham et al., 2007; Thompson, 1969).
+Corn, as a primary crop, plays a vital role in the U.S.
+agriculture, supporting food security as well as the animal feed and biofuel industries (Graham et al., 2007; Thompson, 1969).
 Accurate yield prediction is crucial for effective resource management and economic stability (Kucharik and Ramankutty, 2005).
-However, achieving precise prediction across large areas and diverse environmental conditions is challenging, making it necessary to pursue advanced research in corn yield prediction techniques (Weiss et al., 2020).
-Traditional process-based (PB) models (Weiss et al., 2007) are built on a deep understanding of the physical, chemical, and biological processes within crop growth systems (Weiss et al., 2004).
+Achieving accurate prediction across large areas and diverse environmental conditions is challenging (Weiss et al., 2020).
+Traditional process-based models (Weiss et al., 2007) are built on a deep understanding of the physical, chemical, and biological processes within crop growth systems (Weiss et al., 2004).
 These models use equations derived from scientific principles to simulate the behavior of crops, weather, and soil interactions (Weiss et al., 2002), making them highly interpretable and reliable in scenarios where the underlying processes are well understood (Puntel et al., 2016; Shahhosseini et al., 2021).
-Several promising modeling frameworks are designed for simulating and predicting agricultural production systems, such as The Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), WOrld FOod STudies (WOFOST) (Van Diepen et al., 1989), and The Agricultural Policy / Environmental eXtender (APEX) (Williams and Izaurralde, 2010).
-Many recent studies (Zhen et al., 2023, 2022) have successfully used these models to predict crop yields.
-However, PB models have several limitations.
+Several promising modeling frameworks are designed for simulating and predicting agricultural production systems, such as the Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), the Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), the WOrld FOod STudies (WOFOST) (Van Diepen et al., 1989), and the Agricultural Policy / Environmental eXtender (APEX) (Williams and Izaurralde, 2010).
+Many studies (Zhen et al., 2023, 2022) have successfully used these models to predict crop yields.
+However, process-based models have several limitations.
 Plant growth is highly complex, and these models rely on a limited set of fixed input features, which are insufficient to accurately capture the intricacies of this process.
-Moreover, PB models require extensive manual parameter tuning, limiting their applicability for large-scale applications across vast regions.
+Moreover, process-based models require extensive manual parameter tuning, limiting their applicability for applications across vast regions.
 In recent years, remote sensing (RS) has provided several benefits for corn yield prediction (Lobell et al., 2015).
-It allows for the collection of large-scale and real-time data across vast agricultural areas, which, in conjunction with machine learning (ML) methods, provides an accurate and efficient way for large-scale yield prediction.
-Deep learning (Goodfellow et al., 2016), a widely used ML method, has the advantage of leveraging vast amounts of RS data by automatically learning features from large datasets.
-This removes the need for manual feature engineering, enabling the model to capture complex patterns and relationships between environmental factors and crop yield (He et al., 2016).
-Many studies have applied deep learning to crop yield prediction and shown its advantages.
+It enables the collection of large-scale real-time data across vast agricultural areas.
+This data, typically combined with machine learning (ML), facilitates accurate and efficient large-scale yield prediction.
+Deep learning (DL) (Goodfellow et al., 2016), a widely used ML method, has the advantage of leveraging a large amount of RS data by automatically learning features from big datasets.
+This removes the need for manual feature engineering, enabling the model to capture complex patterns and relationships between crop environmental factors and crop yield (He et al., 2016).
+Many studies applying DL to crop yield prediction have shown its advantages in performance improvement.
 You et al.
@@
 (2021a) developed a county-level corn yield prediction model using a Bayesian Neural Network across 12 states in the Corn Belt, showing the potential for large-scale corn yield prediction tasks.
-However, despite the popularity of deep learning models in corn yield prediction tasks, they have certain limitations.
-First, deep learning models require a large amount of data to learn the relationships between inputs and outputs.
-Second, deep learning models often function as a black box, lacking reliable interpretability of the prediction results.
-Therefore, finding ways to expand our data sources is a critical factor in achieving accurate predictions.
-Additionally, there are complex biological processes among soil, weather, and corn, which are challenging for deep learning models to fully capture.
-Given these challenges, combining PB models with deep learning models has become an important research topic.
-Knowledge-guided machine learning (KGML) (Karpatne et al., 2022) aims to integrate scientific knowledge in ML frameworks to achieve better generalizability, scientific consistency, and explainability of results.
+However, despite the popularity of DL models in corn yield prediction tasks, they have certain limitations.
+First, DL models require a large amount of data to learn the relationships between inputs and outputs.
+Additionally, DL models often function as a black box, lacking reliable interpretability of the prediction results.
+Specifically, crop growth involves complex biological processes among soil, weather, and corn, which are challenging for DL models to fully capture.
+Therefore, how to leverage advantages of process-based models and DL models to address these challenges has become an important research topic.
+Knowledge-guided machine learning (KGML) (Karpatne et al., 2022) aims to integrate scientific knowledge in ML frameworks to achieve better performance, scientific consistency, and explainability of results.
 KGML involves three main approaches: Knowledge-Guided Learning, Knowledge-Guided Architecture, and Knowledge-Guided Pretraining.
@@
 By initializing model parameters using data that encodes scientific principles, this method reduces the need for large amounts of observational data and improves performance in data-scarce scenarios (Licheng et al., 2022; Chen et al., 2023).
-Building on these approaches, KGML has been increasingly applied in various domains, including agriculture, hydrology, and climate science, enabling improved data-driven decision-making and predictive modeling.
-Recently, many studies have used KGML to combine PB models and ML models for crop yield prediction (Kimball et al., 2023; Burroughs et al., 2023).
-Since PB models already provide insights into the crop growth process and generate high-quality simulated data, Knowledge-Guided Pretraining is the most popular approach (He et al., 2023; Yang et al., 2023).
-These studies aim to use ML to identify patterns between crop features and yield from large datasets, while also deriving interpretability from the data simulated by PB models.
-Specifically, KGML models leverage PB models to simulate intermediate variables representing key plant growth processes, enhancing their ability to capture domain-specific knowledge.
-Then, ML models are applied to simulate specific parts of the process.
-By incorporating these intermediate variables, the model can better represent the underlying biological and environmental mechanisms.
+Building on these approaches, KGML has been increasingly used in various domains, such as agriculture, hydrology, and biology, enabling improved data-driven decision-making and predictive modeling.
+Recently, many studies have used KGML to combine process-based and ML models for crop yield predictions (Kimball et al., 2023; Burroughs et al., 2023).
+Since process-based models already provide insights into the crop growth process and generate high-quality simulated data, Knowledge-Guided Pretraining is the most popular approach (He et al., 2023; Yang et al., 2023).
+These studies aim to identify patterns between agricultural features and crop yields from large datasets using ML, while leveraging data simulated by process-based models to provide interpretability.
+Specifically, intermediate variables in crop growth generated by process-based models are used to train ML models, allowing the ML models to simulate specific parts of the crop growth process.
 For example, He et al.
-(2023) integrated physical knowledge from PB models and enabled the model to adjust for temporal data shifts over the years.
+(2023) integrated physical knowledge from process-based models and enabled the model to adjust for temporal data shifts over the years.
 Yang et al.
@@
 (2024) presented a knowledge-guided computer vision framework designed to monitor and simulate the growth of individual strawberry fruits, which can predict growth at the individual fruit level by integrating image-based data and dynamic simulations.
-Although these KGML models have made significant progress in crop yield prediction, all existing KGML-based studies on crop yield prediction have overlooked the critical role of soil, particularly how soil moisture contributes to plant growth and affects yield.
+Although these KGML models have made significant progress in crop yield prediction, all existing KGML-based studies on crop yield predictions have overlooked the critical role of soil, particularly how soil moisture contributes to corn growth and affects yield.
 Soil moisture is influenced by temperature, precipitation, and other environmental factors and directly affects crop growth.
-Unlike RS data, soil moisture has a cumulative impact on yield over time.
-As a key drought indicator, low soil moisture causes water stress, which limits corn growth and reduces yields (Unganai and Kogan, 1998).
-Nonetheless, both PB and ML models have certain limitations in modeling soil moisture.
-In PB models, weather data serve as input, while soil moisture is treated as output (McCown et al., 1996; Williams and Izaurralde, 2010), failing to capture its direct impact on crop yield.
-In ML models, a key challenge is the limited soil moisture data.
-For example, as the most commonly used soil moisture dataset, the SPL4SMGP.007 SMAP dataset (Entekhabi et al., 2010) only covers data from 2015 onward, limiting ML models’ ability to learn from extensive historical data and impacting the model performance.
-Therefore, we integrated PB and ML models through KGML to better model soil moisture for corn yield prediction.
-Additionally, we examined how soil moisture variations over time and across space affected corn yield in drought-prone areas, where drought often reduced yields and hindered prediction accuracy (Cooper et al., 2014).
+As a key indicator of crop water supply, low soil moisture causes water stress, which limits corn growth and reduces yields (Unganai and Kogan, 1998).
+Unlike most RS data, such as NDVI, which only reflects the growth status of plants at a specific moment, soil moisture has a cumulative impact on yield over time.
+Nonetheless, a significant challenge for ML models is the limited availability of soil moisture data.
+For example, the widely used SPL4SMGP.007 SMAP soil moisture dataset (Reichle et al., 2022) only covers the period from 2015 onward, which limits the ability of ML models to learn from extensive historical data and impacts model performance.
+Therefore, we present a Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework that explicitly incorporated the soil moisture impact in the ML model and provided interpretability.
+The model consisted of two major components: a Weather-to-Soil (W2S) encoder modeling the relationship between weather inputs and soil moisture outputs, and an attention module (Vaswani, 2017) integrating features from multiple sources and dynamically weighting their influences on corn yield across time and space.
+To train this model, we adopted a two-step process.
+First, we used the APSIM model to generate a simulated field-level dataset for pretraining.
+Then, the model was finetuned using a county-level dataset from Google Earth Engine (GEE) (Gorelick et al., 2017) and United States Department of Agriculture (USDA) National Agricultural Statistics Service (NASS).
+Furthermore, we conducted research into the relationship between soil moisture and corn yield prediction in drought-affected areas, where drought often reduced yields and caused overestimation (Cooper et al., 2014).
 Researchers have long been interested in studying the relationship between soil moisture, drought, and corn yield prediction (Ines et al., 2013; Bushong et al., 2016; Mladenova et al., 2017).
-However, these studies either have a narrow scope (Vergopolan et al., 2021), resulting in limited data that cannot be effectively generalized to broader applications, or they focus solely on end-of-season yield prediction (Pignotti et al., 2023).
-In this paper, we proposed the KGML-SM framework, which integrated PB and ML models for corn yield prediction while incorporating the role of soil moisture in plant growth.
-The framework consisted of two key components: a Weather-to-Soil (W2S) encoder that modeled the relationship between weather inputs and soil moisture outputs, and an attention module (Vaswani, 2017) that integrated features from multiple sources and dynamically weighted their influence on corn yield across different growth stages and drought conditions.
-Additionally, we designed a drought-aware yield prediction loss function, informed by the observed relationship between drought and yield reduction, to further enhance model performance under drought conditions.
-To train this model, we adopted a two-step process.
-First, we used the APSIM model to generate simulated soil moisture and corn yield data, which were then combined with the input weather data for pretraining.
-Then, we finetuned the model with soil moisture data from the SMAP product, corn yield data from official statistics, as well as weather data and vegetation indices (VIs) from Google Earth Engine (GEE) (Gorelick et al., 2017), which enabled it to adapt to real-world variability and enhance predictive accuracy.
-Our research aimed to improve the accuracy and interpretability of corn yield predictions, which ultimately contributed to more resilient agricultural systems in the face of climate variability.
+However, these studies either had a narrow scope, or did not utilize this connection for model optimization.
+we designed a drought-aware loss function to further enhance model performance under drought conditions and mitigate overestimation problem.
+Finally, based on the relationship between soil moisture and corn yield prediction, we provided explanations for the prediction errors of the KGML-SM model and offered directions for future model optimization.
+Our research aimed to improve the accuracy and interpretability of corn yield predictions, ultimately contributing to more resilient agricultural systems under extreme weather variability.
+Our main contributions are as follows:
+  • 1.
+    This paper introduced the KGML-SM framework. This framework integrated process-based and ML models for corn yield prediction while explicitly incorporating the influence of soil moisture.
+  • 2.
+    A drought-aware loss function was designed. This function enhanced model performance under drought conditions and mitigated overestimation.
+  • 3.
+    Based on the relationship between drought, soil moisture, and corn yield prediction, we provided interpretability for the prediction errors of the KGML-SM model and offered directions for future model optimization.
 2 Data acquisition
 In this study, we developed two datasets for corn yield prediction.
-The first, a field-level dataset, was generated using APSIM (McCown et al., 1996) and used for pretraining.
-The second, a county-level dataset, was derived from GEE (Gorelick et al., 2017) and USDA NASS and employed for finetuning.
+The first dataset was a field-level dataset generated using APSIM (McCown et al., 1996) and used for pretraining.
+The second dataset was a county-level dataset derived from GEE (Gorelick et al., 2017) and USDA NASS and used for finetuning.
 The workflow was divided into three main steps: simulation, pretraining, and finetuning.
@@
 PRISM c Gridded weather data (Daly et al., 2015, 2008)
-MODIS c Vegetation indices (Justice et al., 1998; Schaaf and Wang, 2015)
+MODIS MCD43A4 c Vegetation indices (Schaaf and Wang, 2015)
 MODIS MCD18A1.061 c Radiation data (Wang, 2021)
 USDA-CDL c Crop data layer (USDA-NASS, 2017)
-SMAP c Soil moisture data (Entekhabi et al., 2010)
+SMAP c Soil moisture data (Reichle et al., 2022)
 USDA NASS yield c County-level yield (USDA, 2020)
@@
 Our research focused on corn yield prediction across the U.S.
-Corn Belt, selecting twelve states as our study area: North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan.
-These states are crucial agricultural regions in the U.S., known for their significant contributions to corn production.
+Corn Belt.
+Twelve states are selected as our study area, including North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan.
+These states are crucial agricultural states in the U.S., known for their significant contributions to corn production.
 We generated a five-year average yield map (Fig.
@@
 In this section, we explain how APSIM was used to generate a field-level simulated dataset for model pretraining.
-For the simulation of the APSIM field-level dataset, we used data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides U.S.
+We used data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides U.S.
 agricultural and environmental data.
-We downloaded all available station data from the IEM, which included data from thousands of stations across 12 states from 1980 to 2023 (Table 2).
+We extracted all available station data from the IEM, which included data from thousands of stations across 12 states from 1980 to 2023 (Table 2).
 Table 2: The number of stations selected for each state.
-State IA IL IN KS MI MN MO ND NE OH SD WI
-Station number 113 120 89 150 140 140 130 79 128 100 114 147
-2.2.1 Input and output of APSIM
-The APSIM model simulates crop yield using four weather data inputs: maximum temperature (Tmax), minimum temperature (Tmin), precipitation (PPT), and radiation (Radn).
-It also requires certain management parameters, which we describe in detail later (Sec.
+State IL IN IA KS MI MN MO NE ND OH SD WI
+Station number 120 89 113 150 140 140 130 128 79 100 114 147
+2.2.1 Input and output of APSIM simulation
+The APSIM model simulates corn yield based on four weather data inputs: maximum and minimum temperature (Tmax and Tmin), precipitation (PPT), and radiation (Radn).
+The simulation also requires the station’s latitude, longitude, and year.
+Some management parameters are also necessary, including start and end of sowing window, plant population, fertilizer amount, and initial soil water, etc (Sec.
 3.3.1).
-Maximum and minimum temperatures influence the rate of plant development and growth stages, impacting processes like photosynthesis and respiration.
-Precipitation data are essential for modeling soil moisture levels, which directly affect water availability for crops.
-Radiation is a critical factor in photosynthesis, as it provides the energy needed for plant growth.
-These weather data are available from the IEM website (Herzmann et al., 2004), which collects and distributes environmental data from various observing networks.
+Tmax and Tmin influence the rate of plant development, and affect processes like photosynthesis and respiration.
+PPT are essential for modeling soil moisture levels, which directly affect water availability for crops.
+Radn is a critical factor in photosynthesis, as it provides the energy needed for plant growth.
+These weather data of IEM stations are available from the IEM website (Herzmann et al., 2004).
 The APSIM model takes weather data as input to simulate root zone soil moisture (SM_rootzone), surface soil moisture (SM_surface), and corn yield.
@@
 SM_surface plays a key role in seed germination and early growth stages.
-2.2.2 Field-level dataset for pretraining
-Besides APSIM’s input and output variables, we also extracted two useful features from IEM: prediction year and location (longitude and latitude).
-Then, we calculated the five-year historical average yield based on the average local simulated corn yield over the past five years.
-We believed these three features might contribute to improving the accuracy of our corn yield prediction model.
-After APSIM generated the simulation data, we combined its input weather data, output soil moisture, and corn yield, along with other features to construct the APSIM field-level dataset (Table 3).
-This dataset was used for pretraining our KGML model, aiming to help the model learn the internal processes of corn growth.
-Table 3: Summary of variables comprising the APSIM field-level dataset, including inputs, outputs, and other features
+2.2.2 Variables summary in field-level dataset for pretraining
+We calculated a five-year historical average yield based on the simulated corn yield to establish a local baseline for model accuracy improvement.
+We also included the prediction year and location (longitude and latitude) of IEM stations in the field-level dataset.
+Because soil properties and topography vary across space and time, year and location can help the ML model capture this variation.
+Then we combined the weather data, simulated soil moisture and corn yield, along with other variables to construct the APSIM field-level dataset (Table 3).
+Table 3: Summary of variables comprising the APSIM field-level dataset, including inputs and simulated outputs
 Category Variables
@@
 Corn yield Corn yield (t/ha) (simulated)
-Others Prediction year
-Location (Latitude and longitude)
-Historical average corn yield (t/ha) (simulated)
+Others Prediction year (input)
+Location (Latitude and longitude) (input)
+Historical average yield (t/ha) (simulated)
 2.3 GEE county-level dataset
-In this section, we introduce how to construct a county-level dataset for finetuning from GEE.
-All the variables in the APSIM field-level dataset were also included.
-To enhance the ML model with RS data, we included VIs from satellite imagery: Green Chlorophyll Index (GCI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI).
+In this section, we introduce how to construct a county-level dataset for finetuning.
+Details of the studied counties are provided in Table 4.
+This dataset includes all the variable types in the APSIM field-level dataset (Table 3), including weather, soil moisture, and other features, but from different sources.
+To enhance ML model performance with RS data, we also included four Vegetation Indices (VIs) from satellite imagery: Green Chlorophyll Index (GCI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI).
+Table 4: The average number of counties selected for each state in GEE county-level dataset.
+State IL IN IA KS MI MN MO NE ND OH SD WI
+County number 91 74 87 74 57 64 74 69 48 72 39 62
 2.3.1 Vegetation indices
-The Moderate Resolution Imaging Spectroradiometer (MODIS) dataset (Justice et al., 1998; Schaaf and Wang, 2015) is a vital source of satellite-derived data with 500 m resolution, providing consistent, high-quality observations of the Earth’s surface.
-MODIS captures data in multiple spectral bands and offers various products like VIs.
+The Terra and Aqua combined Moderate Resolution Imaging Spectroradiometer (MODIS) Land Cover Climate Modeling Grid Version 6 data product (MCD12Q1 v6) (Schaaf and Wang, 2015) is a vital source of satellite-derived data with 500 m resolution, providing consistent, high-quality observations of the Earth’s surface.
+MODIS MCD12Q1 v6 captures data in multiple spectral bands and offers various products like VIs.
 These indices are essential for monitoring vegetation health, biomass, water content, and chlorophyll levels across different regions and time scales.
 They help improve yield prediction by indicating vegetation growth and crop health.
-The Green Chlorophyll Index (GCI) is used to estimate the chlorophyll content in plants, which directly relates to plant health and productivity.
+The Green Chlorophyll Index (GCI) (Gitelson et al., 2005) is used to estimate the chlorophyll content in plants, which directly relates to plant health and productivity.
 It is calculated using the near-infrared (NIR) and green spectral bands.
@@
 where $NIR$ represents the surface reflectance in the Near Infrared band, and $Green$ represents the surface reflectance in the Green band.
-The Enhanced Vegetation Index (EVI) is designed to optimize the vegetation signal by enhancing sensitivity in high biomass areas and reducing atmospheric and canopy background noise.
+The Enhanced Vegetation Index (EVI) (Huete et al., 2002) is designed to optimize the vegetation signal by enhancing sensitivity in high biomass areas and reducing atmospheric and canopy background noise.
 It uses the red, blue, and NIR spectral bands.
@@
 where $Red$ and $Blue$ represent the surface reflectances of the Red band and Blue band, respectively.
-The Normalized Difference Water Index (NDWI) measures the moisture content in vegetation and soil by using the NIR and shortwave infrared (SWIR) bands.
+The Normalized Difference Water Index (NDWI) (Gao, 1996) measures the moisture content in vegetation and soil by using the NIR and shortwave infrared (SWIR) bands.
 NDWI is crucial for monitoring water stress in plants, assessing drought conditions, and detecting changes in water bodies.
@@
 where $SWIR$ represents the shortwave infrared band.
-The Normalized Difference Vegetation Index (NDVI) is one of the most widely used indices for measuring vegetation greenness and biomass.
+The Normalized Difference Vegetation Index (NDVI) (KRIEGLER, 1969) is one of the most widely used indices for measuring vegetation greenness and biomass.
 It is calculated using the NIR and red spectral bands.
@@
 This dataset is widely used in agricultural research, hydrology, and weather studies due to its fine spatial resolution and comprehensive coverage, making it an essential tool for understanding and predicting weather-related impacts on crop yield and other environmental processes.
-The MCD18A1 Version 6.1 (Wang, 2021) is a MODIS Terra and Aqua combined Downward Shortwave Radiation gridded Level 3 product that provides us with reliable radiation data (Radn).
-It is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours.
+The MCD18A1 Version 6.1 (Wang, 2021) is a MODIS Terra and Aqua combined Downward Shortwave Radiation gridded Level 3 product.
+The reliable radiation data (Radn) is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours.
 Downward Shortwave Radiation is incident solar radiation over land surfaces in the shortwave spectrum (300-4,000 nanometers) and is an important variable in land-surface models that address a variety of scientific and applied issues.
 2.3.3 Soil moisture
-The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Entekhabi et al., 2010) plays a critical role in our research on the relationship between drought and corn yield prediction.
-By providing detailed measurements of soil moisture at both the surface (0–5 cm) and root zone levels (0–100 cm), SMAP data allows us to assess the availability of water in the soil, a key factor influencing crop growth and resilience during drought conditions.
-2.3.4 County-level dataset for finetuning
-In addition to the primary weather and vegetation parameters, we also incorporated features such as the prediction year, location (latitude and longitude), and the 5-year historical average yield (USDA, 2020) into our GEE county-level dataset, as in the APSIM field-level dataset.
-By integrating these features, the model gained a more comprehensive understanding of the environmental and temporal factors affecting corn yield, leading to more reliable predictions.
+The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Reichle et al., 2022) plays a critical role in our research on the relationship between drought and corn yield prediction.
+By providing detailed measurements of soil moisture at both the surface (0–5 cm) and root zone levels (0–100 cm), SMAP level 4 data allows us to assess the availability of water in the soil, a key factor influencing crop growth and resilience during drought conditions.
+2.3.4 Variables summary in county-level dataset for finetuning
+Additional features included the prediction year, location (latitude and longitude), and the 5-year historical average yield (USDA, 2020).
+These features were also added to the GEE county-level dataset, consistent with their inclusion in the APSIM field-level dataset.
 All the variables in GEE county-level dataset are listed in Table 5.
@@
 3 Methodology
-The overall pipeline of our KGML-SM is shown in Fig.
+The overall pipeline of the KGML-SM is shown in Fig.
 2.
-Our deep learning model consisted of two main components: the W2S encoder, which learned the relationship between weather data and soil moisture, and the attention-based feature-weighting module, which captured how various features influence corn yield.
-We first pretrained the model using the APSIM field-level dataset, then finetuned it with the GEE county-level dataset.
+This architecture comprised two principal components: the W2S encoder, designed to capture the relationship between weather data and soil moisture, and the attention-based feature-weighting module, which learned how various features influence corn yield.
+Initially, the model was pretrained on the APSIM field-level dataset, followed by finetuning using the GEE county-level dataset.
 Figure 2: The pipeline of the proposed KGML-SM model.
@@
 3.1), followed by an introduction to the model components (Sec.
-3.2); next, we explain how field-level dataset is generated for pretraining (Sec.
-3.3) and county-level dataset for finetuning and testing (Sec.
-3.4); finally, we introduce how the KGML model is trained and used for prediction (Sec.
+3.2); next, we explain how the APSIM field-level dataset is generated for pretraining (Sec.
+3.3) and the GEE county-level dataset for finetuning and testing (Sec.
+3.4); finally, we introduce how the KGML-SM model is trained and used for prediction (Sec.
 3.5).
 3.1 Problem formulation
-We formally define the corn yield prediction problem and present its mathematical formulation within a KGML framework.
-For a unique combination of county and year sample $i$, where $i\in\mathcal{N}$, $\mathcal{N}$ represents all counties across all years in the experiment, we define the input features as follows.
-The weather features are denoted as $\mathbf{w_{i}}=[w^{1}_{i},w^{2}_{i},\dots,w^{t}_{i}]$, and the other features (prediction year, location, and historical average yield) as $\mathbf{o_{i}}=[o^{1}_{i},o^{2}_{i},\dots,o^{t}_{i}]$, where $t\in\mathcal{T}$ and $\mathcal{T}$ represents the entire period from corn planting to harvest.
-We also have the simulated soil moisture as $\mathbf{s_{i}}=[s^{1}_{i},s^{2}_{i},\dots,s^{t}_{i}]$ and $y_{i}$ denotes the corn yield of each county in that year as supervision.
-The APSIM field-level dataset, $\mathcal{D}_{\text{sim}}=\{(\mathbf{w_{i}},\mathbf{o_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$, is used for pretraining.
-The vertical bar $\mid$ separates the input variables from the target labels.
-Additionally, the VIs are represented as $\mathbf{v_{i}}=[v^{1}_{i},v^{2}_{i},\dots,v^{t}_{i}]$.
-We denote two subsets of the GEE county-level dataset as $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ respectively, where $\mathcal{D}_{\text{train}}$ is used for finetuning and $\mathcal{D}_{\text{test}}$ for testing, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}=\{(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$.
-Here, we omit the description of the validation set $\mathcal{D}_{\text{val}}$ but introduce its partitioning and usage details in the experimental setup (Sec.
-4.1).
-The objective is to first learn an encoder $f_{\text{W2S}}$ that maps weather inputs to soil moisture: $\mathbf{\hat{s}_{i}}=f_{\text{W2S}}(\mathbf{w_{i}})$.
-Then, the predicted soil moisture $\mathbf{\hat{s}_{i}}$ is used along with other input features to predict yield through an attention module $f_{\text{att}}$: $y_{i}=f_{\text{att}}(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}},\mathbf{\hat{s}_{i}})$.
+The corn yield prediction problem is formally defined in this section, along with its mathematical formulation within the KGML framework.
+Let each unique county-year combination be represented by a sample indexed as $i$ ($i=1,2,\dots,N$), where $N$ is the total number of county-year instances in the experiment.
+For each sample $i$, the input features are specified as follows: The temporal weather features are denoted by the vector $\mathbf{w_{i}}=[w^{1}_{i},w^{2}_{i},\dots,w^{T}_{i}]$.
+Other features, including the prediction year, geographical location, and historical average yield, are aggregated into the vector $\mathbf{o_{i}}=[o^{1}_{i},o^{2}_{i},\dots,o^{T}_{i}]$, In these representations, $T$ corresponds to the total number of discrete time steps covering the duration from corn planting to harvest.
+Additionally, simulated soil moisture data is represented as $\mathbf{s_{i}}=[s^{1}_{i},s^{2}_{i},\dots,s^{T}_{i}]$.
+The corn yield record for sample $i$ is denoted by $y_{i}$ and serves as the supervision for model training.
+The APSIM-generated field-level dataset, specified as $\mathcal{D}_{\text{field/pretrain}}=\{(\mathbf{w_{i}},\mathbf{o_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$, is employed for the pretraining phase.
+In this notation, the vertical bar $\mid$ separates the input variables ($\mathbf{w_{i}},\mathbf{o_{i}}$) from the target labels ($\mathbf{s_{i}},\mathbf{y_{i}}$).
+Furthermore, VIs represented by the vector $\mathbf{v_{i}}=[v^{1}_{i},v^{2}_{i},\dots,v^{T}_{i}]$ are incorporated.
+The county-level dataset derived from GEE is denoted as $\mathcal{D}_{\text{county/finetune}}=\{(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}}\mid\mathbf{s_{i}},\mathbf{y_{i}})\}$ and is utilized for the finetuning.
+The objective is to first build the W2S encoder $f_{\text{W2S}}$ to map weather inputs to soil moisture $\mathbf{\hat{s}_{i}}=f_{\text{W2S}}(\mathbf{w_{i}})$.
+Then, the predicted soil moisture $\mathbf{\hat{s}_{i}}$ is combined with other input features to predict yield via an attention module $f_{\text{att}}$, resulting in the prediction $\hat{y_{i}}=f_{\text{att}}(\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}},\mathbf{\hat{s}_{i}})$.
 The model’s performance is evaluated by comparing the predicted yields $\hat{y}_{i}$ with the actual yields $y_{i}$.
@@
 3.2.1 Weather-to-Soil encoder
-W2S encoder is a module designed to model the influence from weather conditions to soil moisture.
-By capturing the relationship between weather and soil moisture, the W2S encoder enhances the model’s understanding of soil dynamics and the impact of weather variability on soil conditions and corn yield.
-The W2S encoder follows a U-Net-based encoder-decoder structure (Ronneberger et al., 2015), consisting of an encoding function $\mathcal{E}$, a decoding function $\mathcal{D}$, and a final fully connected layer $\mathcal{F}$ for feature transformation.
-Given an input time-series weather input $\mathbf{w_{i}}$, the encoding process can be formulated as:
-$\mathbf{z}_{1}=\mathcal{E}_{1}(\mathbf{w_{i}}),\quad\mathbf{z}_{2}=\mathcal{E}_{2}(\mathcal{P}(\mathbf{z}_{1}))$ (5)
-where $\mathcal{E}_{i}$ denotes the $i$-th convolutional block (LeCun et al., 1995), and $\mathcal{P}$ represents the max-pooling operation (Nagi et al., 2011) that reduces the temporal dimension.
-For simplicity we omit the subscript $i$ for $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$.
-The decoding process reconstructs high-level features through upsampling and convolutional layers:
-$\mathbf{\tilde{z}}_{2}=\mathcal{U}(\mathbf{z}_{2}),\quad\mathbf{\tilde{z}}_{1}=\mathcal{D}_{1}([\mathbf{\tilde{z}}_{2},\mathbf{z}_{1}])$ (6)
-where $\mathcal{U}$ is a transposed convolution (upsampling) operation, $[\mathbf{\hat{z}}_{2},\mathbf{z}_{1}]$ represents skip connections concatenating encoder and decoder feature maps, and $\mathcal{D}_{1}$ is the decoder convolutional block.
-The predicted soil moisture $\mathbf{\hat{s}_{i}}$ is obtained via:
-$\mathbf{\hat{s}_{i}}=\mathcal{F}(\mathcal{D}_{1}([\mathbf{\tilde{z}}_{2},\mathbf{z}_{1}]))$ (7)
+The W2S encoder is a module designed to model the influence of weather conditions on soil moisture.
+By capturing the relationship between weather and soil moisture, the W2S encoder improves the model’s understanding of soil dynamics and the impact of weather variability on soil conditions and corn yield.
+The W2S encoder employs a U-Net-based encoder-decoder architecture, which consists of an encoder, a decoder, and a fully connected layer (Goodfellow et al., 2016) for feature transformation.
+Given a time-series weather input $\mathbf{w_{i}}$, the encoder extracts hierarchical representations by progressively downsampling the temporal dimension.
+The decoder then reconstructs high-level features using upsampling and skip connections that integrate information from the encoder.
+Finally, a fully connected layer transforms the decoded features into the predicted soil moisture output $\mathbf{\hat{s}_{i}}=f_{\text{W2S}}(\mathbf{w_{i}})$.
 3.2.2 Attention module
 The attention mechanism (Vaswani, 2017) is a powerful tool in ML that enables models to focus on the most relevant parts of the input data when making predictions.
-By assigning different levels of importance, or weights, to various input elements, the attention mechanism helps the model prioritize the most crucial information for the task.
+By assigning different levels of importance to various input elements, the attention mechanism helps the model prioritize the most crucial information.
 In KGML-SM, we aim to use the attention mechanism to weight different features, helping us understand each feature’s contribution to yield prediction across different dimensions.
-Specifically, suppose that we have input $x_{i}=[\mathbf{w_{i}},\mathbf{o_{i}},\mathbf{v_{i}},\mathbf{\hat{s}_{i}}]$ for predicting corn yield.
-For each feature embedding $x_{i}$, we use the attention mechanism to learn the corresponding weight $\alpha_{i}$, which is then used to compute the final yield $y_{i}$.
-First, we compute the query $Q_{i}$, key $K_{i}$, and value $V_{i}$ vectors from the feature $x_{i}$ using learned linear transformations:
-$Q_{i}=W_{Q}x_{i},\quad K_{i}=W_{K}x_{i},\quad V_{i}=W_{V}x_{i}$ (8)
-where $W_{Q}$, $W_{K}$, and $W_{V}$ are the learned weight matrices for the query, key, and value, respectively.
+The input for corn yield prediction is formulated as $\mathbf{X_{i}}=[\mathbf{w_{i}};\mathbf{o_{i}};\mathbf{v_{i}};\mathbf{\hat{s}_{i}}]$ representing the concatenation of feature vectors.
+For each feature embedding $\mathbf{X_{i}}$, an attention mechanism is employed to learn the corresponding attention weight $\mathbf{\alpha_{i}}$.
+This weight is subsequently utilized in the computation of the final yield $\hat{y_{i}}$.
+First, we compute the query $\mathbf{Q_{i}}$, key $\mathbf{K_{i}}$, and value $\mathbf{V_{i}}$ vectors from the feature $\mathbf{X_{i}}$ using learned linear transformations:
+$\mathbf{Q_{i}}=\mathbf{W_{Q}}\mathbf{X_{i}},\quad\mathbf{K_{i}}=\mathbf{W_{K}}\mathbf{X_{i}},\quad\mathbf{V_{i}}=\mathbf{W_{V}}\mathbf{X_{i}}$ (5)
+where $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, and $\mathbf{W_{V}}$ are the learned weight matrices for the query, key, and value, respectively.
 Next, we calculate the attention scores by taking the dot product of the query and key, scaled by the square root of the key’s dimension $d_{k}$:
-$\text{Attention}(Q_{i},K_{i})=\frac{Q_{i}\cdot{K_{i}}^{\top}}{\sqrt{d_{k}}}$ (9)
-These attention scores are then passed through a softmax function (Goodfellow et al., 2016) to obtain the weights $\alpha_{i}$:
-$\alpha_{i}=\text{softmax}\left(\frac{Q_{i}\cdot{K_{i}}^{\top}}{\sqrt{d_{k}}}\right)$ (10)
-And The softmax function (Goodfellow et al., 2016) is defined as:
+$\text{Attention}(\mathbf{Q_{i}},\mathbf{K_{i}})=\frac{\mathbf{Q_{i}}\cdot{\mathbf{K_{i}}}^{\top}}{\sqrt{d_{k}}}$ (6)
+These attention scores are then passed through a softmax function (Goodfellow et al., 2016) to obtain the weights $\mathbf{\alpha_{i}}$:
+$\mathbf{\alpha_{i}}=\text{softmax}\left(\frac{\mathbf{Q_{i}}\cdot{\mathbf{K_{i}}}^{\top}}{\sqrt{d_{k}}}\right)$ (7)
+The softmax function converts a vector of $K$ real numbers into a probability distribution of $K$ possible outcomes.
+Given a random input vector $\mathbf{z}=[z_{1},\dots,z_{K}]$ for $i=1,\dots,K$, the softmax function (Goodfellow et al., 2016) is defined as:
 $\text{softmax}(\mathbf{z})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{K}e^{z_{j}}}$ (8)
-Where $\mathbf{z}=[z_{1},\dots,z_{K}]$ is the input vector for $i=1,\dots,K$. $e^{z_{i}}$ is the exponential of the $i$-th element of the input vector $\mathbf{z}$.
-Finally, we compute the weighted sum of the values $V_{i}$ to obtain the final yield prediction:
-$y_{i}=\sum_{t=1}\alpha_{i}V_{i}$ (12)
-This process allows us to capture the importance of each feature $x_{i}$ in predicting the yield, dynamically adjusting the weights based on the learned attention scores.
-3.3 Generating field-level dataset for pretraining
+where $e^{z_{i}}$ is the exponential of the $i$-th element of the input vector $\mathbf{z}$.
+Finally, we compute the weighted sum of the values $\mathbf{V_{i}}$ to obtain the final yield prediction:
+$\hat{y_{i}}=\sum_{t=1}\mathbf{\alpha_{i}}\mathbf{V_{i}}$ (9)
+3.3 Generating APSIM field-level dataset for pretraining
+The variables comprising the field-level dataset were previously introduced (Sec.
+2.2); this section provides additional processing details.
+Besides the input weather data obtained from IEM, specific management parameters were also necessary for model calibration (Sec.
+3.3.1).
+Following the simulation, the resulting data required a optimizing process prior to its utilization (Sec.
+3.3.2).
 3.3.1 APSIM management parameters
-In addition to weather input data, the APSIM model requires management parameters for calibration.
-Simulation methods vary depending on the scale of the study.
-For field-level crop areas, management information is obtained directly from farmers and used as input, while other unknown parameters are adjusted to optimize the model (Shahhosseini et al., 2021).
-For county-level crop areas, a common practice is to identify representative fields within those counties, average their management parameters, and apply these averages to represent the entire county (Puntel et al., 2016).
-When the study area is very large, as in our research on simulating all counties in the U.S.
-Corn Belt, calibrating each county individually became impractical.
-In this case, the typical approach is to define a range for each parameter based on empirical and statistical data, and then randomly combine these parameters within the specified ranges to simulate yield across all counties (Lobell et al., 2015).
+Simulation methodology depends on the study’s scale.
+For field-level areas, management information is obtained directly from farmers as input.
+Other unknown parameters are adjusted to optimize the model (Shahhosseini et al., 2021).
+For county-level areas, identifying representative fields is common practice.
+Their management parameters are averaged, and then applied to represent the entire county (Puntel et al., 2016).
+Calibrating each county individually became impractical for very large study areas like our U.S.
+Corn Belt simulation.
+In such cases, the typical approach defines a parameter range based on empirical and statistical data.
+These parameters are then randomly combined within specified ranges, simulating yield across all counties (Lobell et al., 2015).
 Given the vast area of cornfields in our study, it was not feasible to calibrate these parameters for every individual station.
-Therefore, we established a general range for these parameters to encompass all possible values across different fields.
-This approach ensured that the model remained applicable to a wide range of conditions while maintaining a reasonable level of accuracy in its predictions.
-We selected an appropriate adjustment range for the most important management parameters (Table 5).
+Therefore, we established a general parameter range encompassing all possible values across different fields.
+This approach ensured model applicability across a wide range of conditions, maintaining reasonable prediction accuracy.
+Some reasonable adjustment ranges for key management parameters were collected from prior research and USDA statistics (Table 6).
 Sowing density determines the number of plants per unit area, directly affecting competition for resources such as light, water, and nutrients.
-Based on statistics from the USDA NASS (USDA, 2020), the range for sowing density was set at 6-9 plants/m2.
+Based on statistics from the USDA NASS (USDA, 2020), the range for sowing density was set at 6-9 plants/m2, values of 6,7,8, and 9 were used in the simulation.
 Sowing dates are crucial as they determine the crop’s growth cycle and its interaction with seasonal weather patterns.
@@
 The most commonly used nitrogen fertilizers for corn production in North America are anhydrous ammonia, urea, and urea-ammonium nitrate solutions (Herzmann et al., 2004).
-The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N).
+The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N), values of 200, 250, and 300 were used in the simulation.
 Initial soil water content is important for establishing the starting conditions for the model’s simulation of soil moisture dynamics throughout the growing season.
-The initial soil water content was set between 40% and 60%.
+The initial soil water content was set between 40% and 60%, values of 40%, 50%, and 60% were used in the simulation.
 Table 6: Summary of management parameters in APSIM simulation
@@
 Intial soil water 40%-60% (Lobell et al., 2015)
-3.3.2 Filtering simulated data based on soil moisture
-PB models overlook certain environmental factors and real-world influences, leading to significant discrepancies between some simulated and real-world data.
-In the simulation process, the simulated intermediate variable is soil moisture, and the final simulation output is corn yield.
-It is reasonable to assume that with multiple intermediate variables, the final simulation results will have greater errors than the intermediate variables themselves.
-Additionally, soil moisture is influenced by external factors in a relatively direct manner, unlike yield, which depends on complex plant growth processes.
-Therefore, soil moisture was used as a benchmark to filter the simulated dataset.
-We trained a linear regression (LR) model on the GEE county-level dataset to predict soil moisture from weather data and used it to filter the most accurate simulated data in the APSIM field-level dataset.
-Simulated soil moisture with higher accuracy and its corresponding corn yield were retained to ensure that the model benefited from the most reliable simulations.
-Notably, we used only simulated data from years before the test year, eliminating the risk of future information leakage.
-3.4 Generating county-level dataset for finetuning and testing
-3.4.1 Feature Extraction Within Cropland
-After downloading various feature data from GEE, we needed to identify the cropland areas.
-We used the cropland mask from the Cropland Data Layer (CDL) (USDA-NASS, 2017) to extract the feature data corresponding to cropland.
-The Cropland Data Layer is an annual, raster-based crop-specific land cover dataset produced by the USDA.
-It has a resolution of 30 meters and offers detailed crop type and location data across the United States.
-For our corn yield prediction study in twelve U.S.
-Corn Belt states, we used the CDL to generate a corn field mask in GEE.
-This mask identifies the specific areas where corn is grown, allowing us to focus our analysis on relevant agricultural fields.
+3.3.2 Optimizing APSIM field-level dataset based on soil moisture
+APSIM is not accurate for large-scale corn yield simulation without precise management adjustments, thus requiring optimization of the APSIM field-level pretrain dataset.
+Soil moisture was proposed as a benchmark for filtering high-quality data.
+First, soil moisture dynamics are relatively simple, strongly correlated with precipitation.
+In contrast, corn yield is influenced by multiple factors and complex growth processes, hindering accurate modeling in process-based model.
+Second, simulation processes inherently exhibit cumulative error.
+Due to the relative simplicity of soil moisture modeling, this error is comparatively smaller.
+Therefore, the APSIM field-level dataset was optimized based on the quality of soil moisture data.
+Specifically, an LR model was trained on the GEE county-level dataset.
+It utilized four weather inputs (Tmax, Tmin, PPT, Radn) to predict soil moisture output.
+The trained model subsequently predicted soil moisture for the APSIM field-level dataset using its weather data.
+This prediction was compared against the simulated soil moisture.
+Simulated soil moisture with a mean squared error (MSE) exceeding 0.5 were discarded ensuring dataset quality.
+3.4 Generating GEE county-level dataset for finetuning and testing
+3.4.1 Feature Extraction Within Cornfield
+To focus on the corn portion of the acquired feature data from GEE, the cornfield areas had to be identified.
+The Cropland Data Layer (CDL) cropland mask (USDA-NASS, 2017) was utilized extracting feature data corresponding to cropland.
+The Cropland Data Layer is an annual raster-based dataset.
+It provides crop-specific land cover information produced by the USDA.
+This dataset has a 30-meter resolution.
+The CDL was employed to generate a cornfield mask in GEE, supporting our corn yield prediction study in twelve U.S.
+Corn Belt states.
 3.4.2 Data preprocessing
-Every 16 days, we aggregated all county-level features into a single value to represent agricultural conditions for that period.
-This method followed the experimental setup from our previous research (Ma et al., 2021b; Wang et al., 2025).
-We believed that a 16-day interval was sufficient to capture timely changes in weather and environmental conditions.
-3.5 Developing the KGML framework
+County-level features were aggregated into a single value every 16 days, representing agricultural conditions for that period.
+This process resulted in a data vector for each feature annually for each county.
+This aggregation method followed the experimental setup from previous research (Ma et al., 2021b; Wang et al., 2025).
+3.5 Developing the KGML-SM framework
 3.5.1 Pretraining with APSIM field-level dataset
-The pretraining process began by using a W2S encoder $f_{W2S}$ to learn the relationship between input weather features $w_{i}$ and predicted soil moisture $\hat{s}_{i}$.
-This process was formulated as $\hat{s}_{i}=f_{W2S}(w_{i}),\ \text{where}\ w_{i}\in\mathcal{D}_{\text{sim}}$.
-Then, we used the simulated soil moisture $s_{i}\in\mathcal{D}_{\text{sim}}$ to guide the W2S encoder.
+The pretraining process began by using a W2S encoder $f_{W2S}$ to learn the relationship between input weather features $\mathbf{w_{i}}$ and predicted soil moisture $\mathbf{\hat{s}_{i}}$.
+This process was formulated as $\mathbf{\hat{s}_{i}}=f_{W2S}(\mathbf{w_{i}}),\ \text{where}\ \mathbf{w_{i}}\in\mathcal{D}_{\text{field/pretrain}}$.
+Then, The simulated soil moisture $\mathbf{s_{i}}\in\mathcal{D}_{\text{field/pretrain}}$ is used to guide the W2S encoder.
 The loss function $\mathcal{L}_{\text{SM}}$ was defined as:
-$\mathcal{L}_{\text{SM}}=\frac{1}{N}\sum_{i=1}^{N}(s_{i}-\hat{s}_{i})^{2}$ (13)
-where $N=\lvert\mathcal{N}\rvert$.
-Next, we concatenated the predicted soil moisture with weather data and other features to form the features embedding $x_{i}=[w_{i},o_{i},\hat{s}_{i}]$.
-To predict the final corn yield $\hat{y}_{i}$, we used the attention module $f_{att}$, which took the concatenated features $x_{i}$ as input:
-$\hat{y}_{i}=f_{att}(x_{i})$ (14)
-The model was trained using the simulated corn yield $y_{i}$ from the dataset $\mathcal{D}_{\text{sim}}$, ensuring that predictions align with PB simulations.
-The final yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation.
-It was formulated as:
+$\mathcal{L}_{\text{SM}}=\frac{1}{N}\sum_{i=1}^{N}(\mathbf{s_{i}}-\mathbf{\hat{s}_{i}})^{2}$ (10)
+Next, the predicted soil moisture was concatenated with weather data and other features to form the features embedding $\mathbf{X_{i}}=[\mathbf{w_{i}};\mathbf{o_{i}};\mathbf{\hat{s}_{i}}]$.
+The final corn yield $\hat{y}_{i}$ was predicted using the attention module $f_{att}$, with the concatenated features $\mathbf{X_{i}}$ as input:
+$\hat{y}_{i}=f_{att}(\mathbf{X_{i}})$ (11)
+3.5.2 Drought-aware yield prediction loss function
+Soil moisture is widely recognized to affect crop yield (Ines et al., 2013; Bushong et al., 2016) and numerous studies explored this relationship (Mladenova et al., 2017; Vergopolan et al., 2021; Pignotti et al., 2023).
+However most prior studies lacked an elegant quantitative approach addressing this issue.
+This study introduces a loss function adjusting predicted corn yield based on varying soil moisture levels.
+The final yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation. It was formulated as:
 $\mathcal{L}_{\text{yield}}=\frac{1}{N}\sum_{i=1}^{N}d_{i}\left[(y_{i}-\hat{y}_{i})^{2}+\lambda\max(0,\hat{y}_{i}-y_{i})^{2}\right]$ (12)
@@
 Since soil moisture plays a critical role in crop growth and yield formation, the loss function assigned a higher penalty to drier conditions, encouraging the model to be more responsive to soil moisture variations.
-Additionally, we incorporated an asymmetric penalty term (Ridnik et al., 2021) controlled by the factor $\lambda>0$, which amplifies the loss when the predicted yield $\hat{y}_{i}$ exceeds the true yield $y_{i}$:
+Additionally, there is an asymmetric penalty term (Ridnik et al., 2021) controlled by the factor $\lambda>0$, which amplifies the loss when the predicted yield $\hat{y}_{i}$ exceeds the true yield $y_{i}$:
 $\lambda\max(0,\hat{y}_{i}-y_{i})^{2}$ (14)
@@
 This encourages the model to adopt a conservative approach, reducing the likelihood of overestimating yield, particularly in drought-prone regions where overestimation could lead to inaccurate agricultural planning.
-This formulation ensured that the model not only learned accurate yield predictions but also captured the impact of soil moisture variability and drought stress, leading to more reliable and interpretable results.
+This whole loss function formulation ensured that the model not only learned accurate yield predictions but also captured the impact of soil moisture variability and drought stress, leading to more reliable and interpretable results.
 3.5.3 Finetuning and testing with GEE county-level dataset
-We used the GEE county-level dataset $\mathcal{D}_{\text{train}}$ for finetuning, which contained real-world input features and corresponding USDA NASS yield records.
-The loss functions $\mathcal{L}_{\text{yield}}$ and $\mathcal{L}_{\text{SM}}$ were optimized on this dataset.
-Finally, we evaluated the finetuned model on the testing set $\mathcal{D}_{\text{test}}$, where only the input features were available.
+Following the pretraining of the model on the APSIM field-level dataset $\mathcal{D}_{\text{field/pretrain}}$, finetuning using the GEE county-level dataset $\mathcal{D}_{\text{county/finetune}}$ was required to improve the model for the county-level corn yield prediction task.
+This process involved initially partitioning the data into training, validation, and test sets, followed by predicting corn yield for the target years.
+Specific details regarding the partition and utilization of these datasets are elaborated upon in the experimental setup (Sec.
+4.1).
+The loss functions $\mathcal{L}_{\text{yield}}$ and $\mathcal{L}_{\text{SM}}$ were also optimized on the training and validation datasets.
 4 Experimental Results
@@
 Each experiment was conducted five times with different random seeds, and the final results represented the average across these runs to ensure robustness and reliability.
-We implemented the deep learning models using the PyTorch framework (Paszke et al., 2019) and the traditional ML code with sklearn (Pedregosa et al., 2011).
+We implemented the DL models using the PyTorch framework (Paszke et al., 2019) and the traditional ML code with sklearn (Pedregosa et al., 2011).
 The models were run on A100-SXM4-40GB and A100-SXM4-80GB GPUs.
 For training, we used the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of 0.001 and the ReduceLROnPlateau scheduler.
-Root mean square error ($RMSE$) and the coefficient of determination ($R^{2}$) were used to evaluate the performance of our model.
-$RMSE$ is a measure of the differences between predicted and actual values in a regression model.
-It provides an estimate of the standard deviation of the prediction errors.
-A lower $RMSE$ indicates a better fit of the model to the data.
-The formula for $RMSE$ is:
+Root mean square error ($RMSE$) and the coefficient of determination ($R^{2}$) were used to evaluate the performance of our model. The formulas for $RMSE$ and $R^{2}$ are:
 $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (15)
-$R^{2}$ is a statistical measure that indicates the proportion of the variance in the dependent variable that is predictable from the independent variables.
-An $R^{2}$ value of 1 indicates perfect prediction, while a value of 0 indicates that the model does not explain any of the variance in the response variable.
-The formula for $R^{2}$ is:
 $R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (16)
 where $n$ is the number of observations, $y_{i}$ is the actual value for the $i$-th observation, $\hat{y}_{i}$ is the predicted value for the $i$-th observation, and $\bar{y}$ is the mean of the actual values.
-4.2 Statistical analysis
-Before conducting the experiments, we analyzed the statistics of drought, soil moisture, and corn yield.
-We investigated drought conditions in the United States (Drought.gov, 2024) during August from 2019 to 2023 and present the corresponding maps (Fig.
+4.2 Statistical analysis of drought, soil moisture, and corn yield
+To study the impact of soil moisture on corn yield in the KGML-SM model and provide interpretability, we first analyzed the statistics of drought, soil moisture, and corn yield.
+The analysis aims to determine which regions experienced drought and reduced corn yield, and their relationship with soil moisture.
+4.2.1 Specifying drought area
+We investigated drought conditions in the United States (Drought.gov, 2024) during the corn growth season (June-September) (USDA, 2025) from 2019 to 2023 and present the corresponding maps (Fig.
 3).
-While we focused on August, drought conditions throughout the entire corn growing season from May to October were generally similar.
 In 2019 (Fig.
-3(a)), severe drought was primarily observed in Iowa and Michigan.
-In 2020 (Fig.
-3(b)), severe drought affected Iowa, with North Dakota, South Dakota, Nebraska, and Illinois facing moderate drought.
-2021 (Fig.
-3(c)) saw worsening conditions, with severe drought covering large parts of North Dakota, South Dakota, and Minnesota, while Nebraska and Iowa experienced moderate drought.
+3(a)), drought conditions were predominantly observed in Iowa and Michigan.
+By 2020 (Fig.
+3(b)), Iowa experienced severe drought, while moderate conditions affected North Dakota, South Dakota, Nebraska, and Illinois.
+Conditions deteriorated further in 2021 (Fig.
+3(c)), with severe drought covering large parts of North Dakota, South Dakota, and Minnesota, while Nebraska and Iowa experienced moderate drought.
 In 2022 (Fig.
 3(d)), severe drought persisted in Nebraska and Kansas, with South Dakota and Iowa experiencing moderate drought.
-Finally, in 2023 (Fig.
+In 2023 (Fig.
 3(e)), severe drought was widespread across Nebraska, Kansas, Minnesota, Iowa, and Wisconsin, while Missouri faced moderate drought conditions.
-Figure 3: Drought map of the U.S. in August from (a) 2019 to (e) 2023.
-Before studying the impact of soil moisture on corn yield and its prediction, we specified the relationship between soil moisture and drought.
+Figure 3: Drought map of the U.S. for (1) June - (4) September from (a) 2019 to (e) 2023.
+4.2.2 Soil moisture statistics
+Next, we specified the relationship between soil moisture and drought.
 Fig.
-4 showes the maps of rootzone and surface soil moisture in August from 2019 to 2023.
+4 shows the maps of average rootzone and surface soil moisture during June-September from 2019 to 2023.
 The comparison of the drought maps above with soil moisture data from 2019 to 2023 revealed a strong correlation between drought-affected areas and lower soil moisture levels.
 We also noticed that rootzone moisture was more abundant than surface moisture and that the two spatial distributions were generally similar.
-Figure 4: This figure shows the maps of (1) rootzone and (2) surface soil moisture in August from (a) 2019 to (e) 2023.
-Next, we examined the specific impact of drought on corn yield.
+Figure 4: This figure shows the maps of average (1) rootzone and (2) surface soil moisture during June-September from (a) 2019 to (e) 2023.
+4.2.3 Drought impact on corn yield
+Finally, we examined the specific impact of drought on corn yield.
 We plotted the difference in corn yield for each year from 2019 to 2023 relative to the previous year’s yield (Fig.
@@
 In 2019 (Fig.
-5(a)), drought-affected areas included Iowa and Michigan, however, significant yield reductions were observed only in Michigan.
+5(a)), drought-affected areas included Iowa and Michigan; however, significant yield reductions were observed only in Michigan.
 In 2020 (Fig.
@@
 In 2022 (Fig.
-5(d)), drought-affected states included Nebraska, Kansas, South Dakota, and Iowa, but yield declines were primarily observed in North Dakota, Kansas, and South Dakota.
+5(d)), drought-affected states included Nebraska, Kansas, South Dakota, and Iowa, but yield declines were primarily observed in Nebraska, Kansas, and South Dakota.
 In 2023 (Fig.
@@
 Thus, we concluded that drought consistently led to a decline in corn yield.
-We summarized the drought-affected areas that experienced yield reductions in Table 6.
+Figure 5: The difference in corn yield between each year from (a) 2019 to (e) 2023 and the previous year, with negative values indicating a reduction in yield.
+4.2.4 Statistical summary of drought-affected states
+Based on the above statistical analysis, we summarized the drought-affected areas that experienced yield reductions in Table 7.
 In the experiment, we focused on these states that were affected by drought and experienced yield reductions.
-Figure 5: The difference in corn yield between each year from (a) 2019 to (e) 2023 and the previous year, with negative values indicating a reduction in yield.
 Table 7: Summary of obviously yield-reduced states within drought-affected areas.
@@
 2023 NE,KS,MN,IA,WI,MO IA,MO
-4.3 KGML-SM performance comparison with traditional ML models
-To validate the superiority of our KGML-SM model, we compared it with four traditional machine learning models: LR, multilayer perceptron (MLP), ridge regression (RR), and random forest (RF).
+4.3 Evaluation results
+To validate the superiority of our KGML-SM model, we compared it with some commonly used ML models in remote sensing: LR, multilayer perceptron (MLP), ridge regression (RR), and random forest (RF).
 LR (Bishop and Nasrabadi, 2006) is a simple statistical method that models the relationship between a dependent variable and one or more independent variables by fitting a linear equation to the data.
@@
 RF (Breiman, 2001) is an ensemble learning method that constructs multiple decision trees during training and aggregates their predictions to improve accuracy and robustness.
-The results showed that our KGML-SM model consistently outperformed the other ML models across all years (Table 7).
-Based on the results, RF performed the best among traditional ML models, indicating its strong ability to capture complex relationships in the data.
-RR performed slightly worse than RF, with slightly higher RMSE values, suggesting that regularization helped improve predictions but was not as effective as ensemble learning.
-LR ranked next, showing higher RMSE values, likely due to its inability to model non-linear relationships effectively.
-MLP performed the worst, with the highest RMSE values in most years, indicating that it struggled to generalize well, possibly due to overfitting or insufficient training data.
+The results showed that our KGML-SM model consistently outperformed other ML models across all years.
+As shown in Table 8, RF performed the best among traditional ML models, indicating its strong ability to capture complex relationships in the data (Couronné et al., 2018).
+RR performed slightly worse than RF, with slightly higher RMSE values, suggesting that regularization helped improve predictions but was not as effective as ensemble learning (Schreiber-Gregory, 2018).
+LR ranked next, showing higher RMSE values, likely due to its inability to model non-linear relationships effectively (Manual, 2013).
+MLP performed the worst, with the highest RMSE values in most years, indicating that it struggled to generalize well, possibly due to overfitting or insufficient training data (Caruana et al., 2000).
 Table 8: Comparison with traditional ML models
@@
 2023 1.071 0.807 1.226 0.737 1.225 0.738 1.140 0.776 1.196 0.791
-4.4 Ablation study of different compoments in KGML-SM model
+4.4 Ablation study of different components in KGML-SM model
 To further demonstrate the contribution of each module in our KGML-SM model, we conducted a series of ablation studies.
 We began with a baseline model trained using only the GEE county-level dataset and the attention module (Att) (Wang et al., 2025).
-Next, we incorporated the simulated dataset to evaluate its impact on model performance (Att+sim).
+Next, we incorporated the APSIM field-level dataset pretraining to evaluate its impact on model performance (Att+sim).
 We then introduced the W2S encoder to assess its effectiveness in integrating soil moisture information (Att+sim+W2S).
@@
 Table 9: Comparison of different methods and their components.
-Method Attention module Simulated data W2S encoder Drought-aware loss
+Method Attention module Field-level data pretraining W2S encoder Drought-aware loss
 Att
@@
 KGML-SM
-Through an ablation study of all model components (Table 9), we found that the simulated dataset contributed the most to performance improvement.
-This indicated that our APSIM field-level dataset effectively captured real-world data patterns, playing a crucial role in enabling the model to learn the relationship between corn yield and environmental variables.
-Additionally, our drought-aware loss function significantly enhanced performance, demonstrating that incorporating prior knowledge of the real-world relationship between drought and corn yield can effectively guide the training and prediction of ML models.
+Through the ablation study of all model components (Table 10), we found that the APSIM field-level dataset pretraining contributed the most to performance improvement.
+This indicated that our APSIM field-level dataset effectively captured county-level data patterns, playing a crucial role in enabling the model to learn the relationship between corn yield and agricultural variables.
+Additionally, our drought-aware loss function significantly enhanced performance in most years, demonstrating that incorporating prior knowledge of the relationship between drought and corn yield can effectively guide the training and prediction of ML models.
 Table 10: Ablation study of different components in KGML-SM model
@@
 RMSE R2 RMSE R2 RMSE R2 RMSE R2
-2019 0.964 0.741 0.974 0.752 1.087 0.685 1.268 0.57
+2019 0.964 0.741 0.974 0.732 1.087 0.715 1.268 0.57
 2020 0.98 0.792 0.981 0.783 1.003 0.77 1.011 0.77
 2021 1.104 0.836 1.127 0.811 1.143 0.814 1.195 0.808
-2022 1.085 0.837 1.119 0.802 1.244 0.802 1.315 0.779
+2022 1.085 0.837 1.119 0.802 1.144 0.802 1.315 0.779
 2023 1.071 0.807 1.101 0.805 1.114 0.781 1.201 0.766
@@
 In Fig.
-6, we plotted the error maps for the KGML-SM and LR model results from 2019 to 2023.
+6, we plotted the error maps of the prediction results from 2019 to 2023 for the two best performed models: KGML-SM and RF model.
 In 2019 (Fig.
-6 (a)), the lower part of Michigan shows noticeable overestimation in the attention model, whereas our KGML-SM model performed well.
+6(a)), the lower part of Michigan showed noticeable overestimation in the RF model, whereas our KGML-SM model performed well.
 In 2020 (Fig.
-6 (b)), although the derecho storm in Iowa also had some impact on our KGML-SM model, we found that the LR model exhibited more severe overestimation.
+6(b)), although the derecho storm (Hosseini et al., 2020) in Iowa also had some impact on our KGML-SM model, we found that the RF model exhibited more severe overestimation.
 In 2021 (Fig.
-6 (c)), overestimation in the upper-left region, particularly in North Dakota and South Dakota, is significantly reduced.
+6(c)), overestimation in the upper-left region, particularly in North Dakota and South Dakota, was significantly reduced.
 In 2022 (Fig.
-6 (d)), overestimation decreases notably in South Dakota and Kansas, though the effect in Kansas is less pronounced.
+6(d)), overestimation decreases notably in North Dakota, South Dakota, and Kansas, though the effect in Kansas was less pronounced.
 In 2023 (Fig.
-6 (e)), Missouri showed a clear reduction in overestimation, while the effect in Iowa is less noticeable.
+6(e)), Iowa showed a reduction in overestimation, while the effect in Missouri was less noticeable.
 Overall, our model shows a notable improvement in addressing the issue of overestimation in drought-affected areas.
-Figure 6: The error map for (1) KGML-SM and (2) LR model from (a) 2019 to (e) 2023.
+Figure 6: The error map for (1) KGML-SM and (2) RF model from (a) 2019 to (e) 2023, with blue indicating overestimation.
 4.6 Understanding model bias and prediction-observation alignment
-We generated scatter plots for the KGML model and the LR model to analyze their prediction performance (Fig.
+To analyze model prediction performance, we generated scatter plots for the KGML-SM model and the RF model (Fig.
 7).
-These plots help visualize the relationship between observed and predicted values, revealing patterns of overestimation, underestimation, and potential prediction biases across different years.
+These plots helped visualize the relationship between observed and predicted values, revealing patterns of overestimation, underestimation, and potential prediction biases across different years.
 In 2019 (Fig.
-7 (a)), the KGML-SM model produced a noticeably narrower distribution, indicating a lower spread in prediction errors.
+7(a)), the KGML-SM model presented a noticeably narrower distribution, indicating a lower spread in prediction errors.
 In 2020 (Fig.
-7 (b)), The predictions of the KGML-SM model were noticeably more concentrated along the diagonal and exhibited symmetry on both sides, whereas the LR model produced more dispersed predictions in high-yield regions.
+7(b)), the predictions of the KGML-SM model were noticeably more concentrated along the diagonal and exhibited symmetry on both sides, whereas the RF model produced more dispersed predictions in high-yield regions.
 In 2021 (Fig.
-7 (c)), the LR model exhibited prediction collapse, where certain observed values correspond to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations.
+7(c)), the RF model exhibited prediction collapse, where certain observed values correspond to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations.
 In 2022 (Fig.
-7 (d)), the KGML-SM model maintained a narrower and more concentrated prediction distribution, indicating improved robustness.
-In 2023 (Fig.
-7 (e)), the KGML-SM model continued to avoid prediction collapse, demonstrating more stable performance compared to the attention model.
+7(d)) and 2023 (Fig.
+7(e)), the KGML-SM model maintained a narrower and more concentrated prediction distribution.
 This comparison highlighted the advantage of the KGML-SM model in mitigating prediction collapse and improving overall robustness across different years.
-Figure 7: The scatter plot for (1) KGML-SM and (2) LR model from (a) 2019 to (e) 2023.
+Figure 7: The scatter plot for (1) KGML-SM and (2) RF model from (a) 2019 to (e) 2023.
 5 Discussion
-Specifically, we discussed the role of soil moisture in model prediction from three perspectives:
+In this section, we explore the role of soil moisture in model prediction from four questions:
   • 1.
-    How did soil moisture influence model prediction spatially?
+    (Sec. 5.1) How did soil moisture influence model prediction spatially?
   • 2.
-    How did soil moisture affect model performance throughout the corn growth season?
+    (Sec. 5.2) How did soil moisture affect model performance throughout the corn growth season?
   • 3.
-    how did soil moisture contribute to model prediction in drought and non-drought regions?
-We leveraged the attention mechanism to weight input features, capturing each feature’s contribution to the model, to explore these questions.
+    (Sec. 5.3) how did soil moisture contribute to model prediction in drought and non-drought regions?
+  • 4.
+    (Sec. 5.4) How to interpret the observed inaccuracies in corn yield predictions based on soil moisture?
 5.1 Spatial influence of soil moisture on model prediction
-We visualized the attention scores of soil moisture across twelve states in the U.S.
+To answer the first question, we visualized the attention scores of soil moisture across twelve states in the U.S.
 Corn Belt from June to August over the years 2019 to 2023 (Fig.
@@
 The attention map highlighted how the model’s reliance on soil moisture varied across different growth stages and drought conditions, allowing us to assess whether soil moisture has a greater impact on the model in drought-affected areas.
+To enhance visualization, we normalized attention values within each year.
+Consequently, analysis should focus on attention trends across regions within the same year rather than comparing values across different years.
 Figure 8: The attention visualization of soil moisture on corn yield prediction in different regions.
@@
 These values were then normalized based on the maximum value in this year to ensure a consistent color scale, making the distribution more visually interpretable.
-In June, during the early growth period, the attention scores for soil moisture were generally lower across all four years, suggesting that soil moisture is less critical at this stage.
-This might be due to cooler temperatures and lower evaporation rates, which reduced the immediate need for water.
+In June, during the early growth period, the attention scores for soil moisture were generally lower across all five years, This might be because the temperature was not too high and drought conditions were not severe at this time, leading to a less pronounced correlation between soil moisture and corn yield.
 In July, attention scores increased, particularly in drought-affected areas, as soil moisture became more influential during active growth and vegetative stages.
@@
 From 2019 to 2023, the attention distribution of soil moisture in corn yield prediction exhibited noticeable variations, aligning with the drought severity in different years.
-Note that, for better visualization, we normalized the attention values within each year.
-Therefore, we only need to focus on the differences across regions within the same year, rather than making cross-year comparisons.
 In early 2019 (Fig.
-8(a1)), the attention was concentrated in Iowa, reflecting the drought-affected states.
+8(a1)), attention concentrated in Iowa.
 Subsequently (Fig.
-8(a3)), the regions with higher attention values gradually expanded, but Iowa remained the area with the highest attention.
-In 2020 (Fig.
-8(b)), the high-attention regions expanded to include North Dakota, South Dakota, and Nebraska, corresponding to widespread drought in USDA drought map.
-In 2021 (Fig.
-8(c)), the attention intensified over North Dakota, South Dakota, Minnesota, and Nebraska, mirroring severe drought conditions.
-By 2022 (Fig.
-8(d)), high-attention areas were mainly in Nebraska, Kansas, South Dakota, and Iowa.
-Additionally, over time, there was a trend of drought expanding toward the central region.
+8(a3)), regions with higher attention values expanded gradually, though Iowa retained the highest attention, aligning with observed drought conditions that year.
+During 2020 (Fig.
+8(b)), high-attention regions were initially concentrated in Iowa and Nebraska, later expanding to adjacent areas.
+By 2021 (Fig.
+8(c)), attention intensified over North Dakota, South Dakota, Minnesota, and Nebraska.
+In 2022 (Fig.
+8(d)), high-attention areas were primarily located in Nebraska, Kansas, South Dakota, and Iowa.
 In 2023 (Fig.
-8(e)), the attention peaked across Nebraska, Kansas, Minnesota, Iowa, Wisconsin, and Missouri.
-Finally, in August, it reached its peak, and soil moisture became more important in most states in the central region.
-This above analysis underscored the dynamic role of soil moisture in model prediction, with its importance intensifying during key growth stages and under severe drought conditions.
-The elevated attention scores in drought-prone regions suggested that soil moisture serves as a crucial predictor for accurately estimating corn yield in water-stressed environments.
+8(e1)), attention initially focused on Kansas, expanding thereafter across a majority of states.
+Ultimately, significant attention emerged in Nebraska, Kansas, Minnesota, Iowa, Wisconsin, and Missouri, indicative of substantial drought conditions (Fig.
+8(e3)).
+Overall, the period from 2019 to 2023 exhibited a trend of drought conditions and corresponding attention expanding towards the central United States, consistent with actual observations.
+The reason for the incomplete consistency between the two could be that although drought was present in the region, it did not severely impact corn yield.
+Alternatively, even without drought in a particular area, other factors, such as the derecho storm, might have affected corn yield.
+This analysis underscored the dynamic role of soil moisture in model prediction, with its importance intensifying during key growth stages and under severe drought conditions.
 5.2 Temporal role of soil moisture during the corn growth season
-To illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—every 16 days from June to September.
-We found that the VIs had the highest influence on the model around August, which aligned with findings from previous studies (Johnson, 2014).
-In this period, VIs shows the strongest correlation with corn yield because this period corresponds to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield.
-High VIs in this timeframe indicate optimal chlorophyll content, canopy development, and water availability, which are critical for photosynthesis and grain formation.
-Additionally, we found that weather data showed a significant increase in attention around July in 2020 and 2022, and the periods with noticeable attention spikes aligned with the trends of VIs.
+To answer the second question and illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—every 16 days from June to September in drought-affected states (Fig.
+9).
+We found that the VIs had the highest influence on the model around August, which aligned with findings from previous study (Johnson, 2014).
+In this period, VIs showed the strongest correlation with corn yield because this period corresponded to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield (Huang et al., 2014).
+High VIs in this timeframe indicate optimal chlorophyll content, canopy development, and water availability, which are critical for photosynthesis and grain formation (Wang et al., 2005).
+Additionally, we found that weather data showed a significant increase in attention around July in 2020 (Fig.
+9(b)) and 2022 (Fig.
+9(d)), and the periods with noticeable attention spikes aligned with the trends of VIs.
 This phenomenon was more pronounced in soil moisture, as soil moisture attention showed a strong correlation with VIs in all years except 2022.
 In 2022, a slight increase in soil moisture attention could still be observed near the VIs peak.
-This may be because VIs during these periods are closely related to certain weather data and soil moisture data.
-For example, NDWI increases with higher PPT because more rainfall enhances soil moisture and plant water content, GCI decreases with excessive Tmax because extreme heat can degrade chlorophyll and reduce plant vitality, etc.
+This might be because VIs during these periods are closely related to certain weather data and soil moisture data.
+For example, NDWI increases with higher PPT because more rainfall enhances soil moisture and plant water content (Ashok et al., 2021).
 This indicated that our attention mechanism effectively captured feature importance over time dimension.
@@
 5.3 Statistical impact of soil moisture in drought and non-drought regions
-The box plot illustrates the distribution of soil moisture attention across all counties from 2019 to 2023 (Fig.
+To answer the third question, The box plot was used to illustrate the distribution of soil moisture attention across all counties from 2019 to 2023 (Fig.
 10).
-Each point in this box plot represents a county, displaying the comparison of attention between red-marked drought regions and blue-marked non-drought regions from 2019 to 2023.
-Through this visualization, we aimed to analyze the differences in soil moisture attention between drought-affected and non-drought counties across all years.
-In 2021, the most severe drought year, both the variance and magnitude of soil moisture attention were the highest.
-We also observed that attention values were notably high in 2019, which aligns with our experimental results, where our KGML-SM model achieved the most significant improvement over other models.
-Since our model better captured the impact of soil moisture, it achieved the most significant improvement in years when soil moisture has a greater influence.
-In contrast, the values for 2020, 2022, and 2023 were more concentrated, with fewer extreme values.
-Overall, in all years, soil moisture attention in drought-affected regions was consistently higher than in non-drought regions, indicating that soil moisture had a greater impact on the model in these areas.
+Each point in this box plot represents a county, displaying the comparison of attention between red-marked drought-affected areas and blue-marked non-drought areas from 2019 to 2023 based on Table 7.
+Across all five years of attention values, drought-affected areas exhibited fewer outliers compared to non-drought areas, suggesting that the model’s attention to soil moisture is more stable in drought-affected regions.
+Additionally, in all years, the median of soil moisture attention in drought-affected areas was consistently higher than in non-drought areas, indicating that soil moisture had a greater impact on the model in these areas.
 Figure 10: The boxplot of soil moisture attention in drought-affected areas (D--A) and non-drought areas (N--D) across all counties.
+5.4 Interpreting corn yield prediction errors via soil moisture prediction anomalies
+To answer the fourth question, we plotted the absolute error map of both soil moisture and corn yield predictions (Fig.
+11).
+In our modeling process, we treated soil moisture as an intermediate variable to guide the corn yield output, so when corn yield prediction was inaccurate, the corresponding soil moisture might also exhibit anomalies.
+This analysis aims to conduct interpretability analysis of inaccurate corn yield predictions.
+In 2019 (Fig.
+11(a)), the drought-affected state Michigan exhibited relatively accurate corn yield predictions from the KGML-SM model; however, the soil moisture output was less accurate.
+This suggested that soil moisture was challenging to predict in drought-affected areas.
+In 2020 (Fig.
+11(b)), inaccurate corn yield prediction in Iowa was primarily attributable to the derecho storm’s impact (Hosseini et al., 2020), which soil moisture could not reflect.
+In 2021 (Fig.
+11(c)), drought in North and South Dakota resulted in inaccurate corn yield predictions, and corresponding soil moisture predictions were also found to be inaccurate.
+In 2022 (Fig.
+11(d)), Nebraska, Kansas, and South Dakota experienced drought, and their soil moisture predictions were also inaccurate.
+In 2023 (Fig.
+11(e)), Iowa and Missouri were affected by drought.
+Iowa’s corn yield and soil moisture prediction were accurate, which indicated that our model performed well.
+But the prediction for Missouri exhibited errors, despite relatively accurate corresponding soil moisture predictions for the state.
+According to report (Ward, 2023), some damaging winds and hail significantly impacted corn field, causing root lodging in late June and early July, particularly around Mooresville and Rockport, Missouri.
+Hail damage was also reported at Garden City, Missouri.
+These types of mechanical damage were not reflected in standard soil moisture data and constituted yield-limiting factors likely contributing to the corn yield prediction inaccuracies in Missouri.
+The analysis indicates that a significant portion of the error in corn yield prediction is related to soil moisture prediction error, given the key role soil moisture plays in corn growth; thus, if soil moisture is not accurately predicted, the accumulated error in modeling will lead to inaccurate corn yield prediction.
+Figure 11: The absolute error map for (1) corn yield prediction and (2) soil moisture prediction from (a) 2019 to (e) 2023.
 6 Conclusion
-In this study, we proposed the KGML-SM model, which first supplements the limited soil moisture data with simulated data.
-It then learned the PB relationship between weather and soil moisture.
-Additionally, based on prior knowledge that drought leads to yield overestimation, we introduced a drought-aware loss function to mitigate this issue in drought-affected areas.
-We analyzed 12 states in the U.S.
-Corn Belt to examine the impact of soil moisture on corn yield prediction.
-Our method consistently outperformed baseline models across multiple test years.
-We also explore the spatial and temporal impact of soil moisture on the model by visualizing attention, highlighting when and where the model focuses more on soil moisture.
-In the future, we aim to apply transfer learning to adapt models trained on well-studied regions with simulated data to areas with limited data availability.
+In this study, we proposed the KGML-SM model, where the W2S encoder is designed to capture the influence of weather on soil moisture, and the attention module is employed to weight different input features for final corn yield prediction.
+To address the known issue that drought conditions often lead to yield overestimation, we introduced a drought-aware loss function to mitigate this problem in drought-affected regions.
+We constructed both an APSIM field-level dataset and a GEE county-level dataset, leveraging a process-based model to learn the corn growth process by pretraining the KGML-SM model on the APSIM field-level dataset and finetuning it on the GEE county-level dataset.
+Our analysis covered 12 states in the U.S.
+Corn Belt to investigate the impact of soil moisture on corn yield prediction.
+The proposed method consistently outperformed baseline models across multiple test years.
+Furthermore, we studied the spatial and temporal influence of soil moisture through attention visualization, revealing when and where the model places greater focus on soil moisture.
+Finally, based on the relationship between soil moisture and corn yield prediction, we investigated the causes of inaccuracies in corn yield prediction and provided an explanation.
+In future work, we aim to apply transfer learning techniques to adapt models trained on well-studied regions with abundant simulated data to regions with limited data availability.
+Funding
+This work was supported by the United States Department of Agriculture (USDA) National Institute of Food and Agriculture (NIFA) Agriculture and Food Research Initiative Foundational Program (Award No.
+2022-67021-36468); and the USDA NIFA Hatch Project (Accession No.
+7005141).

### v3

@@
 Abstract
-Remote sensing (RS) technique, enabling the non-contact acquisition of extensive ground observations, is a valuable tool for crop yield predictions.
-Traditional process-based models struggle to incorporate large volumes of RS data, and most users lack understanding of crop growth mechanisms.
-In contrast, machine learning (ML) models are often criticized as “black boxes” due to their limited interpretability.
-To address these limitations, we utilized Knowledge-Guided Machine Learning (KGML), a framework that leverages the strengths of both process-based and ML models.
-Existing works have either overlooked the role of soil moisture in corn growth or did not embed this effect into their models.
-To bridge this gap, we developed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, treating soil moisture as an intermediate variable in corn growth to emphasize its key role in plant development.
-Additionally, based on the prior knowledge that the model may overestimate under drought conditions, we designed a drought-aware loss function that penalized predicted yield in drought-affected areas.
-Our experiments showed that the KGML-SM model outperformed other traditional ML models.
-We explored the relationships between drought, soil moisture, and corn yield prediction by assessing the importance of different features within the model, and analyzing how soil moisture impacts predictions across different regions and time periods.
-Finally we provided interpretability for prediction errors to guide future model optimization.
+Remote sensing (RS) provides large-scale, non-contact observations that are valuable for crop yield prediction.
+Process-based models are based on crop growth mechanisms but often struggle with large RS datasets and require heavy calibration.
+Machine learning (ML) models can use RS data effectively but are often criticized as “black boxes” with low interpretability.
+To overcome these issues, we proposed the Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework, which combined the strengths of both approaches.
+Soil moisture was used as an intermediate variable, linking weather and crop growth and serving as a key factor in drought impacts on yield.
+This design improved interpretability by tracing yield prediction errors to soil moisture estimates.
+We also added a drought-aware loss function that penalized overestimation in dry regions, making the model more robust under drought stress.
+This study focused on the U.S.
+Corn Belt, covering 12 states and more than 800 counties from 2019 to 2023.
+We constructed two datasets: a field-level dataset generated from the Agricultural Production Systems sIMulator (APSIM) and a county-level dataset built with satellite-based MODIS RS data and gridded climate data.
+Model performance was evaluated against USDA-NASS county-level yield records.
+KGML-SM achieved lower errors than traditional ML baselines, with an RMSE of 1.071 t/ha and an $R^{2}$ of 0.807 in 2023.
+Attention-based analysis further revealed the role of drought and soil moisture in yield prediction.
+Overall, KGML-SM improves both accuracy and interpretability, offering insights for future model development and climate-resilient agriculture.
 keywords:
-Corn yield prediction; Knowledge-guided Machine learning; Soil moisture; drought-aware learning
-journal: Journal
+Crop modeling; Remote sensing; Process-based models; Attention mechanism; Drought-aware loss function; Agricultural informatics
+journal: Remote Sensing of Environment
 […]
@@
 […]
-[inst3]organization=Research and Development Division, National Agricultural Statistics Service, United States Department of Agriculture,addressline=Washington, postcode=20250, state=DC, country=USA
+[inst3]organization=U.S. Department of Agriculture, National Agricultural Statistics Service, city=Washington, postcode=20250, state=DC, country=USA
+[…]
+[inst4]organization=U.S. Department of Agriculture, Agricultural Research Service, Genetics and Sustainable Agricultural Research Unit, city=Mississippi State, postcode=39762, state=MS, country=USA
+[…]
+[inst5]organization=NASA Goddard Space Flight Center, city=Greenbelt, postcode=20771, state=MD, country=USA
 1 Introduction
-Corn, as a primary crop, plays a vital role in the U.S.
+Corn, as a primary crop, plays a vital role in U.S.
 agriculture, supporting food security as well as the animal feed and biofuel industries (Graham et al., 2007; Thompson, 1969).
 Accurate yield prediction is crucial for effective resource management and economic stability (Kucharik and Ramankutty, 2005).
-Achieving accurate prediction across large areas and diverse environmental conditions is challenging (Weiss et al., 2020).
-Traditional process-based models (Weiss et al., 2007) are built on a deep understanding of the physical, chemical, and biological processes within crop growth systems (Weiss et al., 2004).
-These models use equations derived from scientific principles to simulate the behavior of crops, weather, and soil interactions (Weiss et al., 2002), making them highly interpretable and reliable in scenarios where the underlying processes are well understood (Puntel et al., 2016; Shahhosseini et al., 2021).
-Several promising modeling frameworks are designed for simulating and predicting agricultural production systems, such as the Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), the Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), the WOrld FOod STudies (WOFOST) (Van Diepen et al., 1989), and the Agricultural Policy / Environmental eXtender (APEX) (Williams and Izaurralde, 2010).
-Many studies (Zhen et al., 2023, 2022) have successfully used these models to predict crop yields.
-However, process-based models have several limitations.
-Plant growth is highly complex, and these models rely on a limited set of fixed input features, which are insufficient to accurately capture the intricacies of this process.
-Moreover, process-based models require extensive manual parameter tuning, limiting their applicability for applications across vast regions.
+Nevertheless, achieving accurate prediction across large areas and diverse environmental conditions remains challenging (Lobell et al., 2015).
+Climate extremes such as drought pose a major obstacle to accuracy, particularly at the county level where heterogeneous management practices and soil conditions introduce additional variability (Rigden et al., 2020).
+Traditional process-based models are grounded in a detailed understanding of the physical, chemical, and biological processes of crop growth (Puntel et al., 2016).
+These models use equations derived from scientific principles to simulate crop, weather, and soil interactions, which makes them highly interpretable and reliable when underlying processes are well characterized (Shahhosseini et al., 2021).
+Prominent frameworks include the Agricultural Production Systems sIMulator (APSIM) (McCown et al., 1996), the Decision Support System for Agrotechnology Transfer (DSSAT) (Jones et al., 2003), the Agricultural Policy/Environmental eXtender (APEX) (Williams and Izaurralde, 2010), and the Ecosys Modelling Project (ecosys) (Grant, 2001), all of which have been successfully applied to crop yield prediction (Zhen et al., 2023, 2022; Asseng et al., 2013; Schauberger et al., 2017).
+Yet plant growth is highly complex, and the reliance on a limited set of fixed inputs prevents these models from fully capturing this complexity (Lobell et al., 2015; Müller et al., 2017).
+In addition, they require extensive manual parameter tuning, which constrains their use across large regions (Kimball et al., 2023; Zhen et al., 2022).
 In recent years, remote sensing (RS) has provided several benefits for corn yield prediction (Lobell et al., 2015).
 It enables the collection of large-scale real-time data across vast agricultural areas.
-This data, typically combined with machine learning (ML), facilitates accurate and efficient large-scale yield prediction.
-Deep learning (DL) (Goodfellow et al., 2016), a widely used ML method, has the advantage of leveraging a large amount of RS data by automatically learning features from big datasets.
-This removes the need for manual feature engineering, enabling the model to capture complex patterns and relationships between crop environmental factors and crop yield (He et al., 2016).
-Many studies applying DL to crop yield prediction have shown its advantages in performance improvement.
-You et al.
-(2017) introduced a scalable approach to crop yield prediction using modern representation learning rather than traditional hand-crafted features, achieving impressive results.
-Wang et al.
-(2018) demonstrated soybean yield predictions in Argentina and obtained reliable results in Brazil using transfer learning, even with limited data.
-Ma et al.
-(2021a) developed a county-level corn yield prediction model using a Bayesian Neural Network across 12 states in the Corn Belt, showing the potential for large-scale corn yield prediction tasks.
-However, despite the popularity of DL models in corn yield prediction tasks, they have certain limitations.
-First, DL models require a large amount of data to learn the relationships between inputs and outputs.
-Additionally, DL models often function as a black box, lacking reliable interpretability of the prediction results.
-Specifically, crop growth involves complex biological processes among soil, weather, and corn, which are challenging for DL models to fully capture.
-Therefore, how to leverage advantages of process-based models and DL models to address these challenges has become an important research topic.
-Knowledge-guided machine learning (KGML) (Karpatne et al., 2022) aims to integrate scientific knowledge in ML frameworks to achieve better performance, scientific consistency, and explainability of results.
-KGML involves three main approaches: Knowledge-Guided Learning, Knowledge-Guided Architecture, and Knowledge-Guided Pretraining.
-Knowledge-Guided Learning incorporates scientific knowledge, such as physical laws, into the learning algorithm by modifying the loss function.
-It guides the optimization process toward scientifically consistent and generalizable solutions (Daw et al., 2022; Bao et al., 2021).
-Knowledge-Guided Architecture embeds scientific knowledge directly into the structure of the ML model.
-By designing knowledge-guided neural network architectures, researchers can enable models to automatically capture physical properties or conservation laws, which improves both interpretability and robustness (Dugdale et al., 2017; Luo et al., 2023).
-Knowledge-Guided Pretraining uses simulated data or self-supervised learning tasks to pretrain ML models.
-By initializing model parameters using data that encodes scientific principles, this method reduces the need for large amounts of observational data and improves performance in data-scarce scenarios (Licheng et al., 2022; Chen et al., 2023).
-Building on these approaches, KGML has been increasingly used in various domains, such as agriculture, hydrology, and biology, enabling improved data-driven decision-making and predictive modeling.
-Recently, many studies have used KGML to combine process-based and ML models for crop yield predictions (Kimball et al., 2023; Burroughs et al., 2023).
-Since process-based models already provide insights into the crop growth process and generate high-quality simulated data, Knowledge-Guided Pretraining is the most popular approach (He et al., 2023; Yang et al., 2023).
-These studies aim to identify patterns between agricultural features and crop yields from large datasets using ML, while leveraging data simulated by process-based models to provide interpretability.
-Specifically, intermediate variables in crop growth generated by process-based models are used to train ML models, allowing the ML models to simulate specific parts of the crop growth process.
-For example, He et al.
-(2023) integrated physical knowledge from process-based models and enabled the model to adjust for temporal data shifts over the years.
+These data, typically combined with machine learning (ML), facilitate accurate and efficient large-scale yield prediction.
+Deep learning (DL), a subset of ML, is based on neural networks and is particularly effective for high-dimensional RS data, as it can automatically extract features from large datasets and capture complex environment–yield relationships without manual engineering (Goodfellow et al., 2016; He et al., 2016).
+Several studies have demonstrated the potential of DL for yield prediction, such as scalable representation learning (You et al., 2017), transfer learning across regions (Wang et al., 2018), and county-level prediction using Bayesian neural networks (Ma et al., 2021a).
+More recently, adaptive multi-modal fusion frameworks have been proposed to integrate heterogeneous data sources and provide improved accuracy (Mena et al., 2025).
+However, DL methods face important limitations: they require large amounts of training data, often function as black boxes with limited interpretability, and struggle to represent the biological processes linking soil, weather, and crop growth.
+In contrast, process-based models explicitly capture these mechanisms but are less suited for large-scale RS integration.
+This complementarity highlights the need for approaches that combine the strengths of both, motivating recent research on combining these two kinds of models.
+A recent study (Sadeh et al., 2024) also emphasized that relying solely on either process-based or ML models is insufficient for robust yield prediction.
+This growing recognition of complementarity has motivated the emergence of knowledge-guided machine learning (KGML), which seeks to integrate process-based understanding into ML frameworks to enhance both accuracy and interpretability.
+KGML (Karpatne et al., 2022) aims to integrate scientific knowledge into ML frameworks to achieve better performance, scientific consistency, and explainability of results.
+This paradigm involves three main approaches: (i) knowledge-guided learning, which incorporates scientific laws into algorithms through modified loss functions (Daw et al., 2022; Bao et al., 2021); (ii) knowledge-guided architectures, which embed knowledge directly into model structures (Dugdale et al., 2017; Luo et al., 2023); and (iii) knowledge-guided pretraining, which uses simulated data or self-supervised tasks to initialize models (Licheng et al., 2022; Chen et al., 2023).
+In crop yield prediction, most KGML applications rely on DL rather than traditional ML approaches (Burroughs et al., 2023; He et al., 2023; Yang et al., 2023).
+Since process-based models already provide insights into crop growth and generate high-quality simulated data, Knowledge-Guided Pretraining has become the most popular strategy.
+These studies typically use ML to learn yield-relevant patterns while incorporating simulated intermediates as mechanistic links from weather and soil to crop growth (He et al., 2023; Yang et al., 2023, 2024).
+Specifically, He et al.
+(2023) applied physics-guided reweighting with ecosystem-process intermediates—ecosystem autotrophic respiration (Ra), ecosystem heterotrophic respiration (Rh), and net ecosystem exchange (NEE)—to improve time-aware robustness and interpretability.
 Yang et al.
-(2023) used both historical and in-season RS data to better predict key agricultural variables such as grain yield and carbon cycling processes.
+(2023) coupled process-model surrogates with Ensemble Kalman Filter (EnKF) data assimilation to fuse leaf area index (LAI), gross primary production (GPP), and evapotranspiration (ET) for yield prediction.
 Yang et al.
-(2024) presented a knowledge-guided computer vision framework designed to monitor and simulate the growth of individual strawberry fruits, which can predict growth at the individual fruit level by integrating image-based data and dynamic simulations.
-Although these KGML models have made significant progress in crop yield prediction, all existing KGML-based studies on crop yield predictions have overlooked the critical role of soil, particularly how soil moisture contributes to corn growth and affects yield.
-Soil moisture is influenced by temperature, precipitation, and other environmental factors and directly affects crop growth.
-As a key indicator of crop water supply, low soil moisture causes water stress, which limits corn growth and reduces yields (Unganai and Kogan, 1998).
-Unlike most RS data, such as NDVI, which only reflects the growth status of plants at a specific moment, soil moisture has a cumulative impact on yield over time.
-Nonetheless, a significant challenge for ML models is the limited availability of soil moisture data.
-For example, the widely used SPL4SMGP.007 SMAP soil moisture dataset (Reichle et al., 2022) only covers the period from 2015 onward, which limits the ability of ML models to learn from extensive historical data and impacts model performance.
-Therefore, we present a Knowledge-Guided Machine Learning with Soil Moisture (KGML-SM) framework that explicitly incorporated the soil moisture impact in the ML model and provided interpretability.
-The model consisted of two major components: a Weather-to-Soil (W2S) encoder modeling the relationship between weather inputs and soil moisture outputs, and an attention module (Vaswani, 2017) integrating features from multiple sources and dynamically weighting their influences on corn yield across time and space.
-To train this model, we adopted a two-step process.
-First, we used the APSIM model to generate a simulated field-level dataset for pretraining.
-Then, the model was finetuned using a county-level dataset from Google Earth Engine (GEE) (Gorelick et al., 2017) and United States Department of Agriculture (USDA) National Agricultural Statistics Service (NASS).
-Furthermore, we conducted research into the relationship between soil moisture and corn yield prediction in drought-affected areas, where drought often reduced yields and caused overestimation (Cooper et al., 2014).
-Researchers have long been interested in studying the relationship between soil moisture, drought, and corn yield prediction (Ines et al., 2013; Bushong et al., 2016; Mladenova et al., 2017).
-However, these studies either had a narrow scope, or did not utilize this connection for model optimization.
-we designed a drought-aware loss function to further enhance model performance under drought conditions and mitigate overestimation problem.
-Finally, based on the relationship between soil moisture and corn yield prediction, we provided explanations for the prediction errors of the KGML-SM model and offered directions for future model optimization.
-Our research aimed to improve the accuracy and interpretability of corn yield predictions, ultimately contributing to more resilient agricultural systems under extreme weather variability.
-Our main contributions are as follows:
-  • 1.
-    This paper introduced the KGML-SM framework. This framework integrated process-based and ML models for corn yield prediction while explicitly incorporating the influence of soil moisture.
-  • 2.
-    A drought-aware loss function was designed. This function enhanced model performance under drought conditions and mitigated overestimation.
-  • 3.
-    Based on the relationship between drought, soil moisture, and corn yield prediction, we provided interpretability for the prediction errors of the KGML-SM model and offered directions for future model optimization.
+(2024) linked image-derived plant traits to a knowledge-guided S-shaped growth curve to forecast fruit growth, yield, and maturity.
+Recently, Miranda et al.
+(2024) developed a physics-informed recurrent neural network for yield loss forecasting, where crop water use (ETa) and drought sensitivity (Ky) were estimated under physical constraints.
+Despite these advances, no study has explicitly embedded soil moisture into the ML model structure.
+Soil moisture is central to crop production because it mediates the effects of weather on plant growth and yield formation (Unganai and Kogan, 1998).
+Unlike vegetation indices (VIs) which capture plant status at a single moment, soil moisture reflects the cumulative effects of precipitation (PPT) and temperature over time, making it a more process-oriented indicator of water availability.
+Embedding soil moisture explicitly in a KGML model not only provides a mechanistic link between weather and yield but also creates diagnostic interpretability, allowing prediction errors to be traced to potential misrepresentation of soil water dynamics.
+To effectively exploit soil moisture for yield prediction, reliable large-scale observations are needed.
+Several soil moisture products have been developed over the past decades and provide valuable long-term global records.
+The ASCAT product offers C-band scatterometer-based soil moisture observations (Wagner et al., 2013).
+In addition to satellite scatterometer data, reanalysis products such as ERA5-Land (Muñoz-Sabater et al., 2021) and MERRA-2 (Gelaro et al., 2017) provide surface and root-zone soil moisture at global scales.
+Furthermore, the ESA product delivers more than 30 years of consistent global daily soil moisture records by merging multiple satellite sensors (Dorigo et al., 2017).
+Building on these advances, the Soil Moisture Active Passive (SMAP) mission (Reichle et al., 2022; Entekhabi et al., 2010) delivers high-accuracy global soil moisture observations with a revisit cycle of two to three days.
+Compared with earlier C- or X-band missions, SMAP’s L-band measurements penetrate deeper into the soil and are less affected by vegetation and atmospheric disturbances, making it one of the most reliable datasets for agricultural applications.
+Previous studies consistently demonstrate a strong link between soil moisture, corn yield, and drought impacts.
+For instance, Ines et al.
+(2013a) showed that assimilating remotely sensed soil moisture and VIs into a crop simulation model substantially improved corn yield prediction.
+In addition, Rigden et al.
+(2020) reported that the combined influence of soil moisture and atmospheric evaporative demand explained most of the observed interannual variability in U.S.
+corn yields, underscoring the need to jointly consider soil and atmospheric drivers.
+At the field scale, de Lara et al.
+(2019) found that soil water content in deeper layers during reproductive stages showed the strongest correlation with corn yield, and that integrating high-resolution VIs improved precision irrigation decisions.
+Other studies have further explored related directions, such as improving retrieval accuracy by combining ASCAT and SMAP observations (Ma et al., 2024), developing cumulative drought indices (CDI) from process-based models to enhance subfield yield predictions (Shuai and Basso, 2022), or applying ML and DL approaches for soil moisture prediction and yield modeling (Teshome et al., 2024; Kang et al., 2020; Yu et al., 2020).
+Collectively, these works (Table 1) underscore the central role of soil moisture in corn production, but most treat it only as an auxiliary covariate or prediction target.
+In contrast, our KGML-SM framework embeds soil moisture explicitly as an intermediate variable, creating a mechanistic link between weather and yield and enabling diagnostic interpretability.
+Building on this design, KGML-SM also introduces a drought-aware loss function to penalize overestimation under water-limited conditions, thereby improving robustness.
+Unlike approaches that simply add soil moisture as another input feature or prediction, our framework provides a structured way to trace yield errors back to soil–weather interactions.
+Moreover, the proposed modules are general: embedding soil moisture as an intermediate variable can be integrated into other crop yield models to enhance interpretability, while the drought-aware loss can be applied more broadly in drought-prone regions.
+Table 1: Summary of KGML applications in crops and soil moisture–related studies in context of this work.
+Reference Content Method
+(He et al., 2023) Using physics-guided neural networks for time-aware fairness in crop yield prediction LSTM with attention; physics-guided reweighting; fairness refinement
+(Yang et al., 2023) Integrating process-based surrogates with multi-source data assimilation for agroecosystem prediction GRU surrogate; EnKF; data fusion
+(Yang et al., 2024) Predicting greenhouse strawberry growth trajectory and yield via knowledge-guided computer vision Faster R-CNN; DenseNet-based trait extraction; S-shaped growth-curve modeling
+(Lobell et al., 2015) Mapping large-area crop yield from satellites via pseudo-observations and calibration Pseudo-observation conversion; regression calibration
+(Shuai and Basso, 2022) Improving subfield corn yield prediction by including in-season water deficit with RS VIs SALUS-simulated CDI by yield-stability zones; Random Forest (RF) per tile-year and a two-window composite
+(Mena et al., 2025) Fusing multi-modal remote-sensing data for optimal subfield yield prediction Multi-modal encoders; gated fusion
+(Miranda et al., 2024) Exploring physics-informed neural networks for crop yield loss forecasting PINN with physics-based constraints
+(Ines et al., 2013a) Assimilating RS soil moisture and vegetation into a crop model for corn yield prediction EnKF with DSSAT-CSM (Maize); sequential SM/LAI assimilation
+(Ma et al., 2024) Integrating ASCAT and SMAP for global surface soil moisture retrieval using ML Model comparison: RF / LSTM / SVM / CNN
+(Rigden et al., 2020) Combining soil moisture and atmospheric evaporative demand to predict US corn yields Statistical modeling with process-based analysis
+(de Lara et al., 2019) Relating soil water content and high-resolution imagery for precision irrigation in corn Linear regression with VIs coupling
+(Teshome et al., 2024) Benchmarking DL vs. ML models for soil moisture prediction under irrigation treatments
+(Kang et al., 2020) Comparing environmental variables and ML algorithms for corn yield in the US Midwest Comparative ML assessment
+(Yu et al., 2020) Predicting multi-depth soil water content during summer corn for irrigation planning Residual bidirectional LSTM
+This study The first study to embed soil moisture explicitly as an intermediate variable within the ML architecture for county-level corn yield prediction; mitigating overestimation under drought area Knowledge-guided pretraining; W2S encoder; drought-aware loss function
+In summary, this study introduces a KGML-SM framework to address the limitations of existing process-based and ML approaches for county-level corn yield prediction.
+The framework explicitly embeds soil moisture as an intermediate variable through a Weather-to-Soil (W2S) encoder and an attention mechanism (Vaswani, 2017), creating a mechanistic link between weather drivers and crop outcomes.
+In addition, a drought-aware loss function penalizes overestimation under water-limited conditions, thereby improving model robustness during drought years.
+By jointly enhancing accuracy, interpretability, and robustness, KGML-SM improves yield prediction.
+To our knowledge, this is the first framework to explicitly embed soil moisture into a KGML model for county-level prediction, offering both methodological innovation and practical relevance for supporting resilient agricultural management under increasing climate variability.
 2 Data acquisition
@@
 The first dataset was a field-level dataset generated using APSIM (McCown et al., 1996) and used for pretraining.
-The second dataset was a county-level dataset derived from GEE (Gorelick et al., 2017) and USDA NASS and used for finetuning.
+The second dataset was a county-level dataset derived from Google Earth Engine (GEE) (Gorelick et al., 2017) and USDA NASS (USDA, 2020), and used for finetuning.
 The workflow was divided into three main steps: simulation, pretraining, and finetuning.
-A summary of all public datasets used in this study is provided in Table 1.
 In this section, we first introduce the study area (Sec.
@@
 2.3).
-Table 1: Public datasets used in this study. a, b, and c indicate different purposes: a for APSIM simulation, b for model pretraining, and c for model finetuning.
-Dataset Use Description References
-Iowa Environmental Mesonet a, b In-situ weather data (Herzmann et al., 2004)
-PRISM c Gridded weather data (Daly et al., 2015, 2008)
-MODIS MCD43A4 c Vegetation indices (Schaaf and Wang, 2015)
-MODIS MCD18A1.061 c Radiation data (Wang, 2021)
-USDA-CDL c Crop data layer (USDA-NASS, 2017)
-SMAP c Soil moisture data (Reichle et al., 2022)
-USDA NASS yield c County-level yield (USDA, 2020)
 2.1 Study area
@@
 Corn Belt.
-Twelve states are selected as our study area, including North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan.
+Twelve states were selected as our study area, including North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan.
 These states are crucial agricultural states in the U.S., known for their significant contributions to corn production.
@@
 1) for these twelve states and considered them well suited for corn yield prediction research.
+For brevity, state names were referred to by their standard abbreviations (e.g., Wisconsin as WI) in the following sections.
 Figure 1: The 5-year average county-level corn yield map in selected states.
@@
 In this section, we explain how APSIM was used to generate a field-level simulated dataset for model pretraining.
-We used data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides U.S.
-agricultural and environmental data.
-We extracted all available station data from the IEM, which included data from thousands of stations across 12 states from 1980 to 2023 (Table 2).
-Table 2: The number of stations selected for each state.
+We used daily weather data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides agricultural and environmental observations across the United States.
+Specifically, we extracted all available station-based records, which include variables such as temperature, PPT, wind, and solar radiation, together with the latitude and longitude of each station.
+This APSIM field-level dataset covers several thousand stations across 12 states from 1980 to 2023 (Table 2).
+Table 2: The number of stations used for APSIM field-level dataset and the average number of counties used for GEE county-level dataset in each state.
 State IL IN IA KS MI MN MO NE ND OH SD WI
-Station number 120 89 113 150 140 140 130 128 79 100 114 147
+Station number
+(APSIM field-level dataset)
+120 89 113 150 140 140 130 128 79 100 114 147
+County number
+(GEE county-level dataset)
+91 74 87 74 57 64 74 69 48 72 39 62
 2.2.1 Input and output of APSIM simulation
-The APSIM model simulates corn yield based on four weather data inputs: maximum and minimum temperature (Tmax and Tmin), precipitation (PPT), and radiation (Radn).
-The simulation also requires the station’s latitude, longitude, and year.
-Some management parameters are also necessary, including start and end of sowing window, plant population, fertilizer amount, and initial soil water, etc (Sec.
-3.3.1).
+The APSIM model simulates corn yield based on four weather data inputs: maximum and minimum temperature (Tmax and Tmin), PPT, and radiation (Radn).
 Tmax and Tmin influence the rate of plant development, and affect processes like photosynthesis and respiration.
@@
 These weather data of IEM stations are available from the IEM website (Herzmann et al., 2004).
+The simulation also requires the station’s latitude, longitude, and year.
+Some management parameters are also necessary, such as the start and end of the sowing window, plant population, fertilizer amount, and initial soil water, which were described in detail in Sec.
+3.3.1.
 The APSIM model takes weather data as input to simulate root zone soil moisture (SM_rootzone), surface soil moisture (SM_surface), and corn yield.
@@
 2.2.2 Variables summary in field-level dataset for pretraining
-We calculated a five-year historical average yield based on the simulated corn yield to establish a local baseline for model accuracy improvement.
-We also included the prediction year and location (longitude and latitude) of IEM stations in the field-level dataset.
+We adopted a five-year averaging strategy consistent with prior studies (Ma et al., 2021a; Wang et al., 2025), which enhances robustness by mitigating short-term variability while retaining long-term soil–climate signals.
+The resulting five-year historical average of simulated corn yield (Historical avg.
+yield) was used as a local baseline to improve model accuracy.
+We also included the prediction year and location (Lat&Long) of IEM stations in the field-level dataset.
 Because soil properties and topography vary across space and time, year and location can help the ML model capture this variation.
-Then we combined the weather data, simulated soil moisture and corn yield, along with other variables to construct the APSIM field-level dataset (Table 3).
-Table 3: Summary of variables comprising the APSIM field-level dataset, including inputs and simulated outputs
-Category Variables
-Weather data Radn (MJ/m2) (input)
-Tmax (${}^{\circ}\mathrm{C}$) (input)
-Tmin (${}^{\circ}\mathrm{C}$) (input)
-PPT (mm) (input)
-Soil moisture SM_surface (simulated)
-SM_rootzone (simulated)
-Corn yield Corn yield (t/ha) (simulated)
-Others Prediction year (input)
-Location (Latitude and longitude) (input)
-Historical average yield (t/ha) (simulated)
+We then combined the weather data, simulated soil moisture and corn yield, along with other variables to construct the APSIM field-level dataset (Table 3).
+VIs were not included in the pretraining dataset because the IEM stations, although co-located with RS imagery, are not necessarily situated in cornfields, making such signals unreliable for yield modeling.
+To avoid introducing noise, VIs were set to zero during pretraining, while in the finetuning stage, actual RS-based VIs were incorporated at the county level where crop type and coverage are reliably defined.
+Table 3: Summary of variables in the APSIM field-level dataset and the GEE county-level dataset
+Category Variables APSIM field-level dataset GEE county-level dataset
+Unit Type Source Unit
+Spatial
+resolution
+Source
+Weather data Radn MJ/m2 Input IEM W/m2 4 km PRISM
+Tmax C Input IEM C 4 km PRISM
+Tmin C Input IEM C 4 km PRISM
+PPT $m$ Input IEM $m$ 4 km PRISM
+Soil moisture SM_surface Simulated APSIM 9 km SMAP
+SM_rootzone Simulated APSIM 9 km SMAP
+Corn yield Yield t/ha Simulated APSIM t/ha USDA NASS
+VIs GCVI 500 m MODIS
+EVI 500 m MODIS
+NDWI 500 m MODIS
+NDVI 500 m MODIS
+Others Prediction year Input IEM NASS
+Location (Lat&Long) Lat&Long Input IEM Lat&Long USDA NASS
+Historical avg. yield t/ha Simulated APSIM t/ha USDA NASS
 2.3 GEE county-level dataset
-In this section, we introduce how to construct a county-level dataset for finetuning.
+In this section, the construction of a county-level dataset for finetuning is described.
 Details of the studied counties are provided in Table 2.
-This dataset includes all the variable types in the APSIM field-level dataset (Table 3), including weather, soil moisture, and other features, but from different sources.
-To enhance ML model performance with RS data, we also included four Vegetation Indices (VIs) from satellite imagery: Green Chlorophyll Index (GCI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI).
-Table 4: The average number of counties selected for each state in GEE county-level dataset.
-State IL IN IA KS MI MN MO NE ND OH SD WI
-County number 91 74 87 74 57 64 74 69 48 72 39 62
+This dataset contained all variable types from the APSIM field-level dataset (Table 3), including weather, soil moisture, and other features, but here they were obtained from RS and reanalysis products available in GEE.
+To enhance our model performance with RS data, four VIs from satellite imagery were also included: Green Chlorophyll Index (GCVI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI).
+Detailed descriptions of these data sources and processing steps are provided in Sec.
+2.3.1.
+Here, we only introduce the variables and data sources of the dataset; the processing methodology is described in Sec.
+3.4.
 2.3.1 Vegetation indices
-The Terra and Aqua combined Moderate Resolution Imaging Spectroradiometer (MODIS) Land Cover Climate Modeling Grid Version 6 data product (MCD12Q1 v6) (Schaaf and Wang, 2015) is a vital source of satellite-derived data with 500 m resolution, providing consistent, high-quality observations of the Earth’s surface.
-MODIS MCD12Q1 v6 captures data in multiple spectral bands and offers various products like VIs.
-These indices are essential for monitoring vegetation health, biomass, water content, and chlorophyll levels across different regions and time scales.
-They help improve yield prediction by indicating vegetation growth and crop health.
-The Green Chlorophyll Index (GCI) (Gitelson et al., 2005) is used to estimate the chlorophyll content in plants, which directly relates to plant health and productivity.
-It is calculated using the near-infrared (NIR) and green spectral bands.
-GCI is valuable for assessing vegetation health, especially for monitoring crop growth and detecting areas with nutrient deficiencies.
-High chlorophyll content typically indicates a healthy, productive plant, making GCI an essential tool in agricultural management.
-GCI is defined as:
-$GCI=\frac{NIR}{Green}-1$ (1)
-where $NIR$ represents the surface reflectance in the Near Infrared band, and $Green$ represents the surface reflectance in the Green band.
-The Enhanced Vegetation Index (EVI) (Huete et al., 2002) is designed to optimize the vegetation signal by enhancing sensitivity in high biomass areas and reducing atmospheric and canopy background noise.
-It uses the red, blue, and NIR spectral bands.
-EVI is particularly useful in regions with dense vegetation, where NDVI might saturate.
-It provides a more accurate representation of vegetation conditions, especially in areas with high leaf area index, making it ideal for monitoring forests and agricultural lands.
-EVI is defined as:
-$EVI=2.5\times\frac{NIR-Red}{NIR+6\times Red-7.5\times Blue+1}$ (2)
-where $Red$ and $Blue$ represent the surface reflectances of the Red band and Blue band, respectively.
-The Normalized Difference Water Index (NDWI) (Gao, 1996) measures the moisture content in vegetation and soil by using the NIR and shortwave infrared (SWIR) bands.
-NDWI is crucial for monitoring water stress in plants, assessing drought conditions, and detecting changes in water bodies.
-It aids in managing irrigation practices and understanding the impact of water availability on crop yields.
-NDWI is defined as:
-$NDWI=\frac{NIR-SWIR}{NIR+SWIR}$ (3)
-where $SWIR$ represents the shortwave infrared band.
-The Normalized Difference Vegetation Index (NDVI) (KRIEGLER, 1969) is one of the most widely used indices for measuring vegetation greenness and biomass.
-It is calculated using the NIR and red spectral bands.
-NDVI is a standard tool for monitoring plant growth, assessing vegetation cover, and detecting changes in biomass over time.
-It is particularly useful for tracking seasonal variations in vegetation and estimating crop yields.
-NDVI is defined as:
-$NDVI=\frac{NIR-RED}{NIR+RED}$ (4)
+The Terra and Aqua combined Moderate Resolution Imaging Spectroradiometer (MODIS) Land Cover Climate Modeling Grid Version 6 product (MCD12Q1 v6) (Schaaf and Wang, 2015) provides satellite-derived data at 500 m resolution and offers consistent, high-quality observations of the Earth’s surface.
+MODIS captures information across multiple spectral bands and enables the derivation of various VIs.
+These indices have been widely used to monitor vegetation health, biomass, water content, and chlorophyll levels across different regions and time scales, thereby improving yield prediction by reflecting crop growth and condition.
+The VIs used in this study are summarized in Table 4 along with their equations and main applications.
+Table 4: Summary of VIs used in this study, including definitions and applications.
+Abbreviations: NIRNIR (near-infrared reflectance); RedRed (red reflectance); BlueBlue (blue reflectance); GreenGreen (green reflectance); SWIRSWIR (shortwave infrared reflectance).
+Band specifications for MODIS are as follows: Red – Band 1 (620–670 nm); NIR – Band 2 (841–876 nm); Blue – Band 3 (459–479 nm); Green – Band 4 (545–565 nm); SWIR – Band 6 (1628–1652 nm).
+Vegetation Index Citation Equation Application
+GCVI (Gitelson et al., 2005) $GCVI=\frac{NIR}{Green}-1$ Chlorophyll content; Crop health; Nutrient status
+EVI (Huete et al., 2002) $EVI=\frac{2.5\times(NIR-Red)}{NIR+6\times Red-7.5\times Blue+1}$ Dense canopy; High biomass; Noise reduction
+NDWI (Gao, 1996) $NDWI=\frac{NIR-SWIR}{NIR+SWIR}$ Soil/vegetation moisture; Drought monitoring; Irrigation management
+NDVI (KRIEGLER, 1969) $NDVI=\frac{NIR-Red}{NIR+Red}$ Vegetation greenness; Biomass; Yield prediction
 2.3.2 Weather data
@@
 The MCD18A1 Version 6.1 (Wang, 2021) is a MODIS Terra and Aqua combined Downward Shortwave Radiation gridded Level 3 product.
-The reliable radiation data (Radn) is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours.
+The reliable radiation data is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours.
 Downward Shortwave Radiation is incident solar radiation over land surfaces in the shortwave spectrum (300-4,000 nanometers) and is an important variable in land-surface models that address a variety of scientific and applied issues.
 2.3.3 Soil moisture
-The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Reichle et al., 2022) plays a critical role in our research on the relationship between drought and corn yield prediction.
+The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Reichle et al., 2022; Entekhabi et al., 2010) plays a critical role in our research on the relationship between drought and corn yield prediction.
 By providing detailed measurements of soil moisture at both the surface (0–5 cm) and root zone levels (0–100 cm), SMAP level 4 data allows us to assess the availability of water in the soil, a key factor influencing crop growth and resilience during drought conditions.
@@
 All the variables in GEE county-level dataset are listed in Table 3.
-Table 5: Summary of variables in GEE county-level dataset
-Category Variables Spatial resolution Source
-Vegetation index Green Chlorophyll Index (GCI) 500m MODIS
-Enhanced Vegetation Index (EVI)
-Normalized Difference Water Index (NDWI)
-Normalized Difference Vegetation Index (NDVI)
-Weather data Radn (W/$m^{2}$)
-Tmax (${}^{\circ}\mathrm{C}$) 4km PRISM
-Tmin (${}^{\circ}\mathrm{C}$)
-PPT (mm)
-Soil moisture SM_surface 9km SMAP
-SM_rootzone
-Corn yield USDA NASS corn yield (t/ha) N/A USDA NASS
-Others Prediction year
-Location (Latitude and longitude)
-Historical average yield (t/ha)
 3 Methodology
@@
 3.3) and the GEE county-level dataset for finetuning and testing (Sec.
-3.4); finally, we introduce how the KGML-SM model is trained and used for prediction (Sec.
-3.5).
+3.4); then, we introduce how the KGML-SM model is trained and used for prediction (Sec.
+3.5); after that, we describe the experimental setup (Sec.
+3.6) and the statistical analysis of drought, soil moisture, and corn yield (Sec.
+3.7).
 3.1 Problem formulation
@@
 The W2S encoder is a module designed to model the influence of weather conditions on soil moisture.
-By capturing the relationship between weather and soil moisture, the W2S encoder improves the model’s understanding of soil dynamics and the impact of weather variability on soil conditions and corn yield.
+By capturing the statistical relationship between weather variables and soil moisture, the W2S encoder improves the representation of soil moisture dynamics at the county scale, which in turn supports more accurate yield prediction under varying weather conditions.
 The W2S encoder employs a U-Net-based encoder-decoder architecture, which consists of an encoder, a decoder, and a fully connected layer (Goodfellow et al., 2016) for feature transformation.
@@
 3.3.1 APSIM management parameters
-Simulation methodology depends on the study’s scale.
-For field-level areas, management information is obtained directly from farmers as input.
-Other unknown parameters are adjusted to optimize the model (Shahhosseini et al., 2021).
-For county-level areas, identifying representative fields is common practice.
-Their management parameters are averaged, and then applied to represent the entire county (Puntel et al., 2016).
+Simulation methodology depended on the study’s scale.
+For field-level areas, management information was obtained directly from farmers as input.
+Other parameters—such as sowing density, fertilizer amount, initial soil water content, and selected cultivar coefficients—were calibrated to optimize model performance (Zhen et al., 2022).
+For county-level areas, identifying representative fields was common practice.
+Their management parameters were averaged, and then applied to represent the entire county (Puntel et al., 2016).
 Calibrating each county individually became impractical for very large study areas like our U.S.
 Corn Belt simulation.
-In such cases, the typical approach defines a parameter range based on empirical and statistical data.
-These parameters are then randomly combined within specified ranges, simulating yield across all counties (Lobell et al., 2015).
+In such cases, the typical approach defined a parameter range based on empirical and statistical data.
+These parameters were then randomly combined within specified ranges, simulating yield across all counties (Lobell et al., 2015).
 Given the vast area of cornfields in our study, it was not feasible to calibrate these parameters for every individual station.
-Therefore, we established a general parameter range encompassing all possible values across different fields.
+Rather than calibrating parameters to specific local conditions, the goal of generating simulated data was to expand the range of training samples, thereby improving the model’s capacity to generalize across diverse environments.
+Although this dataset may appear relatively coarse due to the lack of calibration, such diversity is in fact desirable for pretraining, as it exposes the model to a wider spectrum of conditions and enhances its ability to generalize.
+Importantly, our focus is not on accurately predicting yield at these station sites, but rather on ensuring robust county-level performance after finetuning with observed data.
+At this stage, the simulated dataset primarily serves to expose the model to general patterns of corn growth, while subsequent county-level calibration ensures predictive accuracy at the target scale.
+By defining parameters within broad but reasonable ranges, we were able to generate a diverse set of simulated samples, which serves as the foundation for the parameter specifications described in the following section.
+Therefore, we established a general parameter range encompassing all possible values across different fields (Table 5).
 This approach ensured model applicability across a wide range of conditions, maintaining reasonable prediction accuracy.
-Some reasonable adjustment ranges for key management parameters were collected from prior research and USDA statistics (Table 6).
+Some reasonable adjustment ranges for key management parameters were collected from prior research and USDA statistics.
 Sowing density determines the number of plants per unit area, directly affecting competition for resources such as light, water, and nutrients.
@@
 The most commonly used nitrogen fertilizers for corn production in North America are anhydrous ammonia, urea, and urea-ammonium nitrate solutions (Herzmann et al., 2004).
-The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N), values of 200, 250, and 300 were used in the simulation.
+The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N).
+Values of 200, 250, and 300 were used in the simulation.
 Initial soil water content is important for establishing the starting conditions for the model’s simulation of soil moisture dynamics throughout the growing season.
@@
 3.3.2 Optimizing APSIM field-level dataset based on soil moisture
-APSIM is not accurate for large-scale corn yield simulation without precise management adjustments, thus requiring optimization of the APSIM field-level pretrain dataset.
-Soil moisture was proposed as a benchmark for filtering high-quality data.
-First, soil moisture dynamics are relatively simple, strongly correlated with precipitation.
-In contrast, corn yield is influenced by multiple factors and complex growth processes, hindering accurate modeling in process-based model.
-Second, simulation processes inherently exhibit cumulative error.
-Due to the relative simplicity of soil moisture modeling, this error is comparatively smaller.
-Therefore, the APSIM field-level dataset was optimized based on the quality of soil moisture data.
-Specifically, an LR model was trained on the GEE county-level dataset.
-It utilized four weather inputs (Tmax, Tmin, PPT, Radn) to predict soil moisture output.
-The trained model subsequently predicted soil moisture for the APSIM field-level dataset using its weather data.
-This prediction was compared against the simulated soil moisture.
-Simulated soil moisture with a mean squared error (MSE) exceeding 0.5 were discarded ensuring dataset quality.
+APSIM was not accurate for large-scale corn yield simulation without precise management adjustments, making it necessary to filter the APSIM field-level dataset before using it for pretraining.
+Although the goal was to construct a diverse dataset, retaining highly unrealistic samples would only introduce noise and bias the model away from biologically plausible relationships.
+Filtering was therefore essential to ensure that the dataset remained both diverse and reliable.
+Soil moisture was selected as the benchmark variable because it provides a balance between data availability and process relevance.
+Although it is influenced by multiple factors such as PPT, soil texture, and topography, detailed station-level information on these factors was unavailable across the study region.
+In contrast, weather data were consistently available and strongly correlated with soil moisture, making it a practical proxy for dataset screening.
+Moreover, cumulative simulation errors are typically smaller for soil moisture than for yield, which is affected by many interacting processes.
+Based on these considerations, we used soil moisture quality as the criterion to optimize the APSIM field-level dataset.
+Specifically, a Linear Regression (LR) model was trained on the GEE county-level dataset to predict soil moisture from four weather inputs (Tmax, Tmin, PPT, Radn).
+The trained model then predicted soil moisture predictions for the APSIM field-level dataset, which were compared against the APSIM-simulated values.
+To ensure data quality while retaining sufficient variability, samples with mean squared error (MSE) above 0.5 were discarded.
+This threshold was chosen as a balance: higher-error samples (e.g., $\geq$1) were unreliable compared to typical soil moisture RMSE values, yet a more stringent cutoff could have removed too much variability and reduced the generalization capacity of the pretraining dataset.
 3.4 Generating GEE county-level dataset for finetuning and testing
@@
 To focus on the corn portion of the acquired feature data from GEE, the cornfield areas had to be identified.
-The Cropland Data Layer (CDL) cropland mask (USDA-NASS, 2017) was utilized extracting feature data corresponding to cropland.
-The Cropland Data Layer is an annual raster-based dataset.
-It provides crop-specific land cover information produced by the USDA.
-This dataset has a 30-meter resolution.
-The CDL was employed to generate a cornfield mask in GEE, supporting our corn yield prediction study in twelve U.S.
+The Cropland Data Layer (CDL) cropland mask (USDA-NASS, 2017) was used to extract the corn class for each year from 2015 to 2023.
+The CDL is an annual raster-based dataset with a 30-meter resolution, providing crop-specific land cover information produced by the USDA.
+These annual cornfield masks were employed in GEE to support our corn yield prediction study in twelve U.S.
 Corn Belt states.
 3.4.2 Data preprocessing
-County-level features were aggregated into a single value every 16 days, representing agricultural conditions for that period.
-This process resulted in a data vector for each feature annually for each county.
-This aggregation method followed the experimental setup from previous research (Ma et al., 2021b; Wang et al., 2025).
+For each county, pixel-level feature values from MODIS and SMAP products were first averaged spatially to obtain county-level values.
+These county-level values were then aggregated into a single representative mean value every 16 days during the growing season (April–October), following the experimental setup of previous research (Ma et al., 2021b; Wang et al., 2025).
+This temporal aggregation balanced data availability with noise reduction and provided a consistent seasonal trajectory of vegetation and environmental conditions for each county.
+We noted that MODIS standard products included atmospheric correction and provided quality assurance layers that flag cloud-affected or low-quality pixels.
+In this study, we did not apply additional cloud filtering beyond the standard MODIS product processing; instead, the 16-day temporal aggregation helps to smooth residual noise caused by occasional cloud contamination.
+This procedure resulted in an annual time-series vector of features for each county, which served as input for model training and evaluation.
 3.5 Developing the KGML-SM framework
@@
 3.5.2 Drought-aware yield prediction loss function
-Soil moisture is widely recognized to affect crop yield (Ines et al., 2013; Bushong et al., 2016) and numerous studies explored this relationship (Mladenova et al., 2017; Vergopolan et al., 2021; Pignotti et al., 2023).
+Soil moisture is widely recognized to affect crop yield (Ines et al., 2013b; Bushong et al., 2016) and numerous studies explored this relationship (Mladenova et al., 2017; Vergopolan et al., 2021; Pignotti et al., 2023).
 However most prior studies lacked an elegant quantitative approach addressing this issue.
 This study introduces a loss function adjusting predicted corn yield based on varying soil moisture levels.
-The final yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation. It was formulated as:
+The final objective function jointly optimizes both components and is defined as:
+$\mathcal{L}_{total}=\mathcal{L}_{SM}+\mathcal{L}_{yield}$ (8)
+The yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation. It was formulated as:
 $\mathcal{L}_{\text{yield}}=\frac{1}{N}\sum_{i=1}^{N}d_{i}\left[(y_{i}-\hat{y}_{i})^{2}+\lambda\max(0,\hat{y}_{i}-y_{i})^{2}\right]$ (9)
@@
 $d_{i}=\frac{1}{\bar{s}_{i}+\varepsilon}$ (10)
-where $\bar{s}_{i}$ is the average soil moisture over the growing season for sample $i$, and $\varepsilon$ is a small constant to prevent numerical instability.
-We set $\varepsilon=1$ in our experiment.
+where $\bar{s}_{i}$ is the average soil moisture during the growing season (April–October) for sample $i$, calculated from the input soil moisture dataset (APSIM-simulated data for pretraining, and GEE/SMAP soil moisture for finetuning).
+$\varepsilon$ is a small constant to prevent numerical instability, which we set to 1 in our experiment.
 Since soil moisture plays a critical role in crop growth and yield formation, the loss function assigned a higher penalty to drier conditions, encouraging the model to be more responsive to soil moisture variations.
@@
 By applying a stronger penalty to overestimated yields, the model is encouraged to be more conservative, reducing the risk of unrealistic predictions.
-We set $\lambda=2$ in our experiment.
+The value of hyperparameter $\lambda$ was selected based on validation performance, and we set $\lambda=2$ in our experiment (Sec.
+4.2).
 Specifically, when $\hat{y}_{i}>y_{i}$, the predicted error is scaled by a penalty factor $\lambda$, amplifying the loss in these cases.
@@
 This process involved initially partitioning the data into training, validation, and test sets, followed by predicting corn yield for the target years.
+We ensured independent validation by adopting a temporal split strategy, where models were trained on preceding years and tested on the target year.
 Specific details regarding the partition and utilization of these datasets are elaborated upon in the experimental setup (Sec.
-4.1).
+3.6).
 The loss functions $\mathcal{L}_{\text{yield}}$ and $\mathcal{L}_{\text{SM}}$ were also optimized on the training and validation datasets.
-4 Experimental Results
 3.6 Experimental setup
@@
 The models were run on A100-SXM4-40GB and A100-SXM4-80GB GPUs.
-For training, we used the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of 0.001 and the ReduceLROnPlateau scheduler.
+For pretraining, we used a batch size of 64, a learning rate of 0.001, the Adam optimizer (Kingma and Ba, 2014), and the ReduceLROnPlateau scheduler with a patience of 5, with training stopped once the RMSE dropped below 1.
+For finetuning, we used a smaller batch size of 16 but kept the same learning rate, optimizer, and scheduler, and applied early stopping based on validation loss.
+The full configurations are summarized in Table 6.
+Table 6: Training configurations for pretraining and finetuning.
+Stage
+Batch
+size
+Learning
+rate
+Optimizer Scheduler Epochs Loss function Stopping criterion
+Pretrain 64 0.001 Adam
+ReduceLROnPlateau
+(patience=5)
+$\sim$50 MSE Training RMSE << 1
+Finetune 16 0.001 Adam
+ReduceLROnPlateau
+(patience=5)
+$\sim$30 MSE Early stopping based on validation loss
 Root mean square error ($RMSE$) and the coefficient of determination ($R^{2}$) were used to evaluate the performance of our model. The formulas for $RMSE$ and $R^{2}$ are:
@@
 3.7.1 Specifying drought area
-We investigated drought conditions in the United States (Drought.gov, 2024) during the corn growth season (June-September) (USDA, 2025) from 2019 to 2023 and present the corresponding maps (Fig.
-3).
-In 2019 (Fig.
-3(a)), drought conditions were predominantly observed in Iowa and Michigan.
-By 2020 (Fig.
-3(b)), Iowa experienced severe drought, while moderate conditions affected North Dakota, South Dakota, Nebraska, and Illinois.
-Conditions deteriorated further in 2021 (Fig.
-3(c)), with severe drought covering large parts of North Dakota, South Dakota, and Minnesota, while Nebraska and Iowa experienced moderate drought.
-In 2022 (Fig.
-3(d)), severe drought persisted in Nebraska and Kansas, with South Dakota and Iowa experiencing moderate drought.
-In 2023 (Fig.
-3(e)), severe drought was widespread across Nebraska, Kansas, Minnesota, Iowa, and Wisconsin, while Missouri faced moderate drought conditions.
+To objectively determine drought conditions, we used county‐level data from the U.S.
+Drought Monitor (USDM) (Drought.gov, 2024).
+For each Corn Belt state during the corn growing season (June–September; (USDA, 2025)), the USDM provides categorical drought classifications ranging from None to D4 (Exceptional Drought).
+Fig.
+3 shows the spatial distribution of these drought categories, providing a visual representation of drought intensity across years.
+To quantitatively assess drought severity, we further aggregated the USDM records of county areas under each drought category (None–D4).
+By summing the areas across drought categories, we calculated the proportion of each state’s total area that experienced drought in a given year.
+To complement the maps, Table 7 reports the state‐level proportions of drought‐affected counties, with values exceeding 30% highlighted in bold to emphasize years and regions with particularly widespread drought.
 Figure 3: Drought map of the U.S. for (1) June -(4) September from (a) 2019 to (e) 2023.
+Table 7: Proportion of counties affected by drought (categories D0–D4) in each of the 12 U.S.
+Corn Belt states (2018–2023), derived from U.S.
+Drought Monitor (USDM) county-level records during the corn growing season (June–September).
+Values indicate the fraction of state area under drought, with those exceeding 30% highlighted in bold.
+Year IA IL IN KS MI MN MO ND NE OH SD WI
+2019 17% 16% 15% 9% 15% 11% 1% 37% 1% 4% 0% 1%
+2020 64% 19% 49% 64% 22% 48% 19% 68% 55% 47% 59% 8%
+2021 89% 25% 18% 30% 65% 100% 8% 100% 76% 12% 100% 66%
+2022 62% 38% 47% 88% 30% 23% 61% 8% 97% 9% 73% 37%
+2023 100% 94% 73% 97% 72% 96% 95% 57% 93% 58% 61% 98%
 3.7.2 Soil moisture statistics
@@
 3.7.3 Drought impact on corn yield
-Finally, we examined the specific impact of drought on corn yield.
-We plotted the difference in corn yield for each year from 2019 to 2023 relative to the previous year’s yield (Fig.
-5), where negative values indicate a yield reduction.
-In 2019 (Fig.
-5(a)), drought-affected areas included Iowa and Michigan; however, significant yield reductions were observed only in Michigan.
-In 2020 (Fig.
-5(b)), Iowa faced drought along with a derecho storm (Hosseini et al., 2020), leading to noticeable yield losses.
-In 2021 (Fig.
-5(c)), drought impacted North Dakota, South Dakota, Minnesota, Nebraska, and Iowa, yet significant yield reductions occurred only in North Dakota and South Dakota.
-In 2022 (Fig.
-5(d)), drought-affected states included Nebraska, Kansas, South Dakota, and Iowa, but yield declines were primarily observed in Nebraska, Kansas, and South Dakota.
-In 2023 (Fig.
-5(e)), although Nebraska and Kansas continued to experience drought, corn yield did not decrease significantly compared to the previous year, likely due to the already reduced yield in 2022.
-However, Iowa and Missouri experienced some yield losses due to ongoing drought conditions.
-Thus, we concluded that drought consistently led to a decline in corn yield.
+Finally, we examined the interannual impact of drought on corn yield.
+Fig.
+5 shows the county-level yield differences between each year from 2019 to 2023 and the previous year, where negative values indicate yield reductions.
+To further quantify the extent of yield losses, we calculated for each state the proportion of counties experiencing a yield decline greater than 1 t/ha relative to the previous year.
+These results are summarized in Table 8, with values above 30% highlighted in bold to emphasize states and years with widespread yield reductions.
 Figure 5: The difference in corn yield between each year from (a) 2019 to (e) 2023 and the previous year, with negative values indicating a reduction in yield.
+Table 8: Proportion of counties with yield reductions greater than 1 t/ha relative to the previous year across 12 Corn Belt states (2019–2023), with those exceeding 30% highlighted in bold.
+Year IA IL IN KS MI MN MO ND NE OH SD WI
+2019 27% 78% 67% 15% 21% 36% 10% 50% 32% 54% 42% 30%
+2020 54% 1% 2% 20% 8% 5% 17% 33% 16% 7% 4% 7%
+2021 1% 0% 4% 16% 0% 50% 42% 87% 3% 0% 71% 0%
+2022 33% 1% 30% 76% 42% 3% 28% 0% 85% 26% 39% 12%
+2023 21% 23% 1% 15% 11% 30% 36% 17% 8% 0% 2% 17%
 3.7.4 Statistical summary of drought-affected states
 Based on the above statistical analysis, we summarized the drought-affected areas that experienced yield reductions in Table 9.
-In the experiment, we focused on these states that were affected by drought and experienced yield reductions.
+This table shows the states where drought and yield loss overlapped.
+In 2019, drought was limited, but many counties still had large yield reductions.
+This was mainly due to delayed planting and record prevented planting caused by heavy rainfall and flooding (American Farm Bureau Federation, 2019; Farm Service Agency, U.S.
+Department of Agriculture, 2019).
+In 2020, fewer yield losses were observed, likely because many states had already lost yield in 2019, leaving less room for further decline.
+In 2021 and 2022, drought-affected states also showed clear yield reductions, which matches the expected pattern.
+In 2023, drought was widespread across all states, but only a few states had obvious yield losses, because severe drought in 2022 had already caused large yield reductions.
 Table 9: Summary of obviously yield-reduced states within drought-affected areas.
-Year Area type Drought-affected states Obviously yield-reduced states
-2019 IA,MI MI
-2020 IA,ND,SD,NE,IL IA
-2021 ND,SD,MN,NE,IA ND,SD
-2022 NE,KS,SD,IA NE,KS,SD
-2023 NE,KS,MN,IA,WI,MO IA,MO
+Year Area type Drought-affected states Obviously yield-reduced states Intersection
+2019 ND IL,IN,MN,ND,NE,OH,SD,WI ND
+2020 ND,IA,KS,SD,NE,MN,IN,OH IA,ND IA,ND
+2021 MN,ND,SD,IA,NE,WI,MI MN,MO,ND,SD MN,ND,SD
+2022 NE,KS,SD,IA,MO,IN,IL,WI IA,IN,KS,MI,NE,SD IA,IN,KS,NE,SD
+2023 All the states MN,MO MN,MO
+4 Experimental Results
 4.1 Evaluation results
@@
 Table 10: Comparison with traditional ML models
-Year Method KGML-SM LR MLP RR RF
+Year Method KGML-SM LR MLP $R$ RF
 RMSE R2 RMSE R2 RMSE R2 RMSE R2 RMSE R2
@@
 To further demonstrate the contribution of each module in our KGML-SM model, we conducted a series of ablation studies.
-We began with a baseline model trained using only the GEE county-level dataset and the attention module (Att) (Wang et al., 2025).
-Next, we incorporated the APSIM field-level dataset pretraining to evaluate its impact on model performance (Att+sim).
-We then introduced the W2S encoder to assess its effectiveness in integrating soil moisture information (Att+sim+W2S).
-Finally, the impact of the drought-aware loss function was evaluated by comparing the performance of the KGML-SM model with the (Att+sim+W2S) model.
-We specifically compared the differences among models and their components in Table 9.
+We began with an attention-based baseline model without soil moisture inputs (Att w/o SM), which used all features except soil moisture.
+We then included soil moisture in the input features to form the Att model.
+Next, we incorporated the APSIM field-level dataset for pretraining (Att+sim), followed by the addition of the W2S encoder to integrate soil moisture dynamics (Att+sim+W2S).
+To further analyze the contribution of different loss components in our drought-aware loss $\mathcal{L}_{\text{yield}}=\frac{1}{N}\sum_{i=1}^{N}d_{i}\left[(y_{i}-\hat{y}_{i})^{2}+\lambda\max(0,\hat{y}_{i}-y_{i})^{2}\right]$, we first added the Soil Moisture Weighted (SMW) term, represented by $d_{i}$, which assigns greater weight to errors under low soil moisture conditions (Att+sim+W2S+SMW).
+We then introduced the Overestimation penalty (OE) term, $\lambda\max(0,\hat{y}_{i}-y_{i})^{2}$, which penalizes yield overestimation more strongly, particularly under drought conditions.
+Combining both SMW and OE yielded the full KGML-SM framework.
+A summary of the components of each ablation model is provided in Table 11.
 Table 11: Comparison of different methods and their components.
-Method Attention module Field-level data pretraining W2S encoder Drought-aware loss
+Method
+Attention
+module
+Soil
+moisture
+Field-level data
+pretraining
+W2S
+encoder
+SMW
+loss
+OE
+loss
+Att w/o SM
 Att
@@
 Att+sim+W2S
+Att+sim+W2S+SMW
 KGML-SM
@@
 This indicated that our APSIM field-level dataset effectively captured county-level data patterns, playing a crucial role in enabling the model to learn the relationship between corn yield and agricultural variables.
-Additionally, our drought-aware loss function significantly enhanced performance in most years, demonstrating that incorporating prior knowledge of the relationship between drought and corn yield can effectively guide the training and prediction of ML models.
-Table 10: Ablation study of different components in KGML-SM model
-Year Method KGML-SM Att+sim+W2S Att+sim Att
-RMSE R2 RMSE R2 RMSE R2 RMSE R2
-2019 0.964 0.741 0.974 0.732 1.087 0.715 1.268 0.57
-2020 0.98 0.792 0.981 0.783 1.003 0.77 1.011 0.77
-2021 1.104 0.836 1.127 0.811 1.143 0.814 1.195 0.808
-2022 1.085 0.837 1.119 0.802 1.144 0.802 1.315 0.779
-2023 1.071 0.807 1.101 0.805 1.114 0.781 1.201 0.766
-4.5 Spatial Visualization of model performance
+When comparing the baseline attention model without soil moisture (Att w/o SM) to the one including soil moisture inputs (Att), the performance difference was marginal across most years.
+This result suggests that simply adding raw soil moisture values provides limited benefits, highlighting the necessity of our subsequent modules to better exploit soil moisture information.
+Additionally, our drought-aware components also contributed to performance gains.
+While the SMW Loss improved performance in some drought-affected years, the effect was not consistent.
+This is because overestimation of yield can arise not only from drought but also from other natural hazards such as flooding, extreme temperatures, or pest outbreaks, which are not fully captured by soil moisture weighting alone.
+In contrast, the overestimation penalty directly constrains the model against overprediction across diverse adverse conditions, thereby providing more robust improvements.
+Table 12: Ablation study of different components in KGML-SM model (RMSE and R2 across years).
+Method Year 2019 2020 2021 2022 2023
+RMSE R2 RMSE R2 RMSE R2 RMSE R2 RMSE R2
+Att w/o SM 1.258 0.578 1.053 0.752 1.149 0.812 1.315 0.780 1.096 0.784
+Att 1.268 0.570 1.011 0.770 1.195 0.808 1.315 0.779 1.201 0.766
+Att+sim 1.087 0.715 1.003 0.770 1.143 0.814 1.144 0.802 1.114 0.781
+Att+sim+W2S 0.974 0.732 0.981 0.783 1.127 0.811 1.119 0.802 1.101 0.805
+Att+sim+W2S+SMW 1.097 0.711 1.054 0.763 1.112 0.832 1.074 0.821 1.043 0.812
+KGML-SM 0.964 0.741 0.980 0.792 1.104 0.836 1.085 0.837 1.071 0.807
+In DL, loss functions often include hyperparameters that control the relative importance of different error components (Goodfellow et al., 2016).
+In the drought-aware loss, the coefficient $\lambda$ determines the penalty strength applied to yield overestimation.
+Choosing an appropriate value of $\lambda$ is therefore critical: too small a value would fail to constrain overestimation effectively, whereas too large a value could distort the optimization and harm overall accuracy.
+To ensure a principled selection, we performed hyperparameter tuning on the validation set by testing multiple candidate values ($\lambda=0,1,2,5,10$) rather than assigning it arbitrarily.
+The results are summarized in Table 13.
+We found that $\lambda=2$ consistently offered the best trade-off between reducing RMSE and improving $R^{2}$, while both smaller and larger values led to inferior performance.
+Consequently, we adopted $\lambda=2$ as the default setting in the KGML-SM framework.
+Table 13: Ablation study on different values of the overestimation penalty coefficient $\lambda$ (RMSE and R2 for 2019–2023). The best $\lambda$ was selected based on validation performance.
+Year $\lambda=0$ $\lambda=1$ $\lambda=2$ $\lambda=5$ $\lambda=10$
+RMSE R2 RMSE R2 RMSE R2 RMSE R2 RMSE R2
+2019 1.097 0.711 0.982 0.752 0.964 0.741 1.032 0.730 1.115 0.695
+2020 1.054 0.763 0.976 0.775 0.980 0.792 1.021 0.749 1.109 0.710
+2021 1.112 0.832 1.122 0.819 1.104 0.836 1.166 0.785 1.235 0.745
+2022 1.074 0.821 1.108 0.829 1.085 0.837 1.153 0.792 1.223 0.755
+2023 1.043 0.812 1.062 0.801 1.071 0.807 1.184 0.776 1.246 0.740
+To evaluate the effect of filtering the simulated dataset (Sec.
+3.3.2), we compared KGML-SM trained with and without dataset filtering.
+As shown in Table 14, filtering substantially improved model performance across multiple years.
+This demonstrates that eliminating unreliable simulation samples is critical for constructing a robust pretraining dataset and enhancing generalization to county-level observations.
+Table 14: Effect of filtering the simulated dataset on model performance (RMSE and R2 for 2019–2023).
+Method 2019 2020 2021 2022 2023
+RMSE R2 RMSE R2 RMSE R2 RMSE R2 RMSE R2
+Unfiltered 1.120 0.705 1.010 0.768 1.198 0.807 1.305 0.775 1.225 0.760
+Filtered (KGML-SM) 0.964 0.741 0.980 0.792 1.104 0.836 1.085 0.837 1.071 0.807
+4.3 Prediction error spatialization and model bias
 In Fig.
-6, we plotted the error maps of the prediction results from 2019 to 2023 for the two best performed models: KGML-SM and RF model.
-In 2019 (Fig.
-6(a)), the lower part of Michigan showed noticeable overestimation in the RF model, whereas our KGML-SM model performed well.
-In 2020 (Fig.
-6(b)), although the derecho storm (Hosseini et al., 2020) in Iowa also had some impact on our KGML-SM model, we found that the RF model exhibited more severe overestimation.
-In 2021 (Fig.
-6(c)), overestimation in the upper-left region, particularly in North Dakota and South Dakota, was significantly reduced.
-In 2022 (Fig.
-6(d)), overestimation decreases notably in North Dakota, South Dakota, and Kansas, though the effect in Kansas was less pronounced.
-In 2023 (Fig.
-6(e)), Iowa showed a reduction in overestimation, while the effect in Missouri was less noticeable.
-Overall, our model shows a notable improvement in addressing the issue of overestimation in drought-affected areas.
+6, we present error maps of prediction results from 2019 to 2023 for the two best-performing models: KGML-SM and the RF baseline.
+Across all five years, KGML-SM consistently reduced overestimation compared with RF.
+In 2019, while overestimation in drought-affected areas of North Dakota was not substantially alleviated, several eastern states showed reductions.
+This is consistent with the fact that yield losses in 2019 were mainly due to delayed planting caused by prevented planting (American Farm Bureau Federation, 2019; Farm Service Agency, U.S.
+Department of Agriculture, 2019), rather than drought.
+In 2020, although the derecho storm (Hosseini et al., 2020) in Iowa also posed challenges for KGML-SM, the RF model exhibited much more severe overestimation, indicating that KGML-SM is effective not only under drought conditions but also in mitigating overestimation associated with other natural hazards.
+The improvements were particularly pronounced in 2021 and 2022, when RF showed widespread overestimation across the northwestern Corn Belt, corresponding to regions of drought-induced yield reduction in 2021, and across Kansas and Nebraska in 2022.
+In 2023, KGML-SM also reduced overestimation in Minnesota, further demonstrating its robustness.
+These results show the robustness of KGML-SM in mitigating systematic yield overestimation across diverse climatic conditions.
 Figure 6: The error map for (1) KGML-SM and (2) RF model from (a) 2019 to (e) 2023, with blue indicating overestimation.
-4.6 Understanding model bias and prediction-observation alignment
 To analyze model prediction performance, we generated scatter plots for the KGML-SM model and the RF model (Fig.
@@
 In 2021 (Fig.
-7(c)), the RF model exhibited prediction collapse, where certain observed values correspond to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations.
+7(c)), the RF model exhibited prediction collapse, where certain observed values corresponded to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations.
 In 2022 (Fig.
@@
 This comparison highlighted the advantage of the KGML-SM model in mitigating prediction collapse and improving overall robustness across different years.
-Figure 7: The scatter plot for (1) KGML-SM and (2) RF model from (a) 2019 to (e) 2023.
+Figure 7: Scatter plots of predicted versus observed county-level corn yields for KGML-SM (top row) and the RF model (bottom row) from 2019 to 2023.
+Panels (a1)–(e1) correspond to KGML-SM predictions for 2019–2023, while panels (a2)–(e2) correspond to RF predictions for the same years.
+The horizontal axis represents the reported (observed) yield (t/ha), and the vertical axis represents the model-predicted yield (t/ha).
+The color scale indicates the density of counties, and the dashed line shows the 1:1 reference line for comparison.
 5 Discussion
@@
 To enhance visualization, we normalized attention values within each year.
-Consequently, analysis should focus on attention trends across regions within the same year rather than comparing values across different years.
+Consequently, the analysis focused on attention trends across regions within the same year, while cross-year comparisons were not meaningful due to the normalization.
 Figure 8: The attention visualization of soil moisture on corn yield prediction in different regions.
@@
 These values were then normalized based on the maximum value in this year to ensure a consistent color scale, making the distribution more visually interpretable.
-In June, during the early growth period, the attention scores for soil moisture were generally lower across all five years, This might be because the temperature was not too high and drought conditions were not severe at this time, leading to a less pronounced correlation between soil moisture and corn yield.
+In June, during the early growth period, the attention scores for soil moisture were generally lower across all five years, This might have been because the temperature was not too high and drought conditions were not severe at this time, leading to a less pronounced correlation between soil moisture and corn yield.
 In July, attention scores increased, particularly in drought-affected areas, as soil moisture became more influential during active growth and vegetative stages.
 By August, attention scores peaked in drought-affected states, aligning with the critical reproductive phase of corn when adequate soil moisture was essential for kernel development.
-From 2019 to 2023, the attention distribution of soil moisture in corn yield prediction exhibited noticeable variations, aligning with the drought severity in different years.
+From 2019 to 2023, the attention distribution of soil moisture in corn yield prediction exhibited noticeable variations.
 In early 2019 (Fig.
-8(a1)), attention concentrated in Iowa.
-Subsequently (Fig.
-8(a3)), regions with higher attention values expanded gradually, though Iowa retained the highest attention, aligning with observed drought conditions that year.
+8(a1)), attention concentrated in Iowa, which does not align with the drought and yield reduction observed in North Dakota.
+An explanation is the widespread prevented planting reported that year, with the highest levels occurring in parts of eastern South Dakota, northwest Ohio, northeast Illinois, southwest Minnesota, and along the Mississippi (American Farm Bureau Federation, 2019; Farm Service Agency, U.S.
+Department of Agriculture, 2019), which disrupted normal planting schedules and shifted yield risks beyond drought-affected areas.
 During 2020 (Fig.
@@
 The reason for the incomplete consistency between the two could be that although drought was present in the region, it did not severely impact corn yield.
-Alternatively, even without drought in a particular area, other factors, such as the derecho storm, might have affected corn yield.
+Alternatively, even without drought in a particular area, other factors, such as the storm or flooding, might have affected corn yield.
 This analysis underscored the dynamic role of soil moisture in model prediction, with its importance intensifying during key growth stages and under severe drought conditions.
 5.2 Temporal role of soil moisture during the corn growth season
-To answer the second question and illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—every 16 days from June to September in drought-affected states (Fig.
+To answer the second question and illustrate the impact of soil moisture at different stages of the corn growing season, we visualized the attention of three feature types—VIs, weather, and soil moisture—at 16-day intervals from June to September in drought-affected states (Fig.
 9).
 We found that the VIs had the highest influence on the model around August, which aligned with findings from previous study (Johnson, 2014).
-In this period, VIs showed the strongest correlation with corn yield because this period corresponded to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield (Huang et al., 2014).
+In this period, VIs showed the strongest correlation with corn yield as this period corresponded to the vegetative and reproductive growth stages, during which crop health and biomass accumulation significantly impact final yield (Huang et al., 2014).
 High VIs in this timeframe indicate optimal chlorophyll content, canopy development, and water availability, which are critical for photosynthesis and grain formation (Wang et al., 2005).
@@
 Each point in this box plot represents a county, displaying the comparison of attention between red-marked drought-affected areas and blue-marked non-drought areas from 2019 to 2023 based on Table 9.
-Across all five years of attention values, drought-affected areas exhibited fewer outliers compared to non-drought areas, suggesting that the model’s attention to soil moisture is more stable in drought-affected regions.
+Across all five years, drought-affected areas exhibited fewer outliers compared to non-drought areas, suggesting that the model’s attention to soil moisture is more stable in drought-affected regions.
 Additionally, in all years, the median of soil moisture attention in drought-affected areas was consistently higher than in non-drought areas, indicating that soil moisture had a greater impact on the model in these areas.
-Figure 10: The boxplot of soil moisture attention in drought-affected areas (D--A) and non-drought areas (N--D) across all counties.
+Figure 10: The boxplot of soil moisture attention in drought-affected areas (D-A) and non-drought areas (N-D) across all counties.
 5.4 Interpreting corn yield prediction errors via soil moisture prediction anomalies
@@
 11).
-In our modeling process, we treated soil moisture as an intermediate variable to guide the corn yield output, so when corn yield prediction was inaccurate, the corresponding soil moisture might also exhibit anomalies.
-This analysis aims to conduct interpretability analysis of inaccurate corn yield predictions.
+In our modeling process, we treated soil moisture as an intermediate variable to guide the corn yield output, so when corn yield prediction was inaccurate, the corresponding soil moisture might also have exhibited anomalies.
+This analysis aims to interpret inaccurate corn yield predictions.
 In 2019 (Fig.
-11(a)), the drought-affected state Michigan exhibited relatively accurate corn yield predictions from the KGML-SM model; however, the soil moisture output was less accurate.
-This suggested that soil moisture was challenging to predict in drought-affected areas.
+11(a)), the absolute error map shows that corn yield predictions in North Dakota exhibited relatively large errors, while the corresponding soil moisture predictions remained accurate.
+This inconsistency suggests that yield errors were not primarily driven by soil moisture misrepresentation but were instead likely associated with large-scale prevented planting events in 2019 (American Farm Bureau Federation, 2019; Farm Service Agency, U.S.
+Department of Agriculture, 2019), which cannot be fully explained by soil moisture dynamics.
 In 2020 (Fig.
-11(b)), inaccurate corn yield prediction in Iowa was primarily attributable to the derecho storm’s impact (Hosseini et al., 2020), which soil moisture could not reflect.
-In 2021 (Fig.
-11(c)), drought in North and South Dakota resulted in inaccurate corn yield predictions, and corresponding soil moisture predictions were also found to be inaccurate.
-In 2022 (Fig.
-11(d)), Nebraska, Kansas, and South Dakota experienced drought, and their soil moisture predictions were also inaccurate.
+11(b)), inaccurate corn yield prediction in Iowa was primarily attributable to the derecho storm (Hosseini et al., 2020), which soil moisture could not reflect.
+In 2021 and 2022 (Fig.
+11(c–d)), the model exhibited large corn yield prediction errors in drought-affected regions such as North and South Dakota, Nebraska, and Kansas.
+These yield errors coincided with inaccurate soil moisture predictions in the same areas, suggesting that misrepresentation of soil moisture dynamics under drought conditions was a key contributor to yield prediction inaccuracies.
 In 2023 (Fig.
-11(e)), Iowa and Missouri were affected by drought.
-Iowa’s corn yield and soil moisture prediction were accurate, which indicated that our model performed well.
-But the prediction for Missouri exhibited errors, despite relatively accurate corresponding soil moisture predictions for the state.
-According to report (Ward, 2023), some damaging winds and hail significantly impacted corn field, causing root lodging in late June and early July, particularly around Mooresville and Rockport, Missouri.
-Hail damage was also reported at Garden City, Missouri.
-These types of mechanical damage were not reflected in standard soil moisture data and constituted yield-limiting factors likely contributing to the corn yield prediction inaccuracies in Missouri.
-The analysis indicates that a significant portion of the error in corn yield prediction is related to soil moisture prediction error, given the key role soil moisture plays in corn growth; thus, if soil moisture is not accurately predicted, the accumulated error in modeling will lead to inaccurate corn yield prediction.
+11(e)), prediction errors were observed in Missouri and Minnesota.
+In Missouri, however, the corresponding soil moisture predictions remained relatively accurate.
+Field reports documented damaging winds and hail in late June and early July, particularly around Mooresville, Rockport, and Garden City, which caused root lodging and crop damage (Ward, 2023).
+Because such mechanical damage is not captured by standard soil moisture data, it likely contributed to the yield prediction errors in this region.
+Overall, these results indicate that a significant portion of corn yield prediction error is linked to soil moisture prediction error, underscoring the central role of soil moisture in crop growth.
+By explicitly modeling soil moisture as an intermediate variable, KGML-SM provides a diagnostic layer of interpretability: yield prediction errors can be traced to misrepresentation of soil moisture dynamics.
+This design not only improves robustness under drought conditions but also offers a transparent explanation for when and why yield predictions fail.
 Figure 11: The absolute error map for (1) corn yield prediction and (2) soil moisture prediction from (a) 2019 to (e) 2023.
+5.5 Strengths, limitations, and future work
+Our main contributions are as follows:
+  • 1.
+    This paper introduced the KGML-SM framework, which integrates process-based and ML models for corn yield prediction while explicitly incorporating the influence of soil moisture.
+  • 2.
+    A drought-aware loss function was designed to enhance model performance under drought conditions and mitigate overestimation.
+  • 3.
+    Based on the relationship between drought, soil moisture, and corn yield prediction, we provided interpretability for the prediction errors of the KGML-SM model and offered directions for future model optimization.
+While these contributions highlight the novelty and effectiveness of our approach, it is equally important to recognize its limitations and potential directions for future research.
+First, the study relies on SMAP soil moisture data and APSIM simulations, which are limited in temporal coverage and in the diversity of spatial resolution.
+Incorporating multi-scale data, such as finer field-level observations and broader regional products, could improve the model’s generalization across heterogeneous environments.
+Second, the drought-aware loss function is designed to mitigate overestimation under drought conditions, but it does not address other extreme events such as floods or heatwaves.
+Extending the loss function to include these scenarios would make the framework more comprehensive.
+Finally, the experiments are limited to the U.S.
+Corn Belt, and testing the KGML-SM framework in other regions and on different crops will be essential to evaluate its broader applicability.
+Addressing these limitations will further improve the robustness and generalizability of KGML-SM, ultimately contributing to more reliable and interpretable crop yield prediction under diverse environmental conditions.
 6 Conclusion
-In this study, we proposed the KGML-SM model, where the W2S encoder is designed to capture the influence of weather on soil moisture, and the attention module is employed to weight different input features for final corn yield prediction.
-To address the known issue that drought conditions often lead to yield overestimation, we introduced a drought-aware loss function to mitigate this problem in drought-affected regions.
-We constructed both an APSIM field-level dataset and a GEE county-level dataset, leveraging a process-based model to learn the corn growth process by pretraining the KGML-SM model on the APSIM field-level dataset and finetuning it on the GEE county-level dataset.
-Our analysis covered 12 states in the U.S.
+In this study, we propose the KGML-SM model, where the W2S encoder is designed to capture the influence of weather on soil moisture, and the attention module is employed to weight different input features for final corn yield prediction.
+To address the issue that drought conditions often lead to yield overestimation, we introduce a drought-aware loss function to mitigate this problem in drought-affected regions.
+We construct both an APSIM field-level dataset and a GEE county-level dataset, learning the corn growth process by pretraining KGML-SM on the APSIM field-level dataset and then finetuning it on the GEE county-level dataset.
+Our analysis covers 12 states in the U.S.
 Corn Belt to investigate the impact of soil moisture on corn yield prediction.
-The proposed method consistently outperformed baseline models across multiple test years.
-Furthermore, we studied the spatial and temporal influence of soil moisture through attention visualization, revealing when and where the model places greater focus on soil moisture.
-Finally, based on the relationship between soil moisture and corn yield prediction, we investigated the causes of inaccuracies in corn yield prediction and provided an explanation.
+The proposed method consistently outperforms baseline models across multiple test years.
+Furthermore, we study the spatial and temporal influence of soil moisture through attention visualization, revealing when and where the model places greater focus on soil moisture.
+Finally, based on the relationship between soil moisture and corn yield prediction, we investigate the causes of prediction inaccuracies and provide explanations.
 In future work, we aim to apply transfer learning techniques to adapt models trained on well-studied regions with abundant simulated data to regions with limited data availability.